### v3

# Rigidity results for initial data sets satisfying the dominant energy condition

Abstract.
Our work proves rigidity theorems for initial data sets associated with compact smooth spin manifolds with boundary and with compact convex polytopes, subject to the dominant energy condition. For manifolds with smooth boundary, this is based on the solution of a boundary value problem for Dirac operators. For convex polytopes we use approximations by manifolds with smooth boundary.
Key words and phrases:
Scalar curvature, initial data sets, polyhedra, holographic index theorem, spin geometry, embedding into Minkowski spacetime
2020 Mathematics Subject Classification:
53C20, 53C21, 53C27, 58J20, 58J32
C.B. was supported by DFG-SPP 2026 “Geometry at Infinity” and by the University of Augsburg.
S.B. was supported by the National Science Foundation under grant DMS-2403981 and by the Simons Foundation.
T.-K.A.C. was supported by the Croucher Foundation.
B.H. was supported by DFG-SPP 2026 “Geometry at Infinity”, by Columbia University, by the NYU Courant Institute, by the Isaac Newton Institute and by the University of Potsdam
1. Introduction
An initial data set is a triple $(M,g,q)$, where $M$ is a manifold, $g$ is a Riemannian metric on $M$, and $q$ is a symmetric $(0,2)$-tensor on $M$. Denote the scalar curvature of $g$ by $R$. We define
$\mu=\tfrac{1}{2}\,R+\tfrac{1}{2}\,\mathrm{tr}(q)^{2}-\tfrac{1}{2}\,|q|^{2}$
and
$J=\text{\rm div}(q)-\nabla\mathrm{tr}(q).$
We say that $(M,g,q)$ satisfies the dominant energy condition if $\mu\geq|J|$ at each point in $M$.
Example 1.1.
Let $M$ be a spacelike hypersurface in the Minkowski spacetime $\mathbb{R}^{n,1}$. Let $g$ and $q$ denote the induced metric and second fundamental form on $M$, respectively. Then the initial data set $(M,g,q)$ satisfies $\mu=0$ and $J=0$ by the Gauss and Codazzi-Mainardi equations, respectively. In particular, $(M,g,q)$ satisfies the dominant energy condition.
In this paper we will formulate conditions which insure that a given initial data set is of the form described in this example. The study of rigidity for initial data sets under the dominant energy condition can be traced back to Schoen and Yau’s work [SY81], in which they proved the spacetime positive energy theorem using minimal surface techniques in dimension 3. Eichmair, Huang, Lee, and Schoen [EHLS] later extended this work by proving a more general spacetime positive mass theorem up to dimension 7. Huang and Lee [HL] subsequently affirmed the rigidity statement of this theorem studied in [EHLS]. In a different approach, Witten [Wit] used spinors to prove the spacetime positive mass theorem, a method that generalizes to all higher dimensions for spin manifolds. Parker and Taubes [PT] gave a mathematically rigorous formulation of Witten’s argument. Chruściel and Maerten [CM] later provided a rigorous proof of the desired rigidity statement in all dimensions for spin manifolds, building on Beig and Chruściel’s proof [BC] in dimension 3. Lee’s book [Lee] offers a comprehensive exposition on the positive mass theorem.
We consider the following class of manifolds with boundary, specializing Example 1.1.
Example 1.2.
Let $K\subset\mathbb{R}^{n}$ be a compact convex domain with smooth boundary $\partial K$ and exterior normal $N\colon\partial K\to S^{n-1}$. Note that $N$ has degree $+1$. Let $f\colon K\to\mathbb{R}$ be a smooth map satisfying
  1. (a)
    $|df|<1$ on $K$,
  2. (b)
    $f=0$ along $\partial K$,
  3. (c)
    At each point on $\partial K$, we have $df(N)=0$ or $dN=0$.
Let
[…]
be the graph of $f$. Then $M$ is a compact smooth spacelike hypersurface with boundary $\Sigma=\partial K\times\{0\}\subset\mathbb{R}^{n}\times\{0\}\subset\mathbb{R}^{n,1}$, see Figures 1 and 2.
[…]
Figure 1. The case $df(N)=0$ and $dN\neq 0$
[…]
Figure 2. The case $df(N)\neq 0$ and $dN=0$
Let $I\!I$ denote the vector-valued second fundamental form of $\Sigma$, viewed as a submanifold of $\mathbb{R}^{n,1}$ of codimension $2$. Then
$I\!I(X,Y)=-\langle dN(X),Y\rangle\,N$
for all tangent vectors $X,Y\in T\Sigma$.
We denote by $T$ the future-oriented normal field to $M$ in $\mathbb{R}^{n,1}$, where $T$ is normalized so that $\langle T,T\rangle=-1$. Let $g$ denote the induced Riemannian metric and let $q$ denote the second fundamental form of $M$ with respect to $T$. In other words, $q(X,Y)=\langle\bar{D}_{X}T,Y\rangle$ for all $X,Y\in TM$, where $\bar{D}$ denotes the standard flat connection on $\mathbb{R}^{n,1}$. The initial data set $(M,g,q)$ satisfies the dominant energy condition, cf. Example 1.1. At each point on $\Sigma$, we denote by $\nu$ the outward-pointing unit normal vector to $\Sigma$ in $M$. Moreover, we denote by $h$ the second fundamental form of $\Sigma$, viewed as a hypersurface in $M$, with respect to $\nu$. In other words, $h(X,Y)=\langle D_{X}\nu,Y\rangle$ for $X,Y\in T\Sigma$, where $D$ denotes the Levi-Civita connection on $(M,g)$. We will use these sign conventions for $q$ and $h$ throughout our paper.
With this understood,
$I\!I(X,Y)=-h(X,Y)\,\nu+q(X,Y)\,T$
for all tangent vectors $X,Y\in T\Sigma$. Putting these facts together, we conclude that
$h(X,Y)\,\nu-q(X,Y)\,T=\langle dN(X),Y\rangle\,N$
for all tangent vectors $X,Y\in T\Sigma$. In particular,
$h(X,Y)=\langle dN(X),Y\rangle\,\langle N,\nu\rangle$
and
$q(X,Y)=\langle dN(X),Y\rangle\,\langle N,T\rangle$
for all tangent vectors $X,Y\in T\Sigma$. We now use Assumption (c). Hence, at each point on $\Sigma$, we either have $dN=0$, or else $\langle N,\nu\rangle=1$ and $\langle N,T\rangle=0$. Consequently, $h(X,Y)=\langle dN(X),Y\rangle$ and $q(X,Y)=0$ for all tangent vectors $X,Y\in T\Sigma$. In particular, $H=\|dN\|_{\mathrm{tr}}$ and $\mathrm{tr}(q)-q(\nu,\nu)=0$ at each point on $\Sigma$. Here, $H$ denotes the unnormalized mean curvature, i.e. the sum of the principal curvatures, of $\Sigma$ in $M$. In other words, $H$ is equal to the trace of $h$, considered as an endomorphism field. Moreover, $\|\cdot\|_{\mathrm{tr}}$ denotes the trace norm of a linear map, i.e., the sum of all singular values.
Our first result identifies this class of examples as an extremal case of initial data sets.
Theorem I.
Let $(M,g,q)$ be an initial data set. Assume that $M$ is a compact, connected spin manifold, has dimension $n\geq 2$ and nonempty boundary $\partial M=\Sigma$. Let $N\colon\Sigma\to S^{n-1}$ be a smooth map. We assume that
  • $(M,g,q)$ satisfies the dominant energy condition,
  • $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ holds along $\Sigma$,
  • $N$ has positive degree.
Then $(M,g)$ is isometric to a manifold appearing in Example 1.2 such that $q$ is the second fundamental form of $M$ in $\mathbb{R}^{n,1}$. Moreover, the map $N$ agrees, up to an element of $O(n)$, with the exterior normal $\partial K\to S^{n-1}$.
Remark 1.3.
We do not assume that $\Sigma$ is connected. If $\Sigma$ is not connected, the degree of $N\colon\Sigma\to S^{n-1}$ equals the sum of the mapping degrees of $N$ restricted to the components of $\Sigma$ with their orientations induced from that of $M$.
Remark 1.4.
The quantity $H-|\mathrm{tr}(q)-q(\nu,\nu)|$ is related to the notion of marginally trapped surfaces in general relativity. Recall that a marginally outer trapped surface (MOTS) is a hypersurface in an initial data set $(M,g,q)$ satisfying $H+(\mathrm{tr}(q)-q(\nu,\nu))=0$; see [EHLS]*p. 89.
Remark 1.5.
There is also a variant of Theorem I for $n=1$. In this case, $M=[a,b]$ is a compact interval, $g=dt\otimes dt$ is the standard metric, and $q=\kappa g$ where $\kappa\colon[a,b]\to\mathbb{R}$ is a given smooth function. We can choose $N\colon\partial M=\{a,b\}\to S^{0}=\{-1,1\}$ to be bijective. Then all quantities $R$, $\mu$, $H$, $J$, $dN$, and $\mathrm{tr}(q)-q(\nu,\nu)$ vanish. Thus the assumptions of Theorem I are trivially satisfied.
The conclusion is now that there is a spacelike curve $[a,b]\to\mathbb{R}^{1,1}$, parametrized by arc-length and with curvature $\kappa$. Indeed, this can be shown directly by ODE methods. The analogue for the Euclidean plane $\mathbb{R}^{2}$ instead of $\mathbb{R}^{1,1}$ is known as the fundamental theorem of planar curve theory.
For $q=0$, Theorem I gives the following statement:
Corollary 1.6.
Let $(M,g)$ be a compact connected Riemannian spin manifold of dimension $n\geq 2$ and nonempty boundary $\partial M=\Sigma$. Let $N\colon\Sigma\to S^{n-1}$ be a smooth map. We assume that
  • $R\geq 0$ holds on $M$,
  • $H\geq\|dN\|_{\mathrm{tr}}$ holds along $\Sigma$,
  • $N$ has positive degree.
Then $M$ can be isometrically embedded in $\mathbb{R}^{n}$ as a compact convex domain with smooth boundary and exterior normal $N$.
In the second part of this paper, we consider initial data sets of the form $(\Omega,g,q)$, where $\Omega$ is a compact, convex polytope in $\mathbb{R}^{n}$ with non-empty interior. When $q=0$, the dominant energy condition simplifies to the non-negativity of the scalar curvature $R\geq 0$, and the boundary condition reduces to mean convexity $H\geq 0$, allowing us to compare $(\Omega,g)$ with Euclidean polytopes. Gromov [Gro14] initiated the study of scalar curvature comparison for polytopes with nonnegative scalar curvature to explore the notion of scalar curvature lower bounds in low-regularity spaces, and this problem has been extensively investigated; see [Gro14, Gro19, Gro22, Li17, Li19, WXY]. For $q=g$, the dominant energy condition translates to $R\geq-n(n-1)$. Li [Li20] established a polyhedral comparison result for metrics with scalar curvature at least $-n(n-1)$.
To explore the notion of energy density lower bounds in spacetimes with low regularity, we extend Gromov’s initial work [Gro14] to initial data sets under the dominant energy condition. In the case where $\Omega$ is a cube, Tsang [Tsang] addressed this problem in dimension 3 using the spacetime harmonic function approach.
We will employ the strategy developed by the second named author in [Brendle], adapting it to initial data sets. To achieve this, we utilize a boundary value problem for Dirac operators on smooth domains, which have been extensively studied thanks to the work of Hörmander [Hormander], and the first named author and Ballmann [BB, BB13]. We write the given polytope in the form $\Omega=\bigcap_{i\in I}\{u_{i}\leq 0\}$, where $I$ is a finite set and $u_{i}$ are linear functions on $\mathbb{R}^{n}$. For each $i\in I$, let $N_{i}\in S^{n-1}$ and $\nu_{i}$ denote the outward pointing unit normal vectors to the half-space $\{u_{i}\leq 0\}$ with respect to the Euclidean metric and the metric $g$, respectively. We make the following assumption:
Matching Angle Hypothesis.
If $x\in\partial\Omega$ is a point on the boundary and $i_{1},i_{2}\in I$ satisfy $u_{i_{1}}(x)=u_{i_{2}}(x)=0$, then $\langle\nu_{i_{1}}(x),\nu_{i_{2}}(x)\rangle_{g}=\langle N_{i_{1}},N_{i_{2}}\rangle$. Here the inner product $\langle\nu_{i_{1}},\nu_{i_{2}}\rangle_{g}$ is computed with respect to the metric $g$ and $\langle N_{i_{1}},N_{i_{2}}\rangle$ is the standard Euclidean inner product.
Theorem II.
Assume that $n\geq 2$ is an integer. Let $\Omega$ be a compact, convex polytope in $\mathbb{R}^{n}$ with nonempty interior. Let $g$ be a Riemannian metric which is defined on an open set containing $\Omega$. Let $q$ be a symmetric $(0,2)$-tensor, which is defined on an open set containing $\Omega$. We assume that
  • $(\Omega,g,q)$ satisfies the dominant energy condition,
  • $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|$ on the boundary faces of $\Omega$,
  • the Matching Angle Hypothesis is satisfied.
Then $(\Omega,g,q)$ can be isometrically immersed as a spacelike hypersurface in the Minkowski spacetime $\mathbb{R}^{n,1}$ such that $q$ is the second fundamental form. Moreover, each boundary face of $\Omega$ is totally geodesic as a codimension 2 submanifold in the Minkowski spacetime $\mathbb{R}^{n,1}$.
Conjecture 1.7.
Under the assumptions of Theorem II, there exists a smooth map $f\colon\Omega\to\mathbb{R}$ satisfying
  • $|df|<1$ on $\Omega$,
  • $f=0$ along $\partial\Omega$,
and an isometry of $(\Omega,g)$ to
$\textrm{graph}(f)=\{(x,f(x))\colon x\in\Omega\} \subset\mathbb{R}^{n,1}$
with the induced Riemannian metric. Furthermore, under this isometry, $q$ is the second fundamental form of $\textrm{graph}(f)$ in $\mathbb{R}^{n,1}$.
2. A boundary value problem for the Dirac operator in odd dimensions
Throughout this section, we assume that $M$ is a compact, connected spin manifold of dimension $n\geq 2$ with boundary $\partial M=:\Sigma$. Let $g$ be a Riemannian metric on $M$. Suppose that $\nu$ is the unit outward normal field along $\partial M$ and that $N\colon\partial M\to S^{n-1}$ is a smooth map.
Let $m=2^{[\frac{n}{2}]}$. Let $\mathcal{S}$ denote the spinor bundle over $M$. Note that $\mathcal{S}\to M$ is a complex vector bundle of rank $m$ equipped with a Hermitian inner product and compatible connection $\nabla^{\mathcal{S}}$.
Let $S_{0}$ denote the spinor module of the complex Clifford algebra $\mathbb{C}\mathrm{l}(\mathbb{R}^{n})$ of $\mathbb{R}^{n}$. Now both $S_{0}$ and its dual space $S_{0}^{*}$ are $m$-dimensional Hermitian vector spaces.
We define a complex vector bundle $\mathcal{E}$ over $M$ by $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$. The bundle $\mathcal{E}$ has rank $m^{2}$ and inherits a natural Hermitian inner product and compatible connection $\nabla^{\mathcal{E}}$.
Remark 2.1.
The fiber of $\mathcal{E}\to M$ over $x\in M$ is given by
(2.1) $\mathcal{E}|_{x}=\mathcal{S}|_{x}\otimes S_{0}^{*}=\mathrm{Hom}(S_{0},\mathcal{S}|_{x}).$
Thus, a section $s$ of $\mathcal{E}$ can be considered as a field of homomorphisms. Given an element $\sigma\in S_{0}$ we can apply $s$ at each point $x\in M$ to it and obtain a section $x\mapsto s(x)\sigma$ of $\mathcal{S}$.
Denote by $\gamma\colon TM\to\mathrm{End}(\mathcal{S})$ the Clifford multiplication of $(M,g)$ and by $\gamma_{0}\colon\mathbb{R}^{n}\to\mathrm{End}(S_{0})$ the one of $\mathbb{R}^{n}$. This gives an induced Clifford multiplication $\gamma_{0}^{*}\colon\mathbb{R}^{n}\to\mathrm{End}(S_{0}^{*})$, where $\gamma_{0}^{*}(v)$ is defined as the adjoint of $\gamma_{0}(v)$. We define
(2.2) $\chi\colon\mathcal{E}|_{\Sigma}\to\mathcal{E}|_{\Sigma}\quad\mbox{ by }\quad\chi_{x}:=-\gamma(\nu(x))\otimes\gamma_{0}^{*}(N(x)).$
Since $\gamma(\nu)^{2}=\gamma_{0}^{*}(N)^{2}=-1$ we have
$\chi^{2}=1.$
Since both $\gamma(\nu)$ and $\gamma_{0}^{*}(N)$ are skew-adjoint we get
$\chi^{*}=\chi.$
For $\xi\in T\partial M$ the linear map $\gamma(\xi)\otimes 1$ anticommutes with $\chi$ and is an isomorphism if $\xi\neq 0$. Therefore the ±$\pm 1$-eigenspaces of $\chi$ have the same dimension.
Let $\mathcal{D}$ denote the spinorial Dirac operator on $M$, acting on sections of $\mathcal{E}$. This Dirac operator is formally selfadjoint. By Corollary 7.23 in [BB], the conditions $\chi s=s$ and $\chi s=-s$ both form elliptic boundary conditions for $\mathcal{D}$. Since $\gamma(\nu)\otimes 1$ commutes with $\chi$, the boundary conditions $\chi s=s$ and $\chi s=-s$ are adjoint to each other, see [BB]*Section 7.2.
We denote by $L^{2}(M,\mathcal{E})$ the Hilbert space of square integrable sections of $\mathcal{E}$ and by $H^{1}(M,\mathcal{E})$ the Sobolev space of $L^{2}$-sections whose distributional derivatives are also square integrable. If $\mathcal{F}\subset\mathcal{E}|_{\Sigma}$ is a subbundle of $\mathcal{E}$ along the boundary, we write $H^{1}(M,\mathcal{E},\mathcal{F}):=\{u\in H^{1}(M,\mathcal{E})\mid u|_{\Sigma}\in L^{2}(\Sigma,\mathcal{F})\}$. Note that the restriction map extends uniquely to a bounded linear map $H^{1}(M,\mathcal{E})\to L^{2}(\Sigma,\mathcal{E})$ by the trace theorem.
We decompose the bundle
(2.3) $\mathcal{E}|_{\Sigma}=\mathcal{F}^{+}\oplus\mathcal{F}^{-}$
where $\mathcal{F}^{\pm}$ are the eigensubbundles of $\chi$ corresponding to the eigenvalues ±$\pm 1$. We need the following index computation, which generalizes the results in [Brendle].
Proposition 2.2.
Let $M$ have odd dimension $n\geq 3$. Then the operator
$\mathcal{D}\colon H^{1}(M,\mathcal{E},\mathcal{F}^{+})\to L^{2}(M,\mathcal{E})$
is a Fredholm operator. Its Fredholm index equals the degree of $N\colon\Sigma\to S^{n-1}$.
Proof.
Since $n$ is odd, the restriction $\mathcal{S}|_{\Sigma}$ can be canonically identified with the spinor bundle of $\Sigma$. The field of involutions $i\gamma(\nu)$ can be used to decompose
$\mathcal{S}|_{\Sigma}=\mathcal{S}_{\Sigma}^{+}\oplus\mathcal{S}_{\Sigma}^{-}$
where $\mathcal{S}_{\Sigma}^{\pm}$ are the eigensubbundles corresponding to the eigenvalues ±$\pm 1$. Let $\mathcal{S}_{0}$ denote the trivial bundle over $S^{n-1}$ with fiber $S_{0}$, and let $\mathcal{S}_{0}^{*}$ denote the dual bundle over $S^{n-1}$ with fiber $S_{0}^{*}$. With this understood, we may write $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes N^{*}\mathcal{S}_{0}^{*}$. We may decompose the bundle $\mathcal{S}_{0}^{*}$ as
(2.4) $\mathcal{S}_{0}^{*}=\mathcal{S}_{0}^{*,+}\oplus\mathcal{S}_{0}^{*,-},$
where, for each point $v\in S^{n-1}$, $\mathcal{S}_{0}^{*,\pm}$ denote the eigensubbundles of $i\gamma(v)$ corresponding to the eigenvalues ±$\pm 1$. This gives a decomposition
(2.5) $N^{*}\mathcal{S}^{*}=N^{*}\mathcal{S}_{0}^{*,+}\oplus N^{*}\mathcal{S}_{0}^{*,-},$
where $N^{*}\mathcal{S}_{0}^{*,\pm}$ are the eigensubbundles of $i\gamma_{0}^{*}(N)$ corresponding to the eigenvalues ±$\pm 1$. We then have
$\mathcal{F}^{+}$ $=(\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,+})\oplus(\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,-}),$
$\mathcal{F}^{-}$ $=(\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,-})\oplus(\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,+}).$
Note that even though $\mathcal{S}_{0}^{*}$ is a trivial bundle, the two subbundles $\mathcal{S}_{0}^{*,+}$ and $\mathcal{S}_{0}^{*,-}$ are not because the decomposition depends on the base point. The splitting (2.4) is not parallel with respect to the canonical flat connection on $\mathcal{S}_{0}^{*}$, but the Levi-Civita connection of $S^{n-1}$ induces a connection on the dual of its spinor bundle for which it is. We equip $\mathcal{S}_{0}^{*,\pm}$ with these latter connections and $N^{*}\mathcal{S}_{0}^{*,\pm}$ with the corresponding pull-back connections.
We next consider the twisted Dirac operators
$\mathcal{D}^{\Sigma}_{+}$ $\colon H^{1}(\Sigma,\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,+})\to L^{2}(\Sigma,\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,+}),$
$\mathcal{D}^{\Sigma}_{-}$ $\colon H^{1}(\Sigma,\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,-})\to L^{2}(\Sigma,\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,-}).$
Then the operator
(2.6) $\mathcal{D}^{\Sigma}=\begin{pmatrix}0&\mathcal{D}^{\Sigma}_{-}\\ \mathcal{D}^{\Sigma}_{+}&0\end{pmatrix}$
maps sections of $\mathcal{F}^{+}$ to sections of $\mathcal{F}^{-}$.
In the following, $\hat{\mathsf{A}}(\Sigma)$ denotes the $\hat{\mathsf{A}}$-form of the tangent bundle of $\Sigma$ and $\mathsf{ch}(\mathcal{S}_{0}^{*,\pm})$ the Chern character form of the bundle $\mathcal{S}_{0}^{*,\pm}$. The lower index $k$ in $\hat{\mathsf{A}}(\Sigma)_{k}$ indicates the homogeneous part of degree $k$ and similarly for the Chern character. By the Atiyah-Singer index theorem, we have
$\mathsf{ind}(\mathcal{D}^{\Sigma}_{+})$ $=\int_{\Sigma}\hat{\mathsf{A}}(\Sigma)\wedge N^{*}\mathsf{ch}(\mathcal{S}_{0}^{*,+})$
$=\int_{\Sigma}\hat{\mathsf{A}}(\Sigma)_{0}\wedge N^{*}\mathsf{ch}(\mathcal{S}_{0}^{*,+})_{n-1}+\int_{\Sigma}\hat{\mathsf{A}}(\Sigma)_{n-1}\wedge N^{*}\mathsf{ch}(\mathcal{S}_{0}^{*,+})_{0}$
$=\int_{\Sigma}N^{*}\mathsf{ch}(\mathcal{S}_{0}^{*,+})_{n-1}+\mathsf{rank}(\mathcal{S}_{0}^{*,+})\int_{\Sigma}\hat{\mathsf{A}}(\Sigma)_{n-1}$
$=\mathsf{deg}(N)\int_{S^{n-1}}\mathsf{ch}(\mathcal{S}_{0}^{*,+})_{n-1}+\frac{m}{2}\int_{\Sigma}\hat{\mathsf{A}}(\Sigma)_{n-1}$
and similarly
$-\mathsf{ind}(\mathcal{D}^{\Sigma}_{-})$ $=\mathsf{deg}(N)\int_{S^{n-1}}\mathsf{ch}(\mathcal{S}_{0}^{*,-})_{n-1}+\frac{m}{2}\int_{\Sigma}\hat{\mathsf{A}}(\Sigma)_{n-1}.$
Therefore,
$\mathsf{ind}(\mathcal{D}^{\Sigma})$ $=\mathsf{ind}(\mathcal{D}^{\Sigma}_{+})+\mathsf{ind}(\mathcal{D}^{\Sigma}_{-})$
$=\mathsf{deg}(N)\int_{S^{n-1}}(\mathsf{ch}(\mathcal{S}_{0}^{*,+})-\mathsf{ch}(\mathcal{S}_{0}^{*,-}))_{n-1}$
$=2\,\mathsf{deg}(N).$
Now, we observe that $\mathcal{D}$ and $\mathcal{D}^{\Sigma}$ are formally selfadjoint, $\mathcal{D}^{\Sigma}$ is an adapted boundary operator for $\mathcal{D}$, the involution $i\gamma(\nu)\otimes 1$ preserves the splitting (2.3) and anticommutes with $\mathcal{D}^{\Sigma}$. Therefore, the holographic index theorem from Appendix B in [BBHW23] applies and tells us that the operator
$\mathcal{D}\colon H^{1}(M,\mathcal{E},\mathcal{F}^{+})\to L^{2}(M,\mathcal{E})$
is Fredholm and
$\mathsf{ind}(\mathcal{D}\colon H^{1}(M,\mathcal{E},\mathcal{F}^{+})\to L^{2}(M,\mathcal{E}))=\tfrac{1}{2}\,\mathsf{ind}(\mathcal{D}^{\Sigma})=\mathsf{deg}(N).\qed$
3. Spin geometry for initial data sets
Let $M$ be a compact spin manifold of dimension $n\geq 2$ with boundary $\partial M=\Sigma$. Let $g$ be a Riemannian metric on $M$, let $q$ be a symmetric $(0,2)$-tensor field on $M$ and let $N\colon\Sigma\to S^{n-1}$ be a smooth map. Let $\chi$ be defined as in (2.2). We will denote the selfadjoint endomorphism field corresponding to $q$ also by $q\colon TM\to TM$, i.e.
$q(X,Y)=g(q(X),Y).$
Furthermore, we put
(3.1) $\mu:=\tfrac{1}{2}\,R+\tfrac{1}{2}\,\mathrm{tr}(q)^{2}-\tfrac{1}{2}\,|q|^{2}$
and
(3.2) $J:={\mathrm{div}}(q)-\nabla\mathrm{tr}(q)$
where $R$ is the scalar curvature of $g$. We use $q$ to modify the connections on $\mathcal{S}$ and on $\mathcal{E}$.
Definition 3.1.
We define a connection $\nabla^{\mathcal{S},q}$ on $\mathcal{S}$ by
(3.3) $\nabla^{\mathcal{S},q}:=\nabla^{\mathcal{S}}+\frac{i}{2}\gamma\circ q$
on $\mathcal{S}$. We next introduce the homomorphism field $Q\colon\mathcal{E}\to T^{*}M\otimes\mathcal{E}$ which maps spinors to spinor-valued $1$-forms by setting
(3.4) $(Qs)(X):=(\gamma(q(X))\otimes 1)s$
for every section $s$ of $\mathcal{E}$ and every tangent vector $X$. With this understood, we define a connection $\nabla^{\mathcal{E},q}$ on $\mathcal{E}$ by
(3.5) $\nabla^{\mathcal{E},q}:=\nabla^{\mathcal{E}}+\tfrac{i}{2}Q.$
Note that the connection $\nabla^{\mathcal{E},q}$ on $\mathcal{E}$ is the tensor product connection of $\nabla^{\mathcal{S},q}$ with the flat connection on the trivial bundle over $M$ with fiber $S_{0}$.
In order to derive a modified Weitzenböck formula involving $\nabla^{\mathcal{E},q}$ we need some preparation.
Lemma 3.2.
Denote the (pointwise) adjoint of $Q$ by $Q^{*}\colon T^{*}M\otimes\mathcal{E}\to\mathcal{E}$. Then we have
  1. (a)
    $Q^{*}Q=|q|^{2}$,
  2. (b)
    $Q^{*}\nabla^{\mathcal{E}}-(\nabla^{\mathcal{E}})^{*}Q=\gamma(\mathrm{div}(q))\otimes 1$.
Proof.
Let $p\in M$ and let $e_{1},\dots,e_{n}$ be an orthonormal basis of $T_{p}M$. Denote the dual basis of $T^{*}_{p}M$ by $e^{1},\dots,e^{n}$. Then we have for $s\in\mathcal{E}|_{p}$ and $\psi\in T^{*}_{p}M\otimes\mathcal{E}|_{p}$:
$\langle Q^{*}\psi,s\rangle$ $=\langle\psi,Qs\rangle=\Big{\langle}\psi,\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes 1)s\Big{\rangle}$
$=\sum_{j}\langle\psi(e_{j}),(\gamma(q(e_{j}))\otimes 1)s\rangle$
$=-\sum_{j}\langle(\gamma(q(e_{j}))\otimes 1)\psi(e_{j}),s\rangle$
and hence
$Q^{*}\psi=-\sum_{j}(\gamma(q(e_{j}))\otimes 1)\psi(e_{j}).$
Therefore,
$Q^{*}Qs$ $=Q^{*}\Big{(}\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes 1)s\Big{)}$
$=-\sum_{j}(\gamma(q(e_{j}))\otimes 1)(\gamma(q(e_{j}))\otimes 1)s$
$=\sum_{j}|q(e_{j})|^{2}s=|q|^{2}s.$
This proves (a). To show (b), extend the orthonormal basis to a neighborhood of $p$ such that it is synchronous at $p$. Let $s$ be a smooth section of $\mathcal{E}$, defined near $p$. Then we compute at $p$:
$(\nabla^{\mathcal{E}})^{*}Qs$ $=(\nabla^{\mathcal{E}})^{*}\Big{(}\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes 1)s\Big{)}$
$=-\sum_{j}(\gamma((D_{e_{j}}q)(e_{j}))\otimes 1)s-\sum_{j}(\gamma(q(e_{j}))\otimes 1)\nabla_{e_{j}}^{\mathcal{E}}s$
$=-(\gamma(\mathrm{div}(q))\otimes 1)s+Q^{*}\nabla^{\mathcal{E}}s.\qed$
The well-known Weitzenböck formula gives
$\mathcal{D}^{2}=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}+\tfrac{1}{4}R.$
We find a modification, taking $q$ into account.
Proposition 3.3 (Modified Weitzenböck formula).
We have the following operator identity:
$(\mathcal{D}+\tfrac{i}{2}\mathrm{tr}(q))(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))=(\nabla^{\mathcal{E},q})^{*}\nabla^{\mathcal{E},q}+\tfrac{1}{2}\mu+\tfrac{i}{2}\gamma(J)\otimes 1.$
Proof.
We compute, using Lemma 3.2:
$(\nabla^{\mathcal{E},q})^{*}\nabla^{\mathcal{E},q}$ $=((\nabla^{\mathcal{E}})^{*}-\tfrac{i}{2}Q^{*})(\nabla^{\mathcal{E}}+\tfrac{i}{2}Q)$
$=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}-\tfrac{i}{2}(Q^{*}\nabla^{\mathcal{E}}-(\nabla^{\mathcal{E}})^{*}Q)+\tfrac{1}{4}Q^{*}Q$
$=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}-\tfrac{i}{2}\gamma(\mathrm{div}(q))\otimes 1+\tfrac{1}{4}|q|^{2}.$
Furthermore,
$(\mathcal{D}+\tfrac{i}{2}\mathrm{tr}(q))(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))$ $=\mathcal{D}^{2}-\tfrac{i}{2}[\mathcal{D},\mathrm{tr}(q)]+\tfrac{1}{4}\mathrm{tr}(q)^{2}$
$=\mathcal{D}^{2}-\tfrac{i}{2}\gamma(\nabla\mathrm{tr}(q))\otimes 1+\tfrac{1}{4}\mathrm{tr}(q)^{2}.$
These two equations together with the standard Weitzenböck formula yield
$(\mathcal{D}+\tfrac{i}{2}\mathrm{tr}(q))(\mathcal{D}$ $-\tfrac{i}{2}\mathrm{tr}(q))-(\nabla^{\mathcal{E},q})^{*}\nabla^{\mathcal{E},q}$
$=\tfrac{1}{4}R+\tfrac{i}{2}\gamma(\mathrm{div}(q)-\nabla\mathrm{tr}(q))\otimes 1+\tfrac{1}{4}(\mathrm{tr}(q)^{2}-|q|^{2})$
$=\tfrac{1}{2}\mu+\tfrac{i}{2}\gamma(J)\otimes 1.\qed$
Again we denote $\partial M=\Sigma$ and the unit outward normal along $\Sigma$ by $\nu$. Let $H$ denote the (unnormalized) mean curvature. The sign convention is such that the boundary of a Euclidean ball has positive mean curvature.
We define an operator $\mathcal{D}^{\Sigma}_{0}$ acting on sections of $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes S_{0}^{*}$ by
$\mathcal{D}^{\Sigma}_{0}=-\sum_{j=1}^{n-1}(\gamma(\nu)\otimes 1)(\gamma(e_{j})\otimes 1)\,\nabla_{e_{j}}^{\mathcal{E}}-\tfrac{1}{2}H,$
where $e_{1},\ldots,e_{n-1}$ denotes a local orthonormal frame on $\Sigma$ and $\nabla^{\mathcal{E}}$ denotes the connection on $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$ defined above. By Proposition 2.2 in [B96], $\mathcal{D}^{\Sigma}_{0}$ can be identified with a twisted Dirac operator acting on sections of $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes N^{*}\mathcal{S}_{0}^{*}$, where the twist bundle $N^{*}\mathcal{S}_{0}^{*}$ carries the flat connection. Note that $\mathcal{D}^{\Sigma}_{0}$ differs from the operator $\mathcal{D}^{\Sigma}$ defined in (2.6). In the definition of $\mathcal{D}^{\Sigma}$, we used a nonflat connection on the twist bundle $N^{*}\mathcal{S}_{0}^{*}$ that preserves the splittting (2.5). The operators $\mathcal{D}^{\Sigma}$ and $\mathcal{D}^{\Sigma}_{0}$ both anticommute with $\gamma(\nu)\otimes 1$.
Corollary 3.4.
For all $s\in H^{1}(M,\mathcal{E})$ we have
$-\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}$ $+\int_{M}|\nabla^{\mathcal{E},q}s|^{2}+\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes 1)s,s\rangle$
$=\int_{\Sigma}\langle(-\mathcal{D}^{\Sigma}_{0}-\tfrac{1}{2}H+\tfrac{i}{2}(\gamma(q(\nu)-\mathrm{tr}(q)\nu)\otimes 1))s,s\rangle.$
Proof.
Note that
$(\gamma(\nu)\otimes 1)\mathcal{D}s=-\mathcal{D}^{\Sigma}_{0}s-\nabla_{\nu}^{\mathcal{E}}s-\tfrac{1}{2}Hs$
along $\Sigma$. Using this and integrating by parts we find
$\int_{M}$ $\langle(\mathcal{D}+\tfrac{i}{2}\mathrm{tr}(q))(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s,s\rangle$
$=\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{\Sigma}\langle(\gamma(\nu)\otimes 1)(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s,s\rangle$
$=\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{\Sigma}\big{\langle}\big{(}-\mathcal{D}^{\Sigma}_{0}-\nabla_{\nu}^{\mathcal{E}}-\tfrac{1}{2}H-\tfrac{i}{2}\mathrm{tr}(q)(\gamma(\nu)\otimes 1)\big{)}s,s\big{\rangle}.$
Similarly,
$\int_{M}\langle(\nabla^{\mathcal{E},q})^{*}\nabla^{\mathcal{E},q}s,s\rangle$ $=\int_{M}|\nabla^{\mathcal{E},q}s|^{2}-\int_{\Sigma}\langle\nabla_{\nu}^{\mathcal{E},q}s,s\rangle$
$=\int_{M}|\nabla^{\mathcal{E},q}s|^{2}-\int_{\Sigma}\langle(\nabla_{\nu}^{\mathcal{E}}+\tfrac{i}{2}\gamma(q(\nu))\otimes 1)s,s\rangle.$
Inserting these two equations into the modified Weitzenböck formula concludes the proof. ∎
Lemma 3.5.
Along $\Sigma$ we have
$\mathcal{D}^{\Sigma}_{0}\chi+\chi\mathcal{D}^{\Sigma}_{0}=\sum_{j=1}^{n-1}\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j}))$
where $e_{1},\dots,e_{n-1}$ is a local orthonormal tangent frame to $\Sigma$.
Proof.
Recall that $\mathcal{D}^{\Sigma}_{0}$ anticommutes with $\gamma(\nu)\otimes 1$. Hence, we compute for a smooth section $s$ of $\mathcal{E}|_{\Sigma}$:
$\mathcal{D}^{\Sigma}_{0}\chi s$ $=-\mathcal{D}^{\Sigma}_{0}(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))s$
$=(\gamma(\nu)\otimes 1)\mathcal{D}^{\Sigma}_{0}(1\otimes\gamma_{0}^{*}(N))s$
$=\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)\,\nabla_{e_{j}}^{\mathcal{E}}((1\otimes\gamma_{0}^{*}(N))s)-\tfrac{1}{2}H\,(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))s$
$=\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(N))\,\nabla_{e_{j}}^{\mathcal{E}}s-\tfrac{1}{2}H\,(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))s$
$\quad+\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(dN(e_{j})))s$
$=(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))\,\mathcal{D}^{\Sigma}_{0}s+\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(dN(e_{j})))s$
$=-\chi\mathcal{D}^{\Sigma}_{0}s+\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(dN(e_{j})))s.\qed$
Lemma 3.6.
We have
$\Big{|}\Big{\langle}\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j})))s,s\Big{\rangle}\Big{|}\leq\|dN\|_{\mathrm{tr}}\,|s|^{2}$
at each point on $\Sigma$. Here, $e_{1},\dots,e_{n-1}$ denotes a local orthonormal frame on $\Sigma$.
Proof.
Let us fix a point $p\in\Sigma$. We choose orthonormal bases $e_{1},\dots,e_{n-1}$ of $T_{p}\Sigma$ and $e^{0}_{1},\dots,e^{0}_{n-1}$ of $T_{N(p)}S^{n-1}$ such that $dN(e_{j})=\lambda_{j}e^{0}_{j}$, where $\lambda_{j}\geq 0$ denote the singular values of $dN$. Then
$\Big{|}\Big{\langle}\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j})))s,s\Big{\rangle}\Big{|}$ $=\Big{|}\sum_{j=1}^{n-1}\lambda_{j}\,\langle(\gamma(e_{j})\otimes\gamma_{0}^{*}(e^{0}_{j}))s,s\rangle\Big{|}$
$\leq\sum_{j=1}^{n-1}\lambda_{j}\,|(\gamma(e_{j})\otimes\gamma_{0}^{*}(e^{0}_{j}))s|\,|s|$
$=\sum_{j=1}^{n-1}\lambda_{j}\,|s|^{2}.\qed$
Proposition 3.7.
Suppose that $s\in H^{1}(M,\mathcal{E})$ satisfies the boundary condition $\chi s=\varepsilon s$ along $\Sigma$, where $\varepsilon\colon\Sigma\to\{-1,1\}$ is locally constant. Then we have the estimate
$-\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{M}$ $|\nabla^{\mathcal{E},q}s|^{2}+\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes 1)s,s\rangle$
$\leq\tfrac{1}{2}\int_{\Sigma}(\|dN\|_{\mathrm{tr}}-H+|q(\nu,\nu)-\mathrm{tr}(q)|)|s|^{2}.$
Proof.
We first observe that
$2\,\langle\mathcal{D}^{\Sigma}_{0}s,s\rangle=\varepsilon\,\langle\mathcal{D}^{\Sigma}_{0}\chi s,s\rangle+\varepsilon\,\langle\mathcal{D}^{\Sigma}_{0}s,\chi s\rangle=\varepsilon\,\langle\mathcal{D}^{\Sigma}_{0}\chi s+\chi\mathcal{D}^{\Sigma}_{0}s,s\rangle$
at each point on $\Sigma$. Using Lemma 3.5, we obtain
$2\,|\langle\mathcal{D}^{\Sigma}_{0}s,s\rangle|=\Big{|}\Big{\langle}\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j})))s,s\Big{\rangle}\Big{|}$
at each point on $\Sigma$. Lemma 3.6 implies
(3.6) $2\,|\langle\mathcal{D}^{\Sigma}_{0}s,s\rangle|\leq\|dN\|_{\mathrm{tr}}\,|s|^{2}$
at each point on $\Sigma$. Moreover, for each $X\in T\Sigma$ we have
$\langle(\gamma(X)\otimes 1)s,s\rangle$ $=\varepsilon\langle(\gamma(X)\otimes 1)\chi s,s\rangle$
$=-\varepsilon\langle\chi(\gamma(X)\otimes 1)s,s\rangle$
$=-\varepsilon\langle(\gamma(X)\otimes 1)s,\chi s\rangle$
$=-\langle(\gamma(X)\otimes 1)s,s\rangle,$
hence
$\langle(\gamma(X)\otimes 1)s,s\rangle=0.$
Putting $X=q(\nu)^{\text{\rm tan}}$, the part of $q(\nu)$ tangential to $\Sigma$, gives
$\langle(\gamma(q(\nu))\otimes 1)s,s\rangle=q(\nu,\nu)\,\langle(\gamma(\nu)\otimes 1)s,s\rangle.$
We conclude
$\big{|}\langle(\gamma(q(\nu)-\mathrm{tr}(q)\nu)\otimes 1)s,s\rangle\big{|}$ $=\big{|}(q(\nu,\nu)-\mathrm{tr}(q))\,\langle(\gamma(\nu)\otimes 1)s,s\rangle\big{|}$
(3.7) $\leq|q(\nu,\nu)-\mathrm{tr}(q)|\,|s|^{2}.$
Inserting (3.6) and (3.7) into Corollary 3.4 concludes the proof. ∎
Proposition 3.8.
Suppose that $M$ has odd dimension $n\geq 3$ and that the map $N\colon\Sigma\to S^{n-1}$ has positive degree. Then there exists a non-trivial section $s\in C^{\infty}(M,\mathcal{E})$ such that
$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}$
$\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
Moreover, $\chi s=s$ holds at each point on the boundary $\Sigma$.
Proof.
By Proposition 2.2, the operator $\mathcal{D}$ subject to the boundary conditions $\chi s=s$ has positive index. By deformation invariance of the Fredholm index, the same holds for the operator $\mathcal{D}-\frac{i}{2}\mathrm{tr}(q)$. Hence we can find a non-trivial section $s$ of $\mathcal{E}$ such that $\mathcal{D}s-\frac{i}{2}\,\mathrm{tr}(q)\,s=0$ on $M$ and $\chi s=s$ along $\Sigma$. By elliptic regularity, $s$ is smooth. Using Proposition 3.7, we obtain
$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq-\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes 1)s,s\rangle$
$\qquad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
From this, the assertion follows easily. ∎
In the next step, we establish an analogue of Proposition 3.8 in even dimensions. In the remainder of this section, we will assume that $n\geq 2$ is an even integer.
Lemma 3.9.
We can find a smooth map $\varphi\colon S^{n-1}\times S^{1}\to S^{n}$ with the following properties:
  • The map $\varphi$ has degree $1$.
  • $\varphi^{*}g_{S^{n}}\leq g_{S^{n-1}}+4g_{S^{1}}$.
  • For each $t\in S^{1}$, the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is either Lipschitz continuous with a Lipschitz constant strictly less than $1$, or else $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n}$, $x\mapsto(x,0)$.
Proof.
Fix a smooth $2$-Lipschitz function $\beta\colon[-\pi,\pi]\to[-\pi,\pi]$ such that $\beta(t)=-\pi$ for $t$ near $-\pi$, $\beta(t)=0$ for $t$ near 00, and $\beta(t)=\pi$ for $t$ near $\pi$. Moreover, let us fix a point $a\in S^{n-1}$. We consider the map
$S^{n-1}\times[-\pi,\pi]\to S^{n},\,(x,t)\mapsto\begin{cases}(\sin\beta(t)\,a,\cos\beta(t))&\text{\rm for $t\in[-\pi,0]$},\\ (\sin\beta(t)\,x,\cos\beta(t))&\text{\rm for $t\in[0,\pi]$.}\end{cases}$
This gives a map $\varphi\colon S^{n-1}\times S^{1}\to S^{n}$ with the required properties. ∎
In the following, we fix a smooth map $\varphi\colon S^{n-1}\times S^{1}\to S^{n}$ as in Lemma 3.9. Moreover, $r$ will denote a parameter which will be chosen sufficiently large. We consider the product $\tilde{M}=M\times S^{1}$, equipped with the metric $\tilde{g}=g+r^{2}\,dt^{2}$. We denote by $\tilde{q}$ the pull-back $(0,2)$-tensor field of $q$ under the canonical projection $\tilde{M}\to M$. The boundary of $\tilde{M}$ can be identified with $\tilde{\Sigma}=\Sigma\times S^{1}$. We define a map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ by $\tilde{N}(x,t)=\varphi(N(x),t)$ for $x\in\Sigma$ and $t\in S^{1}$. Since the map $\varphi\colon S^{n-1}\times S^{1}\to S^{n}$ is of degree $1$, the degree of the map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ agrees with the degree of the map $N\colon\Sigma\to S^{n-1}$.
Lemma 3.10.
The trace norms of $dN$ and $d\tilde{N}$ are related by
$\|d\tilde{N}\|_{\mathrm{tr}}\leq\|dN\|_{\mathrm{tr}}+2r^{-1}.$
Proof.
By Lemma 3.9, we have $\varphi^{*}g_{S^{n}}\leq g_{S^{n-1}}+4g_{S^{1}}$. Using the definition of $\tilde{N}$, we obtain $\tilde{N}^{*}g_{S^{n}}\leq N^{*}g_{S^{n-1}}+4g_{S^{1}}$. Let us denote the singular values of $N\colon T_{x}\Sigma\to T_{N(x)}S^{n-1}$ by $\lambda_{1},\ldots,\lambda_{n-1}$. Then the eigenvalues of $N^{*}g_{S^{n-1}}$ with respect to the metric $g$ are given by $\lambda_{1}^{2},\ldots,\lambda_{n-1}^{2}$. Hence, the eigenvalues of $N^{*}g_{S^{n-1}}+4g_{S^{1}}$ with respect to the metric $\tilde{g}$ are given by $\lambda_{1}^{2},\ldots,\lambda_{n-1}^{2},4r^{-2}$. Consequently,
$\|d\tilde{N}\|_{\mathrm{tr}}\leq\sum_{j=1}^{n-1}\lambda_{j}+2r^{-1}=\|dN\|_{\mathrm{tr}}+2r^{-1}.\qed$
We consider the bundles $\mathcal{S}$ and $\mathcal{E}$ over $M$ defined above. Let $\tilde{\mathcal{S}}$ denote the pull-back of $\mathcal{S}$ under the canonical projection $\tilde{M}\to M$. Similarly, let $\tilde{\mathcal{E}}$ denote the pull-back of $\mathcal{E}$ under the canonical projection $\tilde{M}\to M$. Note that $\tilde{\mathcal{S}}$ and $\tilde{\mathcal{E}}$ are bundles over $\tilde{M}$. The map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ induces a chirality operator $\tilde{\chi}$. We define $\tilde{\mu}$ and $\tilde{J}$ as in (3.1) and (3.2) on $\tilde{M}$ using $\tilde{g}$ and $\tilde{q}$. Then $\tilde{\mu}$ and $\tilde{J}$ are the pull-backs of $\mu$ and $J$, respectively.
Proposition 3.11.
Suppose that $M$ has even dimension $n\geq 2$ and that the map $N\colon\Sigma\to S^{n-1}$ has positive degree. For each $r>0$, we can find an element $\hat{t}\in S^{1}$ and a non-trivial section $s\in C^{\infty}(M,\mathcal{E})$ such that
$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}$
$\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}+r^{-1}\int_{\Sigma}|s|^{2}.$
Moreover, we have $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ along $\Sigma$. Here, the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$.
Proof.
Since the map $N\colon\Sigma\to S^{n-1}$ has positive degree, it follows that the map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ has positive degree as well. By Proposition 3.8, we can find a section $\tilde{s}\in C^{\infty}(\tilde{M},\tilde{\mathcal{E}})$ such that
$\int_{\tilde{M}}|\tilde{s}|^{2}>0$
and
$\int_{\tilde{M}}|\nabla^{\tilde{\mathcal{E}},\tilde{q}}\tilde{s}|^{2}$ $\leq\tfrac{1}{2}\int_{\tilde{M}}(-\tilde{\mu}+|\tilde{J}|)|s|^{2}$
$\quad+\tfrac{1}{2}\int_{\tilde{\Sigma}}(|\mathrm{tr}(\tilde{q})-\tilde{q}(\tilde{\nu},\tilde{\nu})|+\|d\tilde{N}\|_{\mathrm{tr}}-\tilde{H})\,|\tilde{s}|^{2}.$
Moreover, $\tilde{s}$ satisfies $\tilde{\chi}\tilde{s}=\tilde{s}$ along the boundary $\tilde{\Sigma}$. Consequently, there exists an element $\hat{t}\in S^{1}$ such that
$\int_{M\times\{\hat{t}\}}|\tilde{s}|^{2}>0$
and
$\int_{M\times\{\hat{t}\}}|\nabla^{\tilde{\mathcal{E}},\tilde{q}}\tilde{s}|^{2}$ $\leq\tfrac{1}{2}\int_{M\times\{\hat{t}\}}(-\tilde{\mu}+|\tilde{J}|)|s|^{2}$
$\quad+\tfrac{1}{2}\int_{\tilde{\Sigma}}(|\mathrm{tr}(\tilde{q})-\tilde{q}(\tilde{\nu},\tilde{\nu})|+\|d\tilde{N}\|_{\mathrm{tr}}-\tilde{H})\,|\tilde{s}|^{2}.$
It follows from Lemma 3.10 that $\|d\tilde{N}\|_{\mathrm{tr}}\leq\|dN\|_{\mathrm{tr}}+2r^{-1}$. Restricting $\tilde{s}$ to $M\times\{\hat{t}\}$ yields a section $s\in C^{\infty}(M,\mathcal{E})$ such that
$\int_{M}|s|^{2}>0$
and
$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}$
$\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}+r^{-1}\int_{\Sigma}|s|^{2}.$
Moreover, since $\tilde{\chi}\tilde{s}=\tilde{s}$ along the boundary $\tilde{\Sigma}$, we conclude that $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ at each point on $\Sigma$, where the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$. ∎
Corollary 3.12.
Suppose that $M$ has even dimension $n\geq 2$ and that the map $N\colon\Sigma\to S^{n-1}$ has positive degree. We can find an element $\hat{t}\in S^{1}$ and a non-trivial section $s\in H^{1}(M,\mathcal{E})$ such that
$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}$
$\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
Moreover, we have $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ along $\Sigma$. Here, the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$.
Proof.
Consider an arbitrary sequence $r_{l}\to\infty$. For each $l$, Proposition 3.11 implies the existence of a section $s^{(l)}\in C^{\infty}(M,\mathcal{E})$ such that
$\int_{M}|s^{(l)}|^{2}=1$
and
$\int_{M}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s^{(l)}|^{2}$
$\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}+r_{l}^{-1}\int_{\Sigma}|s^{(l)}|^{2}.$
Moreover, we can find an element $\hat{t}_{l}\in S^{1}$ such that $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{t}_{l})$ along $\Sigma$, where the map $\hat{N}_{l}:\Sigma\to S^{n}$ is defined by $\hat{N}_{l}(x)=\tilde{N}(x,\hat{t}_{l})=\varphi(N(x),\hat{t}_{l})$ for all $x\in\Sigma$. This implies
$\int_{M}|\nabla^{\mathcal{E}}s^{(l)}|^{2}\leq C\int_{M}|s^{(l)}|^{2}+C\int_{\Sigma}|s^{(l)}|^{2},$
where $C$ is a constant that does not depend on $l$. Using a standard interpolation inequality, we can bound
$\int_{\Sigma}|s^{(l)}|^{2}\leq C\,\bigg{(}\int_{M}|s^{(l)}|^{2}\bigg{)}^{\frac{1}{2}}\,\bigg{(}\int_{M}|\nabla^{\mathcal{E}}s^{(l)}|^{2}\bigg{)}^{\frac{1}{2}}+C\int_{M}|s^{(l)}|^{2},$
where $C$ is a constant that does not depend on $l$. Putting these facts together, we obtain
$\int_{M}|\nabla^{\mathcal{E}}s^{(l)}|^{2}\leq C\,\bigg{(}\int_{M}|s^{(l)}|^{2}\bigg{)}^{\frac{1}{2}}\,\bigg{(}\int_{M}|\nabla^{\mathcal{E}}s^{(l)}|^{2}\bigg{)}^{\frac{1}{2}}+C\int_{M}|s^{(l)}|^{2},$
where $C$ is a constant that does not depend on $l$. Using Young’s inequality, we conclude that
$\int_{M}|\nabla^{\mathcal{E}}s^{(l)}|^{2}\leq C\int_{M}|s^{(l)}|^{2},$
where $C$ is a constant that does not depend on $l$. Since $\int_{M}|s^{(l)}|^{2}=1$ for each $l$, it follows that the sequence $s^{(l)}$ is uniformly bounded in $H^{1}(M,\mathcal{E})$.
After passing to a subsequence if necessary, we may assume that $\hat{t}_{l}\to\hat{t}$ for some $\hat{t}\in S^{1}$. Moreover, we may assume that the sequence $s^{(l)}$ converges weakly in $H^{1}(M,\mathcal{E})$ to some limit $s\in H^{1}(M,\mathcal{E})$. Then $s_{l}\to s$ strongly in $L^{2}(M,\mathcal{E})$. Moreover, $s_{l}|_{\Sigma}\to s|_{\Sigma}$ strongly in $L^{2}(\Sigma,\mathcal{E}|_{\Sigma})$. Thus, we conclude that
$\int_{M}|s|^{2}=1$
and
$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}$
$\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
Finally, on the boundary $\Sigma$, we have $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}_{l})$ for each $l$. Passing to the limit as $l\to\infty$, we conclude that $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$, where the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$. ∎
4. Proof of Theorem I
Let $M$, $\Sigma$, $g$, $q$, $N$, $H$ satisfy the assumptions in Theorem I. In particular, we assume that $\mu\geq|J|$ at each point in $M$ and $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ at each point on the boundary $\Sigma$.
Before we prove Theorem I, we prove some preliminary statements in Proposition 4.1.
We denote the second fundamental form of $\Sigma$ in $M$ (viewed as a (1,1)-tensor) by $h$. For each $X\in T\Sigma$, we denote by $q(X)^{\text{\rm tan}}$ the part of $q(X)\in TM$ tangential to $\Sigma$.
Proposition 4.1.
The following assertions hold.
  1. (a)
    There exists a $\nabla^{\mathcal{E},q}$-parallel smooth section $s\in C^{\infty}(M,\mathcal{E})$ satisfying $\chi s=s$ along $\Sigma$. Considered as a homomorphism field, the section $s$ is invertible at each point.
  2. (b)
    At each point $p\in\Sigma$, the eigenvalues of the second fundamental form $h$ agree with the singular values of $dN_{p}:T_{p}\Sigma\to T_{N(p)}S^{n-1}$. In particular, the eigenvalues of $h$ are nonnegative.
  3. (c)
    $q^{\rm tan}=0$ along $\Sigma$.
We prove Proposition 4.1 first in odd and then in even dimensions.
4.1. Proof of Proposition 4.1 in odd dimensions
Suppose that $n\geq 3$ is an odd integer. Since by assumption $\mu\geq|J|$ on $M$ and $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ along $\Sigma$, Proposition 3.8 yields non-trivial $\nabla^{\mathcal{E},q}$-parallel section $s\in C^{\infty}(M,\mathcal{E})$ such that $\chi s=s$ holds along $\Sigma$.
Lemma 4.2.
At each point $p\in\Sigma$, the second fundamental form $h$ and the tensor field $q$ satisfy the relation
$\gamma(h(X))\circ s+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\circ s=s\circ\gamma_{0}(dN(X))$
for all $X\in T_{p}\Sigma$.
Proof.
Since $s$ is $\nabla^{\mathcal{E},q}$-parallel,
(4.1) $\nabla_{X}^{\mathcal{E}}s=-\tfrac{i}{2}Qs(X)=-\tfrac{i}{2}\gamma(q(X))\circ s$
for every vector field $X$ on $M$. Using the boundary condition $\chi s=s$, we obtain
(4.2) $\gamma(\nu)\circ s=s\circ\gamma_{0}(N)$
on $\Sigma$. In the next step, we differentiate (4.2) along a vector field $X\in C^{\infty}(\Sigma,T\Sigma)$. Using (4.1) we find for the left hand side
$\nabla_{X}^{\mathcal{E}}(\gamma(\nu)\circ s)$ $=\gamma(D_{X}\nu)\circ s+\gamma(\nu)\circ\nabla_{X}^{\mathcal{E}}s$
$=\gamma(h(X))\circ s-\tfrac{i}{2}\gamma(\nu)\circ\gamma(q(X))\circ s$
and for the right hand side
$\nabla_{X}^{\mathcal{E}}(s\circ\gamma_{0}(N))$ $=(\nabla_{X}^{\mathcal{S}}s)\circ\gamma_{0}(N)+s\circ\gamma_{0}(dN(X))$
$=-\tfrac{i}{2}\gamma(q(X))\circ s\circ\gamma_{0}(N)+s\circ\gamma_{0}(dN(X)).$
Thus (4.2) implies
$\gamma(h(X))$ $\circ s-s\circ\gamma_{0}(dN(X))$
$=-\tfrac{i}{2}\big{(}\gamma(q(X))\circ s\circ\gamma_{0}(N)-\gamma(\nu)\circ\gamma(q(X))\circ s\big{)}$
$=-\tfrac{i}{2}\big{(}\gamma(q(X))\circ\gamma(\nu)\circ s-\gamma(\nu)\circ\gamma(q(X))\circ s\big{)}$
$=-i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\circ s.$
For the last equality, we have used the fact that the normal component of $q(X)$ commutes with $\gamma(\nu)$ while the tangential component of $q(X)$ anticommutes with $\gamma(\nu)$. ∎
Lemma 4.3.
We have
$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}=\|dN\|_{\mathrm{tr}}\,|s|^{2}$
at each point on $\Sigma$.
Proof.
Let us fix a point $p\in\Sigma$. Let $e_{1},\dots,e_{n-1}$ be an orthonormal basis of $T_{p}\Sigma$. Using Lemma 4.2, we obtain
[…]
$=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}.$
This implies
$H\,|s|^{2}+i\,(\mathrm{tr}(q)-q(\nu,\nu))\,\langle\gamma(\nu)\circ s,s\rangle=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}.$
By assumption, $H-|\mathrm{tr}(q)-q(\nu,\nu)|\geq\|dN\|_{\mathrm{tr}}$ at each point on $\Sigma$. Consequently,
$\|dN\|_{\mathrm{tr}}\,|s|^{2}\leq H\,|s|^{2}-|\mathrm{tr}(q)-q(\nu,\nu)|\,|s|^{2}\leq-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}.$
On the other hand, using Lemma 3.6, we obtain
$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}\leq\|dN\|_{\mathrm{tr}}\,|s|^{2}.$
Putting these facts together, the assertion follows. ∎
Lemma 4.4.
Let $\sigma\in S_{0}$. Then $s\sigma$ is a $\nabla^{\mathcal{S},q}$-parallel section of $\mathcal{S}$. In particular, if $s\sigma$ vanishes at some point in $M$, then $s\sigma$ vanishes at each point in $M$.
Proof.
This follows directly from the fact that $\nabla^{\mathcal{E},q}$ is the tensor product connection of $\nabla^{\mathcal{S},q}$ and the flat connection on the trivial bundle over $M$ with fiber $S_{0}^{*}$. ∎
We define a linear subspace $L\subset S_{0}$ by
$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ at each point on $M$}\}.$
Lemma 4.5.
The subspace $L\subset S_{0}$ is invariant under $\gamma_{0}(N(p))$ for each point $p\in\Sigma$.
Proof.
Let us fix a point $p\in\Sigma$ and an element $\sigma\in L$. Using the boundary condition $\chi s=s$, we obtain
$\gamma(\nu)\circ s=s\circ\gamma_{0}(N)$
at the point $p$. Since $\sigma\in L$, it follows that
$0=\gamma(\nu)s\sigma=s\gamma_{0}(N)\sigma$
at the point $p$. Using Lemma 4.4, we conclude that $\gamma_{0}(N(p))\sigma\in L$. ∎
Lemma 4.6.
We have $L=\{0\}$. Moreover, the section $s$, considered as a homomorphism field, is invertible at each point in $M$.
Proof.
It follows from Lemma 4.5 that $L$ is invariant under $\gamma_{0}(N(p))$ for each point $p\in\Sigma$. Since the map $N\colon\Sigma\to S^{n-1}$ has non-zero degree, we know that the image of $N$ is all of $S^{n-1}$. Consequently, $L$ is invariant under $\gamma_{0}(v)$ for each $v\in\mathbb{R}^{n}$.
The spinor representation of the Clifford algebra $\mathbb{C}\mathrm{l}(\mathbb{R}^{n})\to\mathrm{End}(S_{0})$ is surjective. Therefore, the elements $\gamma_{0}(v)$, $v\in\mathbb{R}^{n}$, generate all of $\mathrm{End}(S_{0})$. Putting these facts together, we conclude that $L$ is invariant under $\mathrm{End}(S_{0})$. Therefore, $L=\{0\}$ or $L=S_{0}$. The latter case is impossible as this would imply that $s$ vanishes identically. To summarize, we have shown that $L=\{0\}$. Using Lemma 4.4, we conclude that $s$ is injective at each point in $M$. For dimensional reasons, $s$ is invertible at each point in $M$. ∎
Lemma 4.7.
Let $p$ be a point on the boundary $\Sigma$. Let $e_{1},\dots,e_{n-1}$ be an orthonormal basis of $T_{p}\Sigma$ and let $e^{0}_{1},\dots,e^{0}_{n-1}$ be an orthonormal basis of $T_{N(p)}S^{n-1}$ such that $dN(e_{j})=\lambda_{j}e^{0}_{j}$, where $\lambda_{j}\geq 0$ denote the singular values of $dN$. Then $h(e_{j})=\lambda_{j}e_{j}$ and $q(e_{j})^{\text{\rm tan}}=0$ for $j=1,\ldots,n-1$.
Proof.
By Lemma 4.3, we know that
$-\sum_{j=1}^{n-1}\lambda_{j}\,\langle\gamma(e_{j})\circ s\circ\gamma_{0}(e^{0}_{j}),s\rangle=\sum_{j=1}^{n-1}\lambda_{j}\,|s|^{2}.$
This implies $-\gamma(e_{j})\circ s\circ\gamma_{0}(e^{0}_{j})=s$ whenever $\lambda_{j}>0$. In other words, $\gamma(e_{j})\circ s=s\circ\gamma_{0}(e^{0}_{j})$ whenever $\lambda_{j}>0$. Applying Lemma 4.2 with $X=e_{j}$, we obtain
$\gamma(h(e_{j}))\circ s+i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\circ s=\lambda_{j}\,s\circ\gamma_{0}(e^{0}_{j})=\lambda_{j}\,\gamma(e_{j})\circ s$
for $j=1,\ldots,n-1$. By Lemma 4.6, $s$ is invertible at each point in $M$. Consequently,
$\gamma(h(e_{j}))+i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)=\lambda_{j}\,\gamma(e_{j})\in\text{\rm End}(\mathcal{S}|_{p})$
for $j=1,\ldots,n-1$. Note that $i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\in\text{\rm End}(\mathcal{S}|_{p})$ is self-adjoint, while $\gamma(h(e_{j}))\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(e_{j})\in\text{\rm End}(\mathcal{S}|_{p})$ are skew-adjoint. Thus, we conclude that $\gamma(h(e_{j}))=\lambda_{j}\,\gamma(e_{j})\in\text{\rm End}(\mathcal{S}|_{p})$ and $i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p})$ for $j=1,\ldots,n-1$. From this, the assertion follows. ∎
In summary, we proved all the assertions in Proposition 4.1 in odd dimensions.
4.2. Proof of Proposition 4.1 in even dimensions
We now turn to the even-dimensional case of Proposition 4.1. Suppose that $n\geq 2$ is an even integer. Since $\mu\geq|J|$ on $M$ and $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ along $\Sigma$, Corollary 3.12 yields a non-trivial $\nabla^{\mathcal{E},q}$-parallel section $s\in H^{1}(M,\mathcal{E})$. Clearly, $s$ is smooth. Moreover, we can find an element $\hat{t}\in S^{1}$ such that $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ along $\Sigma$. Here, the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$.
Lemma 4.8.
At each point $p\in\Sigma$, the second fundamental form $h$ and the tensor field $q$ satisfy the relation
$\gamma(h(X))\circ s+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\circ s=s\circ\gamma_{0}(d\hat{N}(X))$
for all $X\in T_{p}\Sigma$.
Proof.
Analogous to the proof of Lemma 4.2. ∎
Lemma 4.9.
We have
$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}=\|d\hat{N}\|_{\mathrm{tr}}\,|s|^{2}$
at each point on $\Sigma$. Moreover, $\|d\hat{N}\|_{\mathrm{tr}}=\|dN\|_{\mathrm{tr}}$ at each point on $\Sigma$.
Proof.
Let us fix a point $p\in\Sigma$. Let $e_{1},\dots,e_{n-1}$ be an orthonormal basis of $T_{p}\Sigma$. Using Lemma 4.8, we obtain
[…]
$=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}.$
This implies
$H\,|s|^{2}+i\,(\mathrm{tr}(q)-q(\nu,\nu))\,\langle\gamma(\nu)\circ s,s\rangle=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}.$
Consequently,
$\|dN\|_{\mathrm{tr}}\,|s|^{2}\leq H\,|s|^{2}-|\mathrm{tr}(q)-q(\nu,\nu)|\,|s|^{2}\leq-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}.$
As in Lemma 3.6, we can bound
$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}\leq\|d\hat{N}\|_{\mathrm{tr}}\,|s|^{2}.$
Finally, we note that $\|d\hat{N}\|_{\mathrm{tr}}\leq\|dN\|_{\mathrm{tr}}$ by definition of $\hat{N}$. Since $|s|^{2}>0$ at each point on $M$, the assertion follows. ∎
Lemma 4.10.
The map $\varphi(\cdot,\hat{t})\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n}$. In particular, the map $\hat{N}:\Sigma\to S^{n}$ is the composition of the map $N:\Sigma\to S^{n-1}$ with the standard inclusion $S^{n-1}\hookrightarrow S^{n}$.
Proof.
Suppose that the assertion is false. By Lemma 3.9, the map $\varphi(\cdot,\hat{t})$ is Lipschitz continuous with a Lipschitz constant strictly less than $1$. This implies $\|d\hat{N}\|_{\mathrm{tr}}<\|dN\|_{\mathrm{tr}}$ at each point on $\Sigma$ where $\|dN\|_{\mathrm{tr}}>0$. Using Lemma 4.9, we conclude that $\|dN\|_{\mathrm{tr}}=0$ at each point on $\Sigma$. This implies that $N$ is locally constant. This contradicts the assumption that $N$ has non-zero degree. ∎
Lemma 4.11.
Let $\sigma\in S_{0}$. Then $s\sigma$ is a $\nabla^{\mathcal{S},q}$-parallel section of $\mathcal{S}$. In particular, if $s\sigma$ vanishes at some point in $M$, then $s\sigma$ vanishes at each point in $M$.
Proof.
Analogous to the proof of Lemma 4.4. ∎
As above, we define a linear subspace $L\subset S_{0}$ by
$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ at each point on $M$}\}.$
Lemma 4.12.
The subspace $L\subset S_{0}$ is invariant under $\gamma_{0}(\hat{N}(p))$ for each point $p\in\Sigma$.
Proof.
Analogous to the proof of Lemma 4.5. ∎
Lemma 4.13.
We have $L=\{0\}$. Moreover, the section $s$, considered as a homomorphism field, is invertible on each point in $M$.
Proof.
Since the map $N\colon\Sigma\to S^{n-1}$ has non-zero degree, we know that the image of $N\colon\Sigma\to S^{n-1}$ is all of $S^{n-1}$. Using Lemma 4.10, we conclude that the image of the map $\hat{N}:\Sigma\to S^{n}$ is the equator in $S^{n}$. Thus, we conclude that $\text{\rm span}\{\hat{N}(p):p\in\Sigma\}=\mathbb{R}^{n}\subset\mathbb{R}^{n+1}$. By Lemma 4.12, $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n}$. Since the volume element acts as a scalar, we deduce that $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n+1}$. Arguing as in the proof of Lemma 4.6, the assertion follows. ∎
Lemma 4.14.
Let $p$ be a point on the boundary $\Sigma$. Let $e_{1},\dots,e_{n-1}$ be an orthonormal basis of $T_{p}\Sigma$ and let $e^{0}_{1},\dots,e^{0}_{n-1}$ be a set of orthonormal vectors in $T_{\hat{N}(p)}S^{n}$ such that $d\hat{N}(e_{j})=\lambda_{j}e^{0}_{j}$, where $\lambda_{j}\geq 0$ denote the singular values of $d\hat{N}$. Then $h(e_{j})=\lambda_{j}e_{j}$ and $q(e_{j})^{\text{\rm tan}}=0$ for $j=1,\ldots,n-1$.
Proof.
The proof is analogous to the proof of Lemma 4.7. ∎
In summary, we proved all the assertions in Proposition 4.1 in even dimensions.
4.3. Existence of a local isometric immersion into $\mathbb{R}^{n,1}$
In this subsection, we construct a local isometric immersion of $(M,g)$ into the Minkowski spacetime $\mathbb{R}^{n,1}$.
Definition 4.15.
Let $\mathcal{T}=TM\oplus\mathbb{R}$ denote the extended tangent bundle. We define a connection $\nabla^{\mathcal{T},q}$ on $\mathcal{T}$ by
$\nabla_{X}^{\mathcal{T},q}(Y,\varphi)=(D_{X}Y-\varphi\,q(X),X(\varphi)-q(X,Y))$
for every vector field $X$ and every section $(Y,\varphi)$ of $\mathcal{T}$. We define a bundle metric of signature $(n,1)$ on $\mathcal{T}$ by
$\langle(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}=g(Y,Z)-\varphi\psi$
for $(Y,\varphi),(Z,\psi)\in\mathcal{T}$.
Lemma 4.16.
The connection $\nabla^{\mathcal{T},q}$ is compatible with $\langle\cdot,\cdot\rangle_{\mathcal{T}}$.
Proof.
Let $X$ be a vector field, and let $(Y,\varphi)$, $(Z,\psi)$ be two sections of $\mathcal{T}$. Then
$\langle\nabla_{X}^{\mathcal{T},q}(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}+\langle(Y,\varphi),\nabla_{X}^{\mathcal{T},q}(Z,\psi)\rangle_{\mathcal{T}}$
$=g(D_{X}Y,Z)+g(Y,D_{X}Z)-X(\varphi)\,\psi-\varphi\,X(\psi)$
$=X(g(Y,Z)-\varphi\psi)$
$=X(\langle(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}).$
This completes the proof of Lemma 4.16. ∎
Lemma 4.17.
Suppose that $u,v$ are $\nabla^{\mathcal{S},q}$-parallel sections of $\mathcal{S}$. We define a vector field $Y$ on $M$ by the requirement $g(Y,Z)=\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)$ for all $Z\in TM$. Moreover, we define a scalar function $\varphi$ on $M$ by $\varphi=-\text{\rm Re}(\langle u,v\rangle)$. Then $(Y,\varphi)$ is a $\nabla^{\mathcal{T},q}$-parallel section of $\mathcal{T}$.
Proof.
By assumption,
$\nabla_{X}^{\mathcal{S}}u+\tfrac{i}{2}\,\gamma(q(X))\,u=\nabla_{X}^{\mathcal{S}}v+\tfrac{i}{2}\,\gamma(q(X))\,v=0.$
This implies
$g(D_{X}Y,Z)$ $=\text{\rm Re}(i\,\langle\gamma(Z)\,\nabla_{X}^{\mathcal{S}}u,v\rangle)+\text{\rm Re}(i\,\langle\gamma(Z)\,u,\nabla_{X}^{\mathcal{S}}v\rangle)$
$=\tfrac{1}{2}\,\text{\rm Re}(\langle\gamma(Z)\,\gamma(q(X))\,u,v\rangle)-\tfrac{1}{2}\,\text{\rm Re}(\langle\gamma(Z)\,u,\gamma(q(X))\,v\rangle)$
$=\tfrac{1}{2}\,\text{\rm Re}(\langle\gamma(Z)\,\gamma(q(X))\,u,v\rangle)+\tfrac{1}{2}\,\text{\rm Re}(\langle\gamma(q(X))\,\gamma(Z)\,u,v\rangle)$
$=-\text{\rm Re}(\langle u,v\rangle)\,q(X,Z)$
$=\varphi\,q(X,Z)$
and
$X(\varphi)$ $=-\text{\rm Re}(\langle\nabla_{X}^{\mathcal{S}}u,v\rangle)-\text{\rm Re}(\langle u,\nabla_{X}^{\mathcal{S}}v\rangle)$
$=\tfrac{1}{2}\,\text{\rm Re}(i\,\langle\gamma(q(X))\,u,v\rangle)-\tfrac{1}{2}\,\text{\rm Re}(i\,\langle u,\gamma(q(X))\,v\rangle)$
$=\text{\rm Re}(i\,\langle\gamma(q(X))\,u,v\rangle)$
$=g(Y,q(X))$
$=q(X,Y).$
This completes the proof of Lemma 4.17. ∎
Lemma 4.18.
Let $p$ be an arbitrary point in $M$. Consider the set of all elements $(Y,\varphi)\in\mathcal{T}|_{p}$ which can be written in the form $g(Y,Z)=\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)$ and $\varphi=-\text{\rm Re}(\langle u,v\rangle)$, where $u,v\in\mathcal{S}|_{p}$. This set spans $\mathcal{T}|_{p}$.
Proof.
Suppose that the assertion is false. Then there exists a non-zero element $(Z,\psi)\in\mathcal{T}|_{p}$ such that
$\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)+\text{\rm Re}(\psi\,\langle u,v\rangle)=0$
for all $u,v\in\mathcal{S}|_{p}$. From this, we deduce that $i\,\gamma(Z)+\psi\,\text{\rm id}=0\in\text{\rm End}(\mathcal{S}|_{p})$. This implies $Z=0$ and $\psi=0$. This is a contradiction. ∎
Proposition 4.19.
The bundle $\mathcal{T}$ admits a trivialization by $\nabla^{\mathcal{T},q}$-parallel sections.
Proof.
By Proposition 4.1, there exists a $\nabla^{\mathcal{E},q}$-parallel section $s$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $s$ is invertible at each point in $M$. Consequently, the spinor bundle $\mathcal{S}$ admits a trivialization by $\nabla^{\mathcal{S},q}$-parallel sections. Using Lemma 4.17, we obtain a collection of $\nabla^{\mathcal{T},q}$-parallel sections of $\mathcal{T}$. At each point $p\in M$, these $\nabla^{\mathcal{T},q}$-parallel sections span the fiber $\mathcal{T}|_{p}$ (see Lemma 4.18). This completes the proof of Proposition 4.19. ∎
Let $(E^{(0)},\zeta^{(0)}),(E^{(1)},\zeta^{(0)}),\ldots,(E^{(n)},\zeta^{(n)})$ be a collection of $\nabla^{\mathcal{T},q}$-parallel sections that trivialize $\mathcal{T}$. In view of Lemma 4.16, we may write
$\langle(E^{(\alpha)},\zeta^{(\alpha)}),(E^{(\beta)},\zeta^{(\beta)})\rangle_{\mathcal{T}}=m_{\alpha\beta},$
where $m_{\alpha\beta}$ is a constant $(n+1)\times(n+1)$-matrix. The matrix $m_{\alpha\beta}$ is invertible with signature $(n,1)$. By taking suitable linear combinations, we may arrange that
$m_{\alpha\beta}=\begin{cases}-1&\text{\rm for $\alpha=\beta=0$}\\ 1&\text{\rm for $\alpha=\beta>0$}\\ 0&\text{\rm for $\alpha\neq\beta$}\end{cases}$
for $\alpha,\beta=0,1,\ldots,n$. Then
$\langle(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,\langle(E^{(\alpha)},\zeta^{(\alpha)}),(Y,\varphi)\rangle_{\mathcal{T}}\,\langle(E^{(\beta)},\zeta^{(\beta)}),(Z,\psi)\rangle_{\mathcal{T}}$
for all sections $(Y,\varphi)$ and $(Z,\psi)$ of $\mathcal{T}$. Choosing $\varphi=\psi=0$, this implies
(4.3) $g(Y,Z)=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,g(E^{(\alpha)},Y)\,g(E^{(\beta)},Z)$
for all $Y,Z\in TM$. Moreover, choosing $\varphi=0$ and $Z=0$, we find
(4.4) $00=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,g(E^{(\alpha)},Y)\,\zeta^{(\beta)}=$ 0
for all $Y\in TM$. Finally, choosing $Y=Z=0$ and $\varphi=\psi=1$ yields
(4.5) $-1=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,\zeta^{(\alpha)}\,\zeta^{(\beta)}.$
Since $(E^{(\alpha)},\zeta^{(\alpha)})$ is $\nabla^{\mathcal{T},q}$-parallel, it follows that the covariant derivative satisfies $DE^{(\alpha)}=\zeta^{(\alpha)}q$ and is hence symmetric. In particular, the one-form $g(E^{(\alpha)},\cdot)$ is closed. We may locally write $g(E^{(\alpha)},\cdot)=df^{(\alpha)}$ for some scalar function $f^{(\alpha)}$. It follows from (4.3) that $g=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,df^{(\alpha)}\otimes df^{(\beta)}$. In other words, the map $F=(f^{(0)},f^{(1)},\ldots,f^{(n)})$ is a local isometric immersion of $(M,g)$ into the Minkowski spacetime $\mathbb{R}^{n,1}$. For each vector $X\in TM$, we have
$dF(X)$ $=(df^{(0)}(X),df^{(1)}(X),\ldots,df^{(n)}(X))$
$=(g(E^{(0)},X),g(E^{(1)},X),\ldots,g(E^{(n)},X))\in\mathbb{R}^{n,1}.$
It follows from (4.4) and (4.5) that the unit normal vector to the immersion $F$ is given by $(\zeta^{(0)},\zeta^{(1)},\ldots,\zeta^{(n)})\in\mathbb{R}^{n,1}$. Using the fact that $(E^{(\alpha)},\zeta^{(\alpha)})$ is $\nabla^{\mathcal{T},q}$-parallel, it is easy to see that the second fundamental form associated with $F$ is equal to $q$.
Note that the isometric immersion $F=(f^{(0)},f^{(1)},\ldots,f^{(n)})$ is indeed globally well-defined if $M$ is simply connected. If $M$ is not simply connected, then $F$ may not be globally well-defined, but its differential is globally well-defined.
4.4. Proof of Theorem I in the simply connected case
In this subsection, we prove Theorem I in the special case that $M$ is simply connected. In this case, there exists an isometric immersion $F$ of $M$ into the Minkowski spacetime $\mathbb{R}^{n,1}$ as a spacelike hypersurface. Let $T$ denote the normal vector to $M$ in $\mathbb{R}^{n,1}$, normalized so that $\langle T,T\rangle_{\mathbb{R}^{n,1}}=-1$. The second fundamental form is equal to $q$, so that $q(X,Y)=\langle\bar{D}_{X}T,Y\rangle_{\mathbb{R}^{n,1}}$ for $X,Y\in TM$. Let $\mathcal{S}$ denote the spinor bundle of $M$, and let $\nabla^{\mathcal{S}}$ denote the usual connection on $\mathcal{S}$.
Definition 4.20.
We define $\mathcal{W}=\mathcal{S}\oplus\mathcal{S}$. We define a connection $\nabla^{\mathcal{W}}$ on $\mathcal{W}$ by
$\nabla_{X}^{\mathcal{W}}(u,v)=(\nabla_{X}^{\mathcal{S},q}u,\nabla_{X}^{\mathcal{S},-q}v),$
where $X$ denotes a vector field on $M$, and $u,v$ are sections of $\mathcal{S}$.
Definition 4.21.
For each point $p\in M$, we define a map $\gamma^{\mathcal{W}}:\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ by
$\gamma^{\mathcal{W}}(T)\,(u,v)=(-iv,iu)$
and
$\gamma^{\mathcal{W}}(X)\,(u,v)=(\gamma(X)\,v,\gamma(X)\,u),$
where $X\in TM$ and $u,v\in\mathcal{S}$.
Lemma 4.22.
The map $\gamma^{\mathcal{W}}:\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ satisfies the Clifford relations with respect to the Lorentzian inner product on $\mathbb{R}^{n,1}$. In other words,
$\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)+\gamma^{\mathcal{W}}(b)\,\gamma^{\mathcal{W}}(a)=-2\,\langle a,b\rangle_{\mathbb{R}^{n,1}}\,\text{\rm id}$
for all $a,b\in\mathbb{R}^{n,1}$.
Proof.
Note that
$\gamma^{\mathcal{W}}(X)\,\gamma^{\mathcal{W}}(Y)+\gamma^{\mathcal{W}}(Y)\,\gamma^{\mathcal{W}}(X)=-2\,g(X,Y)\,\text{\rm id}$
for all $X,Y\in TM$. Moreover,
$\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(X)+\gamma^{\mathcal{W}}(X)\,\gamma^{\mathcal{W}}(T)=0$
for all $X\in TM$ and
$(\gamma^{\mathcal{W}}(T))^{2}=\text{\rm id}.$
This completes the proof of Lemma 4.22. ∎
Lemma 4.23.
Suppose that $a$ is a constant vector in $\mathbb{R}^{n,1}$. Then
$\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(a)\,(u,v))=\gamma^{\mathcal{W}}(a)\,\nabla_{X}^{\mathcal{W}}(u,v)$
for all $X\in TM$. In particular, $\gamma^{\mathcal{W}}(a)$ maps parallel sections of $\mathcal{W}$ to parallel sections of $\mathcal{W}$.
Proof.
We decompose $a=Y+fT$, where $Y$ is a section of $TM$ and $f$ is a scalar function on $M$. Then $\bar{D}_{X}Y+\bar{D}_{X}(fT)=0$, where $\bar{D}$ denotes the standard flat connection on $\mathbb{R}^{n,1}$. Using the formula $\bar{D}_{X}Y=D_{X}Y+q(X,Y)\,T$, we obtain
$\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(Y)\,(u,v))-\gamma^{\mathcal{W}}(Y)\,\nabla_{X}^{\mathcal{W}}(u,v)-\gamma^{\mathcal{W}}(\bar{D}_{X}Y)\,(u,v)$
$=\nabla_{X}^{\mathcal{W}}(\gamma(Y)\,v,\gamma(Y)\,u)-\gamma^{\mathcal{W}}(Y)\,(\nabla_{X}^{\mathcal{S},q}u,\nabla_{X}^{\mathcal{S},-q}v)$
$-\gamma^{\mathcal{W}}(D_{X}Y)\,(u,v)-q(X,Y)\,\gamma^{\mathcal{W}}(T)\,(u,v)$
(4.6) $=(\nabla_{X}^{\mathcal{S},q}(\gamma(Y)\,v),\nabla_{X}^{\mathcal{S},-q}(\gamma(Y)\,u))-(\gamma(Y)\,\nabla_{X}^{\mathcal{S},-q}v,\gamma(Y)\,\nabla_{X}^{\mathcal{S},q}u)$
$-(\gamma(D_{X}Y)\,v,\gamma(D_{X}Y)\,u)-q(X,Y)\,(-iv,iu)$
$=0$
for all $X,Y\in TM$. Using the formula $\bar{D}_{X}T=q(X)$, we obtain
$\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(T)\,(u,v))-\gamma^{\mathcal{W}}(T)\,\nabla_{X}^{\mathcal{W}}(u,v)-\gamma^{\mathcal{W}}(\bar{D}_{X}T)\,(u,v)$
$=\nabla_{X}^{\mathcal{W}}(-iv,iu)-\gamma^{\mathcal{W}}(T)\,(\nabla_{X}^{\mathcal{S},q}u,\nabla_{X}^{\mathcal{S},-q}v)-\gamma^{\mathcal{W}}(q(X))\,(u,v)$
$=(-i\,\nabla_{X}^{\mathcal{S},q}v,i\,\nabla_{X}^{\mathcal{S},-q}u)-(-i\,\nabla_{X}^{\mathcal{S},-q}v,i\,\nabla_{X}^{\mathcal{S},q}u)-(\gamma(q(X))\,v,\gamma(q(X))\,u)$
$=0$
for all $X\in TM$. This implies
(4.7) $\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(fT)\,(u,v))-\gamma^{\mathcal{W}}(fT)\,\nabla_{X}^{\mathcal{W}}(u,v)-\gamma^{\mathcal{W}}(\bar{D}_{X}(fT))\,(u,v)=0$
for all $X\in TM$. In the next step, we add (4.4) and (4.7). Using the identity $\bar{D}_{X}Y+\bar{D}_{X}(fT)=0$, we conclude that
$\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(a)\,(u,v))-\gamma^{\mathcal{W}}(a)\,\nabla_{X}^{\mathcal{W}}(u,v)=0$
for all $X\in TM$. This completes the proof of Lemma 4.23. ∎
Definition 4.24.
For each point $p\in M$, we define a map $\gamma^{\mathcal{W}}:\Lambda^{2}\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ by
$\gamma^{\mathcal{W}}(a\wedge b)=\frac{1}{2}\,(\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)-\gamma^{\mathcal{W}}(b)\,\gamma^{\mathcal{W}}(a)).$
Note that $\gamma^{\mathcal{W}}(a\wedge b)=\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)$ whenever $\langle a,b\rangle_{\mathbb{R}^{n,1}}=0$.
Lemma 4.25.
Suppose that $p$ is a point in $M$ and $\omega\in\Lambda^{2}\mathbb{R}^{n,1}$ is a two-form with the property that $\gamma^{\mathcal{W}}(\omega)=0\in\text{\rm End}(\mathcal{W}|_{p})$. Then $\omega=0$.
Proof.
We decompose $\omega=\mu+T\wedge\eta$, where $\mu\in\Lambda^{2}T_{p}M$ and $\eta\in T_{p}M$. By assumption,
00 0 $=\gamma^{\mathcal{W}}(\omega)\,(u,v)$
$=\gamma^{\mathcal{W}}(\mu)\,(u,v)+\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\eta)\,(u,v)$
$=(\gamma(\mu)\,u-i\,\gamma(\eta)\,u,\gamma(\mu)\,v+i\,\gamma(\eta)\,v)$
for all $u,v\in\mathcal{S}|_{p}$. Therefore, $\gamma(\mu)-i\,\gamma(\eta)=0\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(\mu)+i\,\gamma(\eta)=0\in\text{\rm End}(\mathcal{S}|_{p})$. This implies $\gamma(\mu)=0\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(\eta)=0\in\text{\rm End}(\mathcal{S}|_{p})$. Thus, $\mu=0$ and $\eta=0$. This completes the proof of Lemma 4.25. ∎
By Proposition 4.1, there exists a $\nabla^{\mathcal{E},q}$-parallel section $s$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $s$ is invertible at each point in $M$. Moreover, $s$ satisfies the boundary condition $\gamma(\nu)\circ s=s\circ\gamma_{0}(N)$ at each point on the boundary $\Sigma$.
Replacing $q$ by $-q$, we can find a $\nabla^{\mathcal{E},-q}$-parallel section $t$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $t$ is invertible at each point in $M$. Moreover, $t$ satisfies the boundary condition $\gamma(\nu)\circ t=t\circ\gamma_{0}(N)$ at each point on the boundary $\Sigma$.
We denote by $\mathcal{Y}$ the set of all pairs $(v,\omega)\in\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}$ with the property that
$i\,\gamma^{\mathcal{W}}(\omega)\circ(s,t)=(s\circ\gamma_{0}(v),-t\circ\gamma_{0}(v))$
at each point on $M$. Clearly, $\mathcal{Y}$ is a linear subspace of $\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}$. Let $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ denote the image of $\mathcal{Y}$ under the projection $\mathcal{Y}\to\Lambda^{2}\mathbb{R}^{n,1},\,(v,\omega)\mapsto\omega$.
Proposition 4.26.
We have $\mathcal{Y}=\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p)):p\in\Sigma\}$. The projection $\mathcal{Y}\to\mathbb{R}^{n},\,(v,\omega)\mapsto v$ is bijective. In particular, $\mathcal{Y}$ has dimension $n$.
Proof.
Using the boundary conditions satisfied by $s$ and $t$, we compute
$i\,\gamma^{\mathcal{W}}(T\wedge\nu)\circ(s,t)$ $=i\,\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\nu)\circ(s,t)$
$=(\gamma(\nu)\circ s,-\gamma(\nu)\circ t)$
$=(s\circ\gamma_{0}(N),-t\circ\gamma_{0}(N))$
at each point on $\Sigma$. Let us fix a boundary point $p\in\Sigma$, and let $v=N(p)\in\mathbb{R}^{n}$, $a=T(p)\in\mathbb{R}^{n,1}$, and $b=\nu(p)\in\mathbb{R}^{n,1}$. Then
$i\,\gamma^{\mathcal{W}}(a\wedge b)\circ(s,t)=(s\circ\gamma_{0}(v),-t\circ\gamma_{0}(v))$
at the point $p$. Since the left hand side and the right hand side are parallel sections with respect to the same connection, we conclude that
$i\,\gamma^{\mathcal{W}}(a\wedge b)\circ(s,t)=(s\circ\gamma_{0}(v),-t\circ\gamma_{0}(v))$
at each point in $M$. Thus, $(v,a\wedge b)\in\mathcal{Y}$. To summarize, we have shown that
$\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p)):p\in\Sigma\}\subset\mathcal{Y}.$
Since the map $N:\Sigma\to S^{n-1}$ is surjective, it follows that the projection $\mathcal{Y}\to\mathbb{R}^{n},\,(v,\omega)\mapsto v$ is surjective. On the other hand, it follows from Lemma 4.25 that the projection $\mathcal{Y}\to\mathbb{R}^{n},\,(v,\omega)\mapsto v$ is injective. Putting these facts together, we conclude that
$\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p)):p\in\Sigma\}=\mathcal{Y}.$
This completes the proof of Proposition 4.26. ∎
Corollary 4.27.
We have $\mathcal{Z}=\{\mathbb{R}\cdot(T(p)\wedge\nu(p)):p\in\Sigma\}$. Moreover, $\mathcal{Z}$ has dimension $n$.
Proof.
The first statement follows directly from Proposition 4.26. The second statement follows from the fact that $\mathcal{Y}$ has dimension $n$ and the projection $\mathcal{Y}\to\Lambda^{2}\mathbb{R}^{n,1},\,(v,\omega)\mapsto\omega$ is injective. ∎
Proposition 4.28.
There exists a basis $\{b_{0},b_{1},\ldots,b_{n}\}$ of $\mathbb{R}^{n,1}$ such that $b_{0}$ is timelike and $\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{0}\wedge b_{2},\ldots,b_{0}\wedge b_{n}\}$.
Proof.
By Theorem 6 in [LA], there are two possibilities:
Case 1: We have $n=3$ and we can find three linearly independent vectors $b_{0},b_{1},b_{2}\in\mathbb{R}^{3,1}$ with the property that $\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{1}\wedge b_{2},b_{2}\wedge b_{0}\}$. We can find two linearly independent vectors $\tilde{b}_{1},\tilde{b}_{2}\in\text{\rm span}\{b_{0},b_{1},b_{2}\}$ such that $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}$ is spacelike. Then $\tilde{b}_{1}\wedge\tilde{b}_{2}\in\mathcal{Z}$. By Corollary 4.27, we can find a point $p\in\Sigma$ with the property that $\tilde{b}_{1}\wedge\tilde{b}_{2}$ is a non-zero scalar multiple of $T(p)\wedge\nu(p)$. In particular, $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}=\text{\rm span}\{T(p),\nu(p)\}$. This contradicts the fact that $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}$ does not contain a timelike vector.
Case 2: There exists a basis $\{b_{0},b_{1},\ldots,b_{n}\}$ of $\mathbb{R}^{n,1}$ with the property that $\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{0}\wedge b_{2},\ldots,b_{0}\wedge b_{n}\}$. We claim that $b_{0}$ is timelike. Suppose that $b_{0}$ is spacelike or null. We can find a vector $\tilde{b}\in\mathbb{R}^{n,1}$ such that $\tilde{b}$ is orthogonal to $b_{0}$ and $\tilde{b}$ is spacelike. This implies that $b_{0}$ and $\tilde{b}$ are linearly independent. Moreover, $b_{0}\wedge\tilde{b}\in\mathcal{Z}$. By Corollary 4.27, we can find a point $p\in\Sigma$ with the property that $b_{0}\wedge\tilde{b}$ is a non-zero scalar multiple of $T(p)\wedge\nu(p)$. In particular, $\text{\rm span}\{b_{0},\tilde{b}\}=\text{\rm span}\{T(p),\nu(p)\}$. This contradicts the fact that $\text{\rm span}\{b_{0},\tilde{b}\}$ does not contain a timelike vector. ∎
By scaling, we may assume that $\langle b_{0},b_{0}\rangle_{\mathbb{R}^{n,1}}=-1$. By applying a Lorentz transformation, we may assume that $b_{0}=(1,0,\ldots,0)\in\mathbb{R}^{n,1}$.
For each point $p\in\Sigma$, we know that $T(p)\wedge\nu(p)\in\text{\rm span}\{b_{0}\wedge b_{1},b_{0}\wedge b_{2},\ldots,b_{0}\wedge b_{n}\}$, hence $b_{0}\in\text{\rm span}\{T(p),\nu(p)\}$. Consequently, $\langle b_{0},X\rangle_{\mathbb{R}^{n,1}}=0$ for every tangent vector $X\in T\Sigma$. In particular, each connected component of $\Sigma$ is contained in a hyperplane in $\mathbb{R}^{n,1}$ which is orthogonal to $b_{0}$.
For each point $p\in\Sigma$, we can find a spacelike unit vector $B(p)\in\mathbb{R}^{n,1}$ such that $B(p)$ is orthogonal to $b_{0}$ and $T(p)\wedge\nu(p)=b_{0}\wedge B(p)$. This defines a smooth map $B:\Sigma\to S^{n-1}\subset\mathbb{R}^{n,1}$. The map $B$ can be interpreted as the unit normal to $\Sigma$ within the hyperplane orthogonal to $b_{0}$.
Using Proposition 4.26, we obtain $\mathcal{Y}=\{\mathbb{R}\cdot(N(p),b_{0}\wedge B(p)):p\in\Sigma\}$. This implies that the map $B:\Sigma\to S^{n-1}\subset\mathbb{R}^{n,1}$ is the composition of the map $N:\Sigma\to S^{n-1}$ with a linear isometry of $\mathbb{R}^{n}$. In particular, $\|dB\|_{\text{\rm tr}}=\|dN\|_{\text{\rm tr}}$ at each point on $\Sigma$.
Proposition 4.29.
If $p$ is a point on the boundary $\Sigma$ with $\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\neq 0$, then $h=0$, $dN=0$, and $dB=0$ at the point $p$.
Proof.
Let us fix a point $p\in\Sigma$, and let $\{e_{1},\ldots,e_{n-1}\}$ denote an orthonormal basis of $T_{p}\Sigma$. Differentiating the identity $T\wedge\nu=b_{0}\wedge B$ gives
$\sum_{j=1}^{n-1}h(e_{i},e_{j})\,T\wedge e_{j}+\sum_{j=1}^{n-1}q(e_{i},e_{j})\,e_{j}\wedge\nu=b_{0}\wedge dB(e_{i})$
for each $i=1,\ldots,n-1$. By Proposition 4.1, we know that $q(e_{i},e_{j})=0$ for $i,j=1,\ldots,n-1$. This gives
00 0 $=\sum_{j=1}^{n-1}h(e_{i},e_{j})\,\langle T\wedge e_{j},\nu\wedge e_{k}\rangle_{\mathbb{R}^{n,1}}$
$=\langle b_{0}\wedge dB(e_{i}),\nu\wedge e_{k}\rangle_{\mathbb{R}^{n,1}}$
$=\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\,\langle dB(e_{i}),e_{k}\rangle_{\mathbb{R}^{n,1}}.$
Hence, if $\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\neq 0$ at the point $p$, then $dB=0$ at the point $p$. This implies $dN=0$ at the point $p$. Using Proposition 4.1, it follows that $h=0$ at the point $p$. ∎
Proposition 4.30.
The composition of the isometric immersion $F\colon(M,g)\to\mathbb{R}^{n,1}$ with the projection $\pi\colon\mathbb{R}^{n,1}\to\mathbb{R}^{n}$, $(t,x)\mapsto x$, is a diffeomorphism onto a compact convex domain $K\subset\mathbb{R}^{n}$ with smooth boundary. In particular, $F$ is an isometric embedding onto the graph of a smooth real valued function defined on $K$.
Proof.
We define a Riemannian metric $g_{0}$ on $M$ by
$g_{0}(X,Y)=g(X,Y)+\langle b_{0},X\rangle_{\mathbb{R}^{n,1}}\,\langle b_{0},Y\rangle_{\mathbb{R}^{n,1}}$
for $X,Y\in TM$. In other words, $g$ is the pull-back of the Euclidean metric under the immersion $\pi\circ F\colon M\to\mathbb{R}^{n}$. In particular, the metric $g_{0}$ is flat. It follows from Proposition 4.29 that the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ coincides with the second fundamental form of $\partial M$ with respect to the metric $g$. Using Proposition 4.1, we conclude that the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ is positive semi-definite at each point on $\partial M$. Moreover, since $N$ has non-zero degree, we can find a point on $\partial M$ where the Jacobian determinant of $N$ is non-zero, and at that point the second fundamental of $\partial M$ with respect to $g_{0}$ is positive definite. Theorem A.1 now implies that there exists an isometry of $(M,g_{0})$ onto a compact convex domain in $\mathbb{R}^{n}$ with smooth boundary. The isometric immersion $\pi\circ F\colon(M,g_{0})\to\mathbb{R}^{n}$ coincides with this isometry up to a rigid motion of $\mathbb{R}^{n}$. This implies the assertion. ∎
This completes the proof of Theorem I in the special case when $M$ is simply connected.
4.5. Proof of Theorem I in the general case
In this final subsection, we treat the case when $M$ is not simply connected. In this case, the isometric immersion $F:M\to\mathbb{R}^{n,1}$ is only locally defined, but its differential its globally defined. This is sufficient to define the bundle $\mathcal{W}$, the connection $\nabla^{\mathcal{W}}$, and the map $\gamma^{\mathcal{W}}:\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$. The arguments in Subsection 4.4 now imply that there exists a vector $b_{0}\in\mathbb{R}^{n,1}$ such that $\langle b_{0},b_{0}\rangle_{\mathbb{R}^{n,1}}=-1$ and $\langle b_{0},dF(X)\rangle_{\mathbb{R}^{n,1}}=0$ for every tangent vector $X\in T\Sigma$.
We now define a Riemannian metric $g_{0}$ on $M$ by
$g_{0}(X,Y)=g(X,Y)+\langle b_{0},dF(X)\rangle_{\mathbb{R}^{n,1}}\,\langle b_{0},dF(Y)\rangle_{\mathbb{R}^{n,1}}$
for $X,Y\in TM$. Note that $g_{0}$ is a globally defined metric on $M$. As in Section 4.4, we see that $g_{0}$ is flat. Moreover, the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ coincides with the second fundamental form of $\partial M$ with respect to the metric $g$. In particular, the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ is positive semi-definite at each point on $\partial M$, and we can find a point on $\partial M$ where the second fundamental of $\partial M$ with respect to $g_{0}$ is positive definite. Using Theorem A.1, it follows that $M$ is diffeomorphic to a ball. This contradicts the assumption that $M$ is not simply connected.
5. Approximating a compact, convex polytope by smooth domains
Throughout this section, we fix an integer $n\geq 2$. We further assume that $\Omega\subset\mathbb{R}^{n}$ is a compact, convex polytope with non-empty interior. We write $\Omega=\bigcap_{i\in I}\{u_{i}\leq 0\}$, where $I$ is a finite set and the $u_{i}$ are linear functions on $\mathbb{R}^{n}$. After eliminating redundant inequalities, we may assume that the following condition is satisfied.
Assumption 5.1.
For each $i_{0}\in I$, the set
$\{u_{i_{0}}>0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}\leq 0\}$
is non-empty.
Let $g$ be a Riemannian metric which is defined on an open set containing $\Omega$, and let $q$ be a symmetric $(0,2)$-tensor. For each $i\in I$, $\nabla u_{i}$ will denote the gradient of $u_{i}$ with respect to the metric $g$, $D^{2}u_{i}$ will denote the Hessian of $u_{i}$ with respect to the metric $g$, $|\nabla u_{i}|_{g}$ will denote the norm of the gradient of $u_{i}$ with respect to the metric $g$, and $\nu_{i}=\frac{\nabla u_{i}}{|\nabla u_{i}|_{g}}$ will denote the unit normal vector field, with respect to the metric $g$, to the level sets of $u_{i}$. For each $i\in I$, we denote by $N_{i}\in S^{n-1}$ the outward-pointing unit normal vector to the halfspace $\{u_{i}\leq 0\}$ with respect to the Euclidean metric.
For each $\lambda>0$, we define
$\Omega_{\lambda}=\bigg{\{}\sum_{i\in I}e^{\lambda u_{i}}\leq 1\bigg{\}}\subset\Omega.$
If $\lambda$ is sufficiently large, then $\Omega_{\lambda}$ is a compact, convex domain in $\mathbb{R}^{n}$ with smooth boundary $\Sigma_{\lambda}=\partial\Omega_{\lambda}$. The sets $\Omega_{\lambda}$ form an increasing family of sets. Moreover,
$\bigcup_{\lambda>\lambda_{0}}\Omega_{\lambda}=\bigcap_{i\in I}\{u_{i}<0\}.$
In the remainder of this section, $|\cdot|_{g}$ will denote the norm taken with respect to the metric $g$, and ||$\cdot$|| ⋅ | will denote the norm taken with respect to the Euclidean metric. The outward-pointing unit normal vector to the domain $\Omega_{\lambda}$ with respect to the metric $g$ is given by
$\nu=\frac{\sum_{i\in I}e^{\lambda u_{i}}\,\nabla u_{i}}{\big{|}\sum_{i\in I}e^{\lambda u_{i}}\,\nabla u_{i}\big{|}_{g}}=\frac{\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}\,\nu_{i}}{\big{|}\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}\,\nu_{i}\big{|}_{g}}.$
We define a map $N\colon\Sigma_{\lambda}\to S^{n-1}$ by
(5.1) $N=\frac{\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}\,N_{i}}{\big{|}\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}\,N_{i}\big{|}}.$
The map $N\colon\Sigma_{\lambda}\to S^{n-1}$ is homotopic to the Gauss map of $\Sigma_{\lambda}$ with respect to the Euclidean metric. A homotopy is induced by deforming $g$ to the Euclidean metric. In particular, the map $N$ has degree $1$.
Proposition 5.1 (Prop. 3.9 in [Brendle]).
Let $x\in\Sigma_{\lambda}$. Let $\pi\colon T_{x}\Omega\to T_{x}\Omega$ denote the orthogonal projection to the orthogonal complement of $\nu$, and let $P\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ denote the orthogonal projection to the orthogonal complement of $N$. Then $H-\|dN\|_{\mathrm{tr}}\geq V_{\lambda}$, where the function $V_{\lambda}\colon\Sigma_{\lambda}\to\mathbb{R}$ is defined by
$V_{\lambda}$ $=\lambda\,\frac{\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}^{2}\,|\pi(\nu_{i})|_{g}^{2}}{\big{|}\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}\,\nu_{i}\big{|}_{g}}-\lambda\,\frac{\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}^{2}\,|\pi(\nu_{i})|_{g}\,|P(N_{i})|}{\big{|}\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}\,N_{i}\big{|}}$
$\quad+\frac{\sum_{i\in I}e^{\lambda u_{i}}\,(\Delta u_{i}-(D^{2}u_{i})(\nu,\nu))}{\big{|}\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}\,\nu_{i}\big{|}_{g}}-\frac{\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla(|\nabla u_{i}|_{g})|_{g}\,|P(N_{i})|}{\big{|}\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}\,N_{i}\big{|}}.$
We define a function $W_{\lambda}\colon\Sigma_{\lambda}\to\mathbb{R}$ by
$W_{\lambda}=V_{\lambda}-|\mathrm{tr}(q)-q(\nu,\nu)|$
Proposition 5.2.
Suppose that the Matching Angle Hypothesis is satisfied. Then $\sup_{\Sigma_{\lambda}}\max\{-W_{\lambda},0\}\leq o(\lambda)$ as $\lambda\to\infty$.
Proof.
It follows from Proposition 3.10 in [Brendle] that $\sup_{\Sigma_{\lambda}}\max\{-V_{\lambda},0\}\leq o(\lambda)$ as $\lambda\to\infty$. Since $W_{\lambda}-V_{\lambda}$ is uniformly bounded, the assertion follows. ∎
In the remainder of this section, we will estimate the $L^{\sigma}$-norm $\max\{-W_{\lambda},0\}$ on $\Sigma_{\lambda}\cap B_{r}(p)$, where $\sigma\in[1,\frac{3}{2})$ is a fixed exponent and $B_{r}(p)$ denotes a Euclidean ball of radius $r$. We begin by recalling a basic fact about the area of convex hypersurfaces in $\mathbb{R}^{n}$.
Definition 5.3.
Consider three pairwise distinct elements $i_{1},i_{2},i_{3}\in I$. We denote by $G_{\lambda}^{(i_{1},i_{2},i_{3})}$ the set of all points $x\in\Sigma_{\lambda}$ with the property that $u_{i_{1}}(x)\geq u_{i_{2}}(x)\geq u_{i_{3}}(x)$ and $u_{i_{3}}(x)\geq u_{i}(x)$ for $i\in I\setminus\{i_{1},i_{2},i_{3}\}$.
Lemma 5.4.
For each $i\in I$, we assume that the hypersurface $\{u_{i}=0\}$ satisfies the condition $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$. Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
$\bigg{(}r^{\sigma+1-n}\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}$
for all pairwise distinct elements $i_{1},i_{2},i_{3}\in I$.
Proof.
We adapt the proof of Lemma 3.13 in [Brendle]. Let us consider an arbitrary point $x\in G_{\lambda}^{(i_{1},i_{2},i_{3})}$ with $u_{i_{2}}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$. By definition of $G_{\lambda}^{(i_{1},i_{2},i_{3})}$, it follows that $u_{i}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$ for all $i\in I\setminus\{i_{1}\}$. Using the identity $\sum_{i\in I}e^{\lambda u_{i}(x)}=1$, we obtain $e^{\lambda u_{i_{1}}(x)}\geq 1-C\,e^{-(\lambda r)^{\frac{1}{8}}}$, hence $u_{i_{1}}(x)\geq-C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$. Arguing as in [Brendle], we obtain
$V_{\lambda}\geq\frac{\Delta u_{i_{1}}-(D^{2}u_{i_{1}})(\nu_{i_{1}},\nu_{i_{1}})}{|\nabla u_{i_{1}}|_{g}}-C\lambda\,e^{-(\lambda r)^{\frac{1}{8}}}$
at the point $x$. Moreover, $|\nu-\nu_{i_{1}}|_{g}\leq C\,e^{-(\lambda r)^{\frac{1}{8}}}$ at the point $x$. This implies
$|\mathrm{tr}(q)-q(\nu,\nu)|\leq|\mathrm{tr}(q)-q(\nu_{i_{1}},\nu_{i_{1}})|+C\,e^{-(\lambda r)^{\frac{1}{8}}}$
at the point $x$. Putting these facts together, we obtain
$W_{\lambda}$ $=V_{\lambda}-|\mathrm{tr}(q)-q(\nu,\nu)|$
$\geq\frac{\Delta u_{i_{1}}-(D^{2}u_{i_{1}})(\nu_{i_{1}},\nu_{i_{1}})}{|\nabla u_{i_{1}}|_{g}}-|\mathrm{tr}(q)-q(\nu_{i_{1}},\nu_{i_{1}})|-C\lambda\,e^{-(\lambda r)^{\frac{1}{8}}}$
at the point $x$. Since $u_{i_{1}}(x)\geq-C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$ and $u_{i}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$ for all $i\in I\setminus\{i_{1}\}$, we can find a point $y\in\Omega$ such that $u_{i_{1}}(y)=0$ and $d(x,y)\leq C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$. By assumption, $H-|\mathrm{tr}(q)-q(\nu_{i_{1}},\nu_{i_{1}})|\geq 0$ at the point $y$. This implies
$\frac{\Delta u_{i_{1}}-(D^{2}u_{i_{1}})(\nu_{i_{1}},\nu_{i_{1}})}{|\nabla u_{i_{1}}|_{g}}-|\mathrm{tr}(q)-q(\nu_{i_{1}},\nu_{i_{1}})|\geq 0$
at the point $y$. Consequently,
$\frac{\Delta u_{i_{1}}-(D^{2}u_{i_{1}})(\nu_{i_{1}},\nu_{i_{1}})}{|\nabla u_{i_{1}}|_{g}}-|\mathrm{tr}(q)-q(\nu_{i_{1}},\nu_{i_{1}})|\geq-C\,d(x,y)$
at the point $x$. Putting these facts together, we conclude that
$W_{\lambda}(x)\geq-C\lambda\,e^{-(\lambda r)^{\frac{1}{8}}}$
for each point $x\in G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}$. On the other hand, the intersection $\Sigma_{\lambda}\cap B_{r}(p)$ has area at most $Cr^{n-1}$. Consequently,
$\bigg{(}r^{\sigma+1-n}\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}.$
This completes the proof of Lemma 5.4. ∎
Lemma 5.5.
Assume that the Matching Angle Hypothesis holds. Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
$\bigg{(}r^{\sigma+1-n}\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\geq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap\{u_{i_{3}}\leq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}$
$\leq C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}$
for all pairwise distinct elements $i_{1},i_{2},i_{3}\in I$.
Proof.
The corresponding estimate for $V_{\lambda}$ follows from Lemma 3.14 in [Brendle]. We next observe that $|W_{\lambda}-V_{\lambda}|\leq C\leq C\lambda^{\frac{1}{8}}r^{-\frac{7}{8}}$. Moreover, it is shown in [Brendle] that the set $G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\geq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap\{u_{i_{3}}\leq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}$ has area at most $C\,(\lambda r)^{-\frac{7}{8}}\,r^{n-1}$. From this, the assertion follows. ∎
Lemma 5.6.
Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
$\bigg{(}r^{\sigma+1-n}\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{3}}\geq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}\leq C\,(\lambda r)^{1-\frac{3}{2\sigma}}$
for all pairwise distinct elements $i_{1},i_{2},i_{3}\in I$.
Proof.
The corresponding estimate for $V_{\lambda}$ follows from Lemma 3.15 in [Brendle]. We next observe that $|W_{\lambda}-V_{\lambda}|\leq C\leq C\lambda$. Moreover, it is shown in [Brendle] that the set $G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{3}}\geq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}\cap B_{r}(p)$ has area at most $C\,(\lambda r)^{-\frac{3}{2}}\,r^{n-1}$. From this, the assertion follows. ∎
Proposition 5.7.
For each $i\in I$, we assume that the hypersurface $\{u_{i}=0\}$ satisfies the condition $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$. Moreover, we assume that the Matching Angle Hypothesis is satisfied. Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$.
If $\lambda r$ is sufficiently large, then
$\bigg{(}r^{\sigma+1-n}\int_{\Sigma_{\lambda}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}$
$\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}+C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}+C\,(\lambda r)^{1-\frac{3}{2\sigma}}.$
Proof.
Combining Lemma 5.4, Lemma 5.5, and Lemma 5.6, we conclude that
$\bigg{(}r^{\sigma+1-n}\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}$
$\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}+C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}+C\,(\lambda r)^{1-\frac{3}{2\sigma}}$
for all pairwise distinct elements $i_{1},i_{2},i_{3}\in I$. On the other hand, $\Sigma_{\lambda}=\bigcup_{i_{1},i_{2},i_{3}}G_{\lambda}^{(i_{1},i_{2},i_{3})}$, where the union is taken over all pairwise distinct elements $i_{1},i_{2},i_{3}\in I$. Hence, the assertion follows by summation over $i_{1},i_{2},i_{3}$. This completes the proof of Proposition 5.7. ∎
Corollary 5.8.
For each $i\in I$, we assume that the hypersurface $\{u_{i}=0\}$ satisfies the condition $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$. Moreover, we assume that the Matching Angle Hypothesis is satisfied. Let us fix an exponent $\sigma\in[1,\frac{3}{2})$.
Then
$\sup_{p\in\mathbb{R}^{n}}\sup_{r\leq 1}\bigg{(}r^{\sigma+1-n}\int_{\Sigma_{\lambda}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}\to 0$
as $\lambda\to\infty$.
Proof.
Let us consider an arbitrary sequence $\lambda_{l}\to\infty$. By Proposition 5.2, we can find a sequence of positive real numbers $\delta_{l}\to 0$ such that
$(\delta_{l}\lambda_{l})^{-1}\sup_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}\to 0$
as $l\to\infty$. This implies
$\sup_{p\in\mathbb{R}^{n}}\sup_{r\leq(\delta_{l}\lambda_{l})^{-1}}\bigg{(}r^{\sigma+1-n}\int_{\Sigma_{\lambda_{l}}\cap B_{r}(p)}\max\{-W_{\lambda_{l}},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}$
$\leq(\delta_{l}\lambda_{l})^{-1}\sup_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}\to 0$
as $l\to\infty$. On the other hand, it follows from Proposition 5.7 that
$\sup_{p\in\mathbb{R}^{n}}\sup_{(\delta_{l}\lambda_{l})^{-1}\leq r\leq 1}\bigg{(}r^{\sigma+1-n}\int_{\Sigma_{\lambda_{l}}\cap B_{r}(p)}\max\{-W_{\lambda_{l}},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}\to 0$
as $l\to\infty$. Putting these facts together, the assertion follows. ∎
6. Proof of Theorem II
Let $\Omega$, $g$, $q$, $H$ satisfy the assumptions in Theorem II. Furthermore assume that $\Omega$ satisfies the Matching Angle hypothesis. We obtain the spinor bundle $\mathcal{S}$ and the twisted spinor bundle $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$ over an open set containing $\Omega$.
Let $U$ denote a Euclidean ball such that the closure of $U$ is contained in the interior of $\Omega$. Consider a sequence $\lambda_{l}\to\infty$. Given $l$, let $N^{(l)}\colon\Sigma_{\lambda_{l}}\to S^{n-1}$ denote the map from (5.1). Recall that $N^{(l)}$ has degree $1$. Note that $U\subset\Omega_{\lambda_{l}}$ if $l$ is sufficiently large.
Proposition 6.1.
Suppose that $K>0$ is a given positive real number. Then there exists a uniform constant $C$ (independent of $l$) such that
$\int_{\Omega_{\lambda_{l}}}F^{2}\leq C\int_{\Omega_{\lambda_{l}}}(|\nabla F|-KF)_{+}^{2}+C\int_{U}F^{2}$
for every nonnegative smooth function $F$ on $\Omega_{\lambda_{l}}$.
Proof.
Note that the hypersurface $\Sigma_{\lambda_{l}}=\partial\Omega_{\lambda_{l}}$ can be written as a radial graph with bounded slope. From this, it is easy to see that $\Omega_{\lambda_{l}}$ is bi-Lipschitz equivalent to the Euclidean unit ball, with constants that are independent of $l$. The assertion now follows from Proposition B.1. ∎
Corollary 6.2.
For each section $s\in C^{\infty}(\Omega_{\lambda_{l}},\mathcal{E})$, we have
$\int_{\Omega_{\lambda_{l}}}|s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s|^{2}+C\int_{U}|s|^{2},$
where $C$ is a constant that does not depend on $l$.
Proof.
If we choose $K$ to be a large multiple of $\sup_{\Omega}|q|$, then we can bound
$(|\nabla^{\mathcal{E}}s|-K\,|s|)_{+}\leq|\nabla^{\mathcal{E},q}s|.$
We now apply Proposition 6.1 to the function $F=(\delta^{2}+|s|^{2})^{\frac{1}{2}}$, and send $\delta\to 0$. This gives
$\int_{\Omega_{\lambda_{l}}}|s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}(|\nabla^{\mathcal{E}}s|-K\,|s|)_{+}^{2}+C\int_{U}|s|^{2}.$
Putting these facts together, the assertion follows. ∎
Proposition 6.3.
There exists a uniform constant $C$ (independent of $l$) such that
$\int_{\Sigma_{\lambda_{l}}}F^{2}\leq C\int_{\Omega_{\lambda_{l}}}|\nabla F|^{2}+C\int_{\Omega_{\lambda_{l}}}F^{2}$
for every smooth function $F$ on $\Omega_{\lambda_{l}}$.
Proof.
Note that the hypersurface $\Sigma_{\lambda_{l}}=\partial\Omega_{\lambda_{l}}$ can be written as a radial graph with bounded slope. From this, it is easy to see that $\Omega_{\lambda_{l}}$ is bi-Lipschitz equivalent to the Euclidean unit ball, with constants that are independent of $l$. Hence, the assertion follows from the Sobolev trace theorem. ∎
Corollary 6.4.
For each section $s\in C^{\infty}(\Omega_{\lambda_{l}},\mathcal{E})$, we have
$\int_{\Sigma_{\lambda_{l}}}|s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s|^{2}+C\int_{U}|s|^{2},$
where $C$ is a constant that does not depend on $l$.
Proof.
We apply Proposition 6.3 to the function $F=(\delta^{2}+|s|^{2})^{\frac{1}{2}}$, and send $\delta\to 0$. This gives
$\int_{\Sigma_{\lambda_{l}}}|s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E}}s|^{2}+C\int_{\Omega_{\lambda_{l}}}|s|^{2}.$
Hence, the assertion follows from Corollary 6.2. ∎
Proposition 6.5.
We have
$\int_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}F^{2}\leq o(1)\int_{\Omega_{\lambda_{l}}}|\nabla F|^{2}+o(1)\int_{\Omega_{\lambda_{l}}}F^{2}$
for every smooth function $F$ on $\Omega_{\lambda_{l}}$.
Proof.
This follows from Corollary 5.8 together with the Fefferman-Phong estimate (cf. [Brendle], Appendix A). ∎
Corollary 6.6.
For each section $s\in C^{\infty}(\Omega_{\lambda_{l}},\mathcal{E})$, we have
$\int_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}|s|^{2}\leq o(1)\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s|^{2}+o(1)\int_{U}|s|^{2}.$
Proof.
We apply Proposition 6.5 to the function $F=(\delta^{2}+|s|^{2})^{\frac{1}{2}}$, and send $\delta\to 0$. This gives
$\int_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}|s|^{2}\leq o(1)\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E}}s|^{2}+o(1)\int_{\Omega_{\lambda_{l}}}|s|^{2}.$
Hence, the assertion follows from Corollary 6.2. ∎
Proposition 6.7.
The following assertions hold.
  1. (a)
    There exists a $\nabla^{\mathcal{E},q}$-parallel section $s$ defined over $\Omega$. Considered as a homomorphism field, the section $s$ is invertible at each point in $\Omega$.
  2. (b)
    Each of the boundary faces is totally geodesic with respect to $g$.
  3. (c)
    $q^{\rm tan}=0$ along each of the boundary faces.
Having established Proposition 6.7, the arguments in Subsection 4.3 imply that $(\Omega,g,q)$ can be isometrically immersed as a spacelike hypersurface in the Minkowski spacetime $\mathbb{R}^{n,1}$ such that $q$ is the second fundamental form.
In the remainder of this section, we will give the proof of Proposition 6.7. We first treat the odd-dimensional case, and then the even-dimensional case.
6.1. Proof of Proposition 6.7 in odd dimensions
Suppose that $n\geq 3$ is an odd integer. By Proposition 3.8, we can find a non-trivial section $s^{(l)}\in C^{\infty}(\Omega_{\lambda_{l}},\mathcal{E})$ such that
$\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}$ $\leq\tfrac{1}{2}\int_{\Omega_{\lambda_{l}}}(-\mu+|J|)|s^{(l)}|^{2}$
$+\tfrac{1}{2}\int_{\Sigma_{\lambda_{l}}}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN^{(l)}\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}.$
Moreover, $\chi^{(l)}s^{(l)}=s^{(l)}$ at each point on the boundary $\Sigma_{\lambda_{l}}$. By assumption, $\mu-|J|\geq 0$. Using Corollary 6.6, we obtain
$\int_{\Sigma_{\lambda_{l}}}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN^{(l)}\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}$
$\leq o(1)\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}+o(1)\int_{U}|s^{(l)}|^{2}.$
Putting these facts together, we conclude that
$\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}\leq o(1)\int_{U}|s^{(l)}|^{2}$
if $l$ is sufficiently large. Using Corollary 6.2, we obtain
$\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}\leq C\int_{U}|s^{(l)}|^{2}$
if $l$ is sufficiently large. Since $\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}>0$, it follows that $\int_{U}|s^{(l)}|^{2}>0$ if $l$ is sufficiently large.
By scaling, we can arrange that $\int_{U}|s^{(l)}|^{2}=1$ if $l$ is sufficiently large. Then
(6.1) $\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}\leq o(1)$
and
$\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}\leq C.$
After passing to a subsequence if necessary, the sequence $s^{(l)}$ converges weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$. The limit $s$ is defined on the interior of $\Omega$ and satisfies $\int_{U}|s|^{2}=1$. Using (6.1), we obtain $\nabla^{\mathcal{E},q}s=0$. In particular, $s$ is smooth. Using $\nabla^{\mathcal{E},q}$-parallel transport along radial lines emanating from some point in the interior of $\Omega\subset\mathbb{R}^{n}$, we extend $s$ to a smooth section which is defined on an open set containing $\Omega$ and satisfies $\nabla^{\mathcal{E},q}s=0$ at each point in $\Omega$.
Proposition 6.8.
We have
$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\to 0$
as $l\to\infty$.
Proof.
Using Corollary 6.4, we obtain
$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}(s^{(l)}-s)|^{2}+C\int_{U}|s^{(l)}-s|^{2}.$
Since $\nabla^{\mathcal{E},q}s=0$, we conclude that
$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}+C\int_{U}|s^{(l)}-s|^{2}.$
The assertion now follows from (6.1) and the fact that $s^{(l)}\to s$ weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$. ∎
Corollary 6.9.
We have
$\int_{\Sigma_{\lambda_{l}}}|\chi^{(l)}s-s|^{2}\to 0$
as $l\to\infty$. Here, $\chi^{(l)}$ denotes the boundary operator on $\Sigma_{\lambda_{l}}$.
Proof.
Recall that $s^{(l)}$ satisfies the boundary condition $\chi^{(l)}s^{(l)}=s^{(l)}$ at each point on $\Sigma_{\lambda_{l}}$. Hence, the assertion follows from Lemma 6.8. ∎
For each $i_{0}\in I$, Assumption 5.1 implies that the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is non-empty. Moreover, it follows from Lemma 6.9 that $s$ satisfies the boundary condition
$\gamma(\nu_{i_{0}})\circ s=s\circ\gamma_{0}(N_{i_{0}})$
on $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$.
Lemma 6.10.
Let $\sigma\in S_{0}$. If $s\sigma$ vanishes at some point of $\Omega$, then it vanishes everywhere on $\Omega$.
Proof.
The proof is analogous to the proof of Lemma 4.4. ∎
As above, we define
$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ each point on $\Omega$}\}.$
Lemma 6.11.
The subspace $L\subset S_{0}$ is invariant under $\gamma_{0}(N_{i})$ for each $i\in I$.
Proof.
The proof is analogous to the proof of Lemma 4.5. ∎
Lemma 6.12.
We have $L=\{0\}$. Moreover, the section $s$, considered as a homomorphism field, is invertible on each point in $\Omega$.
Proof.
This follows from Lemma 6.11 together with the fact that $\text{\rm span}\{N_{i}:i\in I\}=\mathbb{R}^{n}$. ∎
Lemma 6.13.
Let $i_{0}\in I$. Along the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$, we have $h(X)=q(X)^{\text{\rm tan}}=0$ whenever $X$ is a tangent vector to that boundary face.
Proof.
Recall that
$\gamma(\nu_{i_{0}})\circ s=s\circ\gamma_{0}(N_{i_{0}})$
on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$. In the next step, we differentiate this relation in tangential direction. Arguing as in the proof of Lemma 4.2, we obtain
$\gamma(h(X))\circ s+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\circ s=0$
whenever $p$ is a point on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ and $X$ is a tangent vector to that boundary face. Using Lemma 6.12, we conclude that
$\gamma(h(X))+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p})$
whenever $p$ is a point on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ and $X$ is a tangent vector to that boundary face. Note that $i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\in\text{\rm End}(\mathcal{S}|_{p})$ is self-adjoint, while $\gamma(h(X))\in\text{\rm End}(\mathcal{S}|_{p})$ is skew-adjoint. Thus, we conclude that
$\gamma(h(X))=0\in\text{\rm End}(\mathcal{S}|_{p}),\quad\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p})$
whenever $p$ is a point on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ and $X$ is a tangent vector to that boundary face. From this, the assertion follows. ∎
6.2. Proof of Proposition 6.7 in even dimensions
Suppose now that $n\geq 2$ is an even integer. By Corollary 3.12, we can find a non-trivial section $s^{(l)}\in H^{1}(\Omega_{\lambda_{l}},\mathcal{E})$ such that
$\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}$ $\leq\tfrac{1}{2}\int_{\Omega_{\lambda_{l}}}(-\mu+|J|)|s^{(l)}|^{2}$
$+\tfrac{1}{2}\int_{\Sigma_{\lambda_{l}}}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN^{(l)}\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}.$
Moreover, we can find an element $\hat{t}_{l}\in S^{1}$ such that $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}^{(l)})$ on $\Sigma_{\lambda_{l}}$, where the map $\hat{N}^{(l)}:\Sigma_{\lambda_{l}}\to S^{n}$ is defined by $\hat{N}^{(l)}(x)=\varphi(N^{(l)}(x),\hat{t}_{l})$ for all $x\in\Sigma_{\lambda_{l}}$. By assumption, $\mu-|J|\geq 0$. Using Corollary 6.6, we obtain
$\int_{\Sigma_{\lambda_{l}}}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN^{(l)}\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}$
$\leq o(1)\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}+o(1)\int_{U}|s^{(l)}|^{2}.$
Putting these facts together, we conclude that
$\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}\leq o(1)\int_{U}|s^{(l)}|^{2}$
if $l$ is sufficiently large. Using Corollary 6.2, we obtain
$\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}\leq C\int_{U}|s^{(l)}|^{2}$
if $l$ is sufficiently large. Since $\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}>0$, it follows that $\int_{U}|s^{(l)}|^{2}>0$ if $l$ is sufficiently large.
By scaling, we can arrange that $\int_{U}|s^{(l)}|^{2}=1$ if $l$ is sufficiently large. Then
(6.2) $\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}\leq o(1)$
and
$\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}\leq C.$
After passing to a subsequence if necessary, we may assume that $\hat{t}_{l}\to\hat{t}$. Moreover, the sequence $s^{(l)}$ converges weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$. The limit $s$ is defined on the interior of $\Omega$ and satisfies $\int_{U}|s|^{2}=1$. Using (6.2), we obtain $\nabla^{\mathcal{E},q}s=0$. In particular, $s$ is smooth. Using $\nabla^{\mathcal{E},q}$-parallel transport along radial lines emanating from some point in the interior of $\Omega\subset\mathbb{R}^{n}$, we extend $s$ to a smooth section which is defined on an open set containing $\Omega$ and satisfies $\nabla^{\mathcal{E},q}s=0$ at each point in $\Omega$.
Proposition 6.14.
We have
$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\to 0$
as $l\to\infty$.
Proof.
The proof is analogous to the proof of Proposition 6.8. ∎
Corollary 6.15.
We have
$\int_{\Sigma_{\lambda_{l}}}|\gamma(\nu)\circ s-s\circ\gamma_{0}(\hat{N}^{(l)})|^{2}\to 0$
as $l\to\infty$. Here, $\chi^{(l)}$ denotes the boundary operator on $\Sigma_{\lambda_{l}}$.
Proof.
Recall that $s^{(l)}$ satisfies the boundary condition $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}^{(l)})$ on $\Sigma_{\lambda_{l}}$ at each point on $\Sigma_{\lambda_{l}}$. Hence, the assertion follows from Lemma 6.14. ∎
For each $i\in I$, we define $\hat{N}_{i}=\varphi(N_{i},\hat{t})\in S^{n}$. For each $i_{0}\in I$, Assumption 5.1 implies that the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is non-empty. Moreover, it follows from Lemma 6.15 that $s$ satisfies the boundary condition
$\gamma(\nu_{i_{0}})\circ s=s\circ\gamma_{0}(\hat{N}_{i_{0}})$
on $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$. By Lemma 3.2 in [Brendle], the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is a dense subset of $\Omega\cap\{u_{i_{0}}=0\}$. Thus, we conclude that
$\gamma(\nu_{i_{0}})\circ s=s\circ\gamma_{0}(\hat{N}_{i_{0}})$
on $\Omega\cap\{u_{i_{0}}=0\}$.
Lemma 6.16.
Suppose that $i_{1},i_{2}\in I$ satisfy $\Omega\cap\{u_{i_{1}}=0\}\cap\{u_{i_{2}}=0\}\neq\emptyset$. Then $\langle N_{i_{1}},N_{i_{2}}\rangle=\langle\hat{N}_{i_{1}},\hat{N}_{i_{2}}\rangle$.
Proof.
We consider a point $p\in\Omega\cap\{u_{i_{1}}=0\}\cap\{u_{i_{2}}=0\}$. Then
$\gamma(\nu_{i_{1}})\circ s=s\circ\gamma_{0}(\hat{N}_{i_{1}})$
and
$\gamma(\nu_{i_{2}})\circ s=s\circ\gamma_{0}(\hat{N}_{i_{2}})$
at the point $p$. This implies
$\langle\nu_{i_{1}},\nu_{i_{2}}\rangle_{g}\,|s|^{2}$ $=\tfrac{1}{2}\,\langle\gamma(\nu_{i_{1}})\circ s,\gamma(\nu_{i_{2}})\circ s\rangle+\tfrac{1}{2}\,\langle\gamma(\nu_{i_{2}})\circ s,\gamma(\nu_{i_{1}})\circ s\rangle$
$=\tfrac{1}{2}\,\langle s\circ\gamma_{0}(\hat{N}_{i_{1}}),s\circ\gamma_{0}(\hat{N}_{i_{2}})\rangle+\tfrac{1}{2}\,\langle s\circ\gamma_{0}(\hat{N}_{i_{2}}),s\circ\gamma_{0}(\hat{N}_{i_{1}})\rangle$
$=\langle\hat{N}_{i_{1}},\hat{N}_{i_{2}}\rangle\,|s|^{2}$
at the point $p$. On the other hand, the Matching Angle Hypothesis implies that $\langle\nu_{i_{1}},\nu_{i_{2}}\rangle_{g}\,|s|^{2}=\langle N_{i_{1}},N_{i_{2}}\rangle\,|s|^{2}$ at the point $p$. Putting these facts together, we conclude that $\langle N_{i_{1}},N_{i_{2}}\rangle\,|s|^{2}=\langle\hat{N}_{i_{1}},\hat{N}_{i_{2}}\rangle\,|s|^{2}$ at the point $p$. Since $|s|^{2}>0$ at each point in $\Omega$, the assertion follows. ∎
Lemma 6.17.
The map $\varphi(\cdot,\hat{t})\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n}$.
Proof.
Suppose that the assertion is false. By Lemma 3.9, the map $\varphi(\cdot,\hat{t})$ is Lipschitz continuous with a Lipschitz constant strictly less than $1$. This contradicts Lemma 6.16. ∎
Lemma 6.18.
Let $\sigma\in S_{0}$. If $s\sigma$ vanishes at some point of $\Omega$, then it vanishes everywhere on $\Omega$.
Proof.
The proof is analogous to the proof of Lemma 4.4. ∎
We again define
$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ each point on $\Omega$}\}.$
Lemma 6.19.
The subspace $L\subset S_{0}$ is invariant under $\gamma_{0}(\hat{N}_{i})$ for each $i\in I$.
Proof.
The proof is analogous to the proof of Lemma 4.5. ∎
Lemma 6.20.
We have $L=\{0\}$. Moreover, the section $s$, considered as a homomorphism field, is invertible on each point in $\Omega$.
Proof.
Note that $\text{\rm span}\{N_{i}:i\in I\}=\mathbb{R}^{n}$. Using Lemma 6.17, it follows that $\text{\rm span}\{\hat{N}_{i}:i\in I\}=\mathbb{R}^{n}\subset\mathbb{R}^{n+1}$. By Lemma 6.19, $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n}$. Since the volume element acts as a scalar, we deduce that $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n+1}$. Arguing as in the proof of Lemma 4.6, the assertion follows. ∎
Lemma 6.21.
Let $i_{0}\in I$. Along the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$, we have $h(X)=q(X)^{\text{\rm tan}}=0$ whenever $X$ is a tangent vector to that boundary face.
Proof.
The proof is analogous to the proof of Lemma 6.13. ∎
Appendix A A metric characterization of the ball
Theorem A.1.
Let $(M,g)$ be a compact connected Riemannian manifold of dimension $n\geq 2$ with boundary such that the following conditions are satisfied:
  • $g$ is flat.
  • The second fundamental form of $\partial M$ is positive semi-definite.
  • There exists a point in $\partial M$, where the second fundamental form is positive definite.
Then $(M,g)$ is isometric to a compact convex domain in $\mathbb{R}^{n}$ with smooth boundary. In particular, $M$ is diffeomorphic to a ball.
Proof.
For small enough $\varepsilon>0$, we consider the smooth manifold
$M^{\prime}:=M\cup_{\sim}(\partial M\times(-\varepsilon,\infty)),$
where $\sim$ identifies $\partial M\times(-\varepsilon,0]\subset\partial M\times(-\varepsilon,\infty)$ with an open neighborhood of $\partial M\subset M$ by the normal exponential map along $\partial M$ with respect to $g$. Denote the metric on $\partial M$ induced by $g$ by $g_{0}$ and the second fundamental form of $\partial M$ (considered as a $(1,1)$-tensor field) by $h$. The Gauss lemma, the Riccati equation for the second fundamental form and the vanishing of the curvature of $M$ imply that on $\partial M\times(-\varepsilon,0]$ the metric $g$ takes the form $g=dt^{2}+g_{t}$ where $g_{t}$ is the metric on $\partial M$ given by $g_{t}(X,Y)=g_{0}((\textup{id}+th)^{2}X,Y)$. Here $t\in(-\varepsilon,\infty)$ is the standard parameter. Since $h$ is positive semi-definite, we can use this formula to extend $g$ to $\partial M\times[0,\infty)$ and hence to a smooth metric $g^{\prime}$ on all of $M^{\prime}$. By Theorem 7.2 in [BGM05], the manifold $(M^{\prime},g^{\prime})$ is flat. Since the projection $(\partial M\times[0,\infty),dt^{2}+g_{t})\to(\partial M,g_{0})$ is $1$-Lipschitz, it is easy to see that $(M^{\prime},g^{\prime})$ is complete.
Let $\tilde{M}^{\prime}$ be the universal covering of $M^{\prime}$ equipped with the complete flat Riemannian metric $\tilde{g}^{\prime}$ induced by $g^{\prime}$. Since $(\tilde{M}^{\prime},\tilde{g}^{\prime})$ is complete, flat and simply connected, it is isometric to Euclidean space $\mathbb{R}^{n}$. Using this isometry, we regard the preimage $\tilde{M}\subset\tilde{M}^{\prime}$ of $M\subset M^{\prime}$ as a closed subset of $\mathbb{R}^{n}$ with smooth boundary $\partial\tilde{M}$. Since the embedding $M\hookrightarrow M^{\prime}$ is a homotopy equivalence, $\tilde{M}$ is connected and $\tilde{M}$ is the universal covering of $M$.
We first assume that $n\geq 3$. By assumption, the second fundamental form of $\partial\tilde{M}\subset\mathbb{R}^{n}$ is positive semi-definite and there exists a connected component $\Sigma\subset\partial\tilde{M}$ with the property that the second fundamental form is positive definite at some point on $\Sigma$. This implies that the sectional curvature of the induced Riemannian metric on $\Sigma\subset\tilde{M}$ is non-negative and positive at some point. This Riemannian metric of $\Sigma$ is complete because it is pulled back from a compact manifold.
By the main theorem in [dCL] (see also [Sacksteder]), $\Sigma=\partial K\subset\mathbb{R}^{n}$ where $K\subset\mathbb{R}^{n}$ is convex closed subset in $\mathbb{R}^{n}$ with smooth boundary. Furthermore, $\Sigma$ is diffeomorphic to $S^{n-1}$ or to $\mathbb{R}^{n-1}$, hence it is the universal covering of a component of $\partial M$ with the induced metric. In the case $\Sigma\approx\mathbb{R}^{n-1}$, the Cheeger-Gromoll splitting theorem (Theorem 3 in [CG]) implies that the induced metric on $\Sigma$ is flat, which is a contradiction.
Hence $\Sigma\approx S^{n-1}$ and $K$ is a compact convex domain in $\mathbb{R}^{n}$ with smooth boundary. As $\tilde{M}$ is connected and complete, we have $\tilde{M}\subset K$ as a closed, hence compact subset. Assume $K\setminus\tilde{M}\neq\emptyset$. Then there exists a connected component $\Sigma^{\prime}\neq\Sigma$ of $\partial\tilde{M}$ and, again by [dCL], $\Sigma^{\prime}=\partial K^{\prime}$ for a convex closed subset $K^{\prime}\subset{\rm int}(K)\subset\mathbb{R}^{n}$. Since $\Sigma^{\prime}\subset\mathrm{int}(K)$, we have that $K^{\prime}\subset\mathrm{int}(K)$. In particular, $K^{\prime}\subset K$ and hence $K^{\prime}$ is compact. Arguing as before, we obtain $\tilde{M}\subset K^{\prime}$. Thus $\tilde{M}\subset\mathrm{int}(K)$, which is a contradiction.
Hence $\tilde{M}=K$ and $\tilde{M}$ is diffeomorphic to a ball. By the Brouwer fixed point theorem, we have $M=\tilde{M}$. This implies the assertion.
Now assume $n=2$. Then the Gauss-Bonnet formula shows that the Euler characteristic of $M$ is positive, which implies that $M$ is diffeomorphic to a $2$-ball and, in particular, simply connected. Since the inclusion $M\hookrightarrow M^{\prime}$ is a homotopy equivalence, we get $\tilde{M}^{\prime}=M^{\prime}$. In particular, $(M^{\prime},g^{\prime})$ is isometric to Euclidean space $\mathbb{R}^{2}$ and $(M,g)\subset(M^{\prime},g^{\prime})$ is isometric to a compact domain $K\subset\mathbb{R}^{2}$ with smooth, mean-convex boundary. This implies the assertion. ∎
Appendix B An auxiliary estimate
Proposition B.1.
Let $\Omega\subset\mathbb{R}^{n}$ be a compact, connected Lipschitz domain. Assume that $U$ is a non-empty open subset of $\Omega$. Suppose that $K>0$ is a given positive real number. Then
$\int_{\Omega}F^{2}\leq C\int_{\Omega}(|\nabla F|-KF)_{+}^{2}+C\int_{U}F^{2}$
for every nonnegative function $F\in H^{1}(\Omega)$. Here, $C$ is a positive constant that depends on $\Omega$, $U$, and $K$, but not on $F$.
Proof.
Suppose the assertion is false. Then we can find a sequence of nonnegative functions $F_{j}\in H^{1}(\Omega)$ such that $\int_{\Omega}F_{j}^{2}=1$, $\int_{U}F_{j}^{2}\to 0$, and
$\int_{\Omega}(|\nabla F_{j}|-KF_{j})_{+}^{2}\to 0.$
This implies
$\int_{\Omega}(-F_{j})_{+}^{2}\to 0.$
Moreover, using the triangle inequality, we obtain
$\int_{\Omega}|\nabla F_{j}|^{2}\leq C.$
After passing to a subsequence, the sequence $F_{j}$ converges to a function $F$ in $L^{2}(\Omega)$. The function $F$ satisfies $\int_{\Omega}F^{2}=1$. Moreover, $F=0$ almost everywhere in $U$, and $F\geq 0$ almost everywhere in $\Omega$.
Let $\eta:\mathbb{R}^{n}\to[0,\infty)$ denote a smooth function which is supported in the unit ball and satisfies $\int_{\mathbb{R}^{n}}\eta=1$. For each $\varepsilon>0$, we define a smooth function $\eta_{\varepsilon}:\mathbb{R}^{n}\to[0,\infty)$ by
$\eta_{\varepsilon}(x)=\frac{1}{\varepsilon^{n}}\eta\left(\frac{x}{\varepsilon}\right)$
for all $x\in\mathbb{R}^{n}$. The function $\eta_{\varepsilon}$ is supported in a ball of radius $\varepsilon$ and $\int_{\mathbb{R}^{n}}\eta_{\varepsilon}=1$.
Let us fix an arbitrary point $p\in U$. Let $\varepsilon$ be a real number such that $0<\varepsilon<d(p,\partial U)$. Then $d(p,\partial\Omega)>\varepsilon$. We denote by $\Omega_{\varepsilon}$ the connected component of the set $\{x\in\Omega:d(x,\partial\Omega)>\varepsilon\}$ containing $p$. Let $F_{j}^{\varepsilon}\colon\Omega_{\varepsilon}\to\mathbb{R}$ denote the convolution of $F_{j}$ with $\eta_{\varepsilon}$, and let $F^{\varepsilon}\colon\Omega_{\varepsilon}\to\mathbb{R}$ denote the convolution of $F$ with $\eta_{\varepsilon}$. Then $|\nabla F_{j}^{\varepsilon}|\leq\eta_{\varepsilon}*|\nabla F_{j}|$ at each point in $\Omega_{\varepsilon}$. This implies
$|\nabla F_{j}^{\varepsilon}|-KF_{j}^{\varepsilon}\leq(\eta_{\varepsilon}*|\nabla F_{j}|)-K\,(\eta_{\varepsilon}*F_{j})\leq\eta_{\varepsilon}*(|\nabla F_{j}|-KF_{j})_{+}$
at each point in $\Omega_{\varepsilon}$. We now pass to the limit as $j\to\infty$, keeping $\varepsilon$ fixed. Since $F_{j}\to F$ in $L^{2}(\Omega)$, we know that $F_{j}^{\varepsilon}\to F^{\varepsilon}$ and $|\nabla F_{j}^{\varepsilon}|\to|\nabla F^{\varepsilon}|$ at each point in $\Omega_{\varepsilon}$. Moreover, since $(|\nabla F_{j}|-KF_{j}\big{)}_{+}\to 0$ in $L^{2}(\Omega)$, it follows that $\eta_{\varepsilon}*(|\nabla F_{j}|-KF_{j})_{+}\to 0$ at each point in $\Omega_{\varepsilon}$. Putting these facts together, we conclude that
$|\nabla F^{\varepsilon}|-KF^{\varepsilon}\leq 0$
at each point in $\Omega_{\varepsilon}$. On the other hand, we know that $F=0$ almost everywhere in $U$. Since $d(p,\partial U)>\varepsilon$, it follows that $F^{\varepsilon}$ vanishes at the point $p$. Since $\Omega_{\varepsilon}$ is connected, standard ODE arguments imply that $F^{\varepsilon}=0$ at each point in $\Omega_{\varepsilon}$.
Finally, we send $\varepsilon\to 0$. Since $\Omega$ is connected, we conclude that $F=0$ almost everywhere in $\Omega$. This contradicts the fact that $\int_{\Omega}F^{2}=1$. ∎

### v4

@@
 1. Introduction
-An initial data set is a triple $(M,g,q)$, where $M$ is a manifold, $g$ is a Riemannian metric on $M$, and $q$ is a symmetric $(0,2)$-tensor on $M$.
-Denote the scalar curvature of $g$ by $R$.
+An initial data set is a triplet $(M,g,q)$, where $M$ is a manifold, $g$ is a Riemannian metric on $M$, and $q$ is a symmetric $(0,2)$-tensor on $M$.
+We denote the scalar curvature of $g$ by $R$.
 We define
@@
 In particular, $(M,g,q)$ satisfies the dominant energy condition.
-In this paper we will formulate conditions which insure that a given initial data set is of the form described in this example.
-The study of rigidity for initial data sets under the dominant energy condition can be traced back to Schoen and Yau’s work [SY81], in which they proved the spacetime positive energy theorem using minimal surface techniques in dimension 3.
-Eichmair, Huang, Lee, and Schoen [EHLS] later extended this work by proving a more general spacetime positive mass theorem up to dimension 7.
-Huang and Lee [HL] subsequently affirmed the rigidity statement of this theorem studied in [EHLS].
-In a different approach, Witten [Wit] used spinors to prove the spacetime positive mass theorem, a method that generalizes to all higher dimensions for spin manifolds.
-Parker and Taubes [PT] gave a mathematically rigorous formulation of Witten’s argument.
-Chruściel and Maerten [CM] later provided a rigorous proof of the desired rigidity statement in all dimensions for spin manifolds, building on Beig and Chruściel’s proof [BC] in dimension 3.
+The study of initial data sets under the dominant energy condition can be traced back to Schoen and Yau’s work [Schoen-Yau], in which they proved the spacetime positive energy theorem using minimal surface techniques in dimension 3.
+Eichmair, Huang, Lee, and Schoen [Eichmair-Huang-Lee-Schoen] later extended this work by proving a more general spacetime positive mass theorem up to dimension 7.
+Huang and Lee [Huang-Lee] subsequently proved the rigidity statement in this theorem.
+In a different approach, Witten [Witten] used spinors to prove the spacetime positive mass theorem, a method that generalizes to all higher dimensions for spin manifolds.
+Parker and Taubes [Parker-Taubes] gave a mathematically rigorous formulation of Witten’s argument.
+Chruściel and Maerten [Chrusciel-Maerten] later provided a rigorous proof of the desired rigidity statement in all dimensions for spin manifolds, building on Beig and Chruściel’s proof [Beig-Chrusciel] in dimension 3.
 Lee’s book [Lee] offers a comprehensive exposition on the positive mass theorem.
+Instead of noncompact asymptotically flat manifolds, we consider compact manifolds with boundary.
+The condition of asymptotic flatness is replaced by suitable boundary conditions.
+In the interior, we still demand the dominant energy condition.
+First rigidity results have been obtained in this context by Eichmair, Galloway and Mendes in [Eichmair-Galloway-Mendes] and by Glöckle in [Gloe].
 We consider the following class of manifolds with boundary, specializing Example 1.1.
@@
 Let $K\subset\mathbb{R}^{n}$ be a compact convex domain with smooth boundary $\partial K$ and exterior normal $N\colon\partial K\to S^{n-1}$.
-Note that $N$ has degree $+1$.
-Let $f\colon K\to\mathbb{R}$ be a smooth map satisfying
+Note that $N$ has degree $1$.
+Let $f\colon K\to\mathbb{R}$ be a smooth map satisfying the following conditions:
   1. (a)
@@
     At each point on $\partial K$, we have $df(N)=0$ or $dN=0$.
-Let
-[…]
-be the graph of $f$. Then $M$ is a compact smooth spacelike hypersurface with boundary $\Sigma=\partial K\times\{0\}\subset\mathbb{R}^{n}\times\{0\}\subset\mathbb{R}^{n,1}$, see Figures 1 and 2.
+Let $M:=\{(x,f(x))\colon x\in K\}\subset\mathbb{R}^{n,1}$ denote the graph of $f$.
+Then $M$ is a compact smooth spacelike hypersurface with boundary $\Sigma=\partial K\times\{0\}\subset\mathbb{R}^{n}\times\{0\}\subset\mathbb{R}^{n,1}$, see Figures 1 and 2.
 […]
@@
 Figure 2. The case $df(N)\neq 0$ and $dN=0$
-Let $I\!I$ denote the vector-valued second fundamental form of $\Sigma$, viewed as a submanifold of $\mathbb{R}^{n,1}$ of codimension $2$. Then
-$I\!I(X,Y)=-\langle dN(X),Y\rangle\,N$
+Let $\mathrm{II}$ denote the vector-valued second fundamental form of $\Sigma$, viewed as a submanifold of $\mathbb{R}^{n,1}$ of codimension $2$. Then
+$\mathrm{II}(X,Y)=-\langle dN(X),Y\rangle\,N$
 for all tangent vectors $X,Y\in T\Sigma$.
-We denote by $T$ the future-oriented normal field to $M$ in $\mathbb{R}^{n,1}$, where $T$ is normalized so that $\langle T,T\rangle=-1$.
+We denote by $T$ the future-oriented normal vector field to $M$ in $\mathbb{R}^{n,1}$, where $T$ is normalized so that $\langle T,T\rangle_{\mathbb{R}^{n,1}}=-1$.
 Let $g$ denote the induced Riemannian metric and let $q$ denote the second fundamental form of $M$ with respect to $T$.
-In other words, $q(X,Y)=\langle\bar{D}_{X}T,Y\rangle$ for all $X,Y\in TM$, where $\bar{D}$ denotes the standard flat connection on $\mathbb{R}^{n,1}$.
+In other words, $q(X,Y)=\langle\bar{D}_{X}T,Y\rangle_{\mathbb{R}^{n,1}}$ for all $X,Y\in TM$, where $\bar{D}$ denotes the standard flat connection on $\mathbb{R}^{n,1}$.
 The initial data set $(M,g,q)$ satisfies the dominant energy condition, cf.
@@
 At each point on $\Sigma$, we denote by $\nu$ the outward-pointing unit normal vector to $\Sigma$ in $M$.
-Moreover, we denote by $h$ the second fundamental form of $\Sigma$, viewed as a hypersurface in $M$, with respect to $\nu$.
-In other words, $h(X,Y)=\langle D_{X}\nu,Y\rangle$ for $X,Y\in T\Sigma$, where $D$ denotes the Levi-Civita connection on $(M,g)$.
+Moreover, we denote by $h$ the second fundamental form of $\Sigma$, viewed as a hypersurface in $M$.
+In other words, $h(X,Y)=\langle D_{X}\nu,Y\rangle_{\mathbb{R}^{n,1}}$ for $X,Y\in T\Sigma$, where $D$ denotes the Levi-Civita connection on $(M,g)$.
 We will use these sign conventions for $q$ and $h$ throughout our paper.
 With this understood,
-$I\!I(X,Y)=-h(X,Y)\,\nu+q(X,Y)\,T$
-for all tangent vectors $X,Y\in T\Sigma$. Putting these facts together, we conclude that
-$h(X,Y)\,\nu-q(X,Y)\,T=\langle dN(X),Y\rangle\,N$
-for all tangent vectors $X,Y\in T\Sigma$. In particular,
-$h(X,Y)=\langle dN(X),Y\rangle\,\langle N,\nu\rangle$
+$h(X,Y)\,\nu-q(X,Y)\,T=-\mathrm{II}(X,Y)=\langle dN(X),Y\rangle\,N$
+for all tangent vectors $X,Y\in T\Sigma$. As $\langle\nu,\nu\rangle_{\mathbb{R}^{n,1}}=1$, $\langle T,T\rangle_{\mathbb{R}^{n,1}}=-1$, and $\langle\nu,T\rangle_{\mathbb{R}^{n,1}}=0$, it follows that
+$h(X,Y)=\langle dN(X),Y\rangle\,\langle N,\nu\rangle_{\mathbb{R}^{n,1}}$
 and
-$q(X,Y)=\langle dN(X),Y\rangle\,\langle N,T\rangle$
+$q(X,Y)=\langle dN(X),Y\rangle\,\langle N,T\rangle_{\mathbb{R}^{n,1}}$
 for all tangent vectors $X,Y\in T\Sigma$.
-We now use Assumption (c).
-Hence, at each point on $\Sigma$, we either have $dN=0$, or else $\langle N,\nu\rangle=1$ and $\langle N,T\rangle=0$.
+Assumption (c) implies that, at each point on $\Sigma$, we either have $dN=0$, or $\langle N,\nu\rangle_{\mathbb{R}^{n,1}}=1$ and $\langle N,T\rangle_{\mathbb{R}^{n,1}}=0$.
 Consequently, $h(X,Y)=\langle dN(X),Y\rangle$ and $q(X,Y)=0$ for all tangent vectors $X,Y\in T\Sigma$.
-In particular, $H=\|dN\|_{\mathrm{tr}}$ and $\mathrm{tr}(q)-q(\nu,\nu)=0$ at each point on $\Sigma$.
+In particular, we have $H=\|dN\|_{\mathrm{tr}}$ and $\mathrm{tr}(q)-q(\nu,\nu)=0$ at each point on $\Sigma$.
 Here, $H$ denotes the unnormalized mean curvature, i.e.
-the sum of the principal curvatures, of $\Sigma$ in $M$.
+the sum of the principal curvatures of $\Sigma$ in $M$.
 In other words, $H$ is equal to the trace of $h$, considered as an endomorphism field.
-Moreover, $\|\cdot\|_{\mathrm{tr}}$ denotes the trace norm of a linear map, i.e., the sum of all singular values.
+Moreover, $\|\cdot\|_{\mathrm{tr}}$ denotes the trace norm of a linear map, i.e.
+the sum of all singular values.
 Our first result identifies this class of examples as an extremal case of initial data sets.
@@
 Let $(M,g,q)$ be an initial data set.
-Assume that $M$ is a compact, connected spin manifold, has dimension $n\geq 2$ and nonempty boundary $\partial M=\Sigma$.
+Assume that $M$ is a compact connected spin manifold, has dimension $n\geq 2$ and nonempty boundary $\partial M=\Sigma$.
 Let $N\colon\Sigma\to S^{n-1}$ be a smooth map.
-We assume that
-  • $(M,g,q)$ satisfies the dominant energy condition,
-  • $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ holds along $\Sigma$,
-  • $N$ has positive degree.
+We assume that the following conditions are satisfied:
+  • $(M,g,q)$ satisfies the dominant energy condition.
+  • $H\geq\|dN\|_{\mathrm{tr}}+|\mathrm{tr}(q)-q(\nu,\nu)|$ holds along $\Sigma$.
+  • The map $N\colon\Sigma\to S^{n-1}$ has positive degree.
 Then $(M,g)$ is isometric to a manifold appearing in Example 1.2 such that $q$ is the second fundamental form of $M$ in $\mathbb{R}^{n,1}$.
@@
 The quantity $H-|\mathrm{tr}(q)-q(\nu,\nu)|$ is related to the notion of marginally trapped surfaces in general relativity.
-Recall that a marginally outer trapped surface (MOTS) is a hypersurface in an initial data set $(M,g,q)$ satisfying $H+(\mathrm{tr}(q)-q(\nu,\nu))=0$; see [EHLS]*p.
+Recall that a marginally outer trapped surface (MOTS) is a hypersurface in an initial data set $(M,g,q)$ satisfying $H+(\mathrm{tr}(q)-q(\nu,\nu))=0$; see [Eichmair-Huang-Lee-Schoen]*p.
 89.
@@
 We can choose $N\colon\partial M=\{a,b\}\to S^{0}=\{-1,1\}$ to be bijective.
-Then all quantities $R$, $\mu$, $H$, $J$, $dN$, and $\mathrm{tr}(q)-q(\nu,\nu)$ vanish.
+Then all quantities $R$, $\mu$, $H$, $J$, dNdN, and $\mathrm{tr}(q)-q(\nu,\nu)$ vanish.
 Thus the assumptions of Theorem I are trivially satisfied.
-The conclusion is now that there is a spacelike curve $[a,b]\to\mathbb{R}^{1,1}$, parametrized by arc-length and with curvature $\kappa$.
+The conclusion is now that there is a spacelike curve $a,b]\to\mathbb{R}^{1,1}$, parametrized by arc-length and with curvature $\kappa$.
 Indeed, this can be shown directly by ODE methods.
@@
 Corollary 1.6.
-Let $(M,g)$ be a compact connected Riemannian spin manifold of dimension $n\geq 2$ and nonempty boundary $\partial M=\Sigma$. Let $N\colon\Sigma\to S^{n-1}$ be a smooth map. We assume that
-  • $R\geq 0$ holds on $M$,
-  • $H\geq\|dN\|_{\mathrm{tr}}$ holds along $\Sigma$,
-  • $N$ has positive degree.
-Then $M$ can be isometrically embedded in $\mathbb{R}^{n}$ as a compact convex domain with smooth boundary and exterior normal $N$.
-In the second part of this paper, we consider initial data sets of the form $(\Omega,g,q)$, where $\Omega$ is a compact, convex polytope in $\mathbb{R}^{n}$ with non-empty interior.
-When $q=0$, the dominant energy condition simplifies to the non-negativity of the scalar curvature $R\geq 0$, and the boundary condition reduces to mean convexity $H\geq 0$, allowing us to compare $(\Omega,g)$ with Euclidean polytopes.
-Gromov [Gro14] initiated the study of scalar curvature comparison for polytopes with nonnegative scalar curvature to explore the notion of scalar curvature lower bounds in low-regularity spaces, and this problem has been extensively investigated; see [Gro14, Gro19, Gro22, Li17, Li19, WXY].
-For $q=g$, the dominant energy condition translates to $R\geq-n(n-1)$.
-Li [Li20] established a polyhedral comparison result for metrics with scalar curvature at least $-n(n-1)$.
-To explore the notion of energy density lower bounds in spacetimes with low regularity, we extend Gromov’s initial work [Gro14] to initial data sets under the dominant energy condition.
-In the case where $\Omega$ is a cube, Tsang [Tsang] addressed this problem in dimension 3 using the spacetime harmonic function approach.
+Let $(M,g)$ be a compact connected Riemannian spin manifold of dimension $n\geq 2$ and nonempty boundary $\partial M=\Sigma$.
+Let $N\colon\Sigma\to S^{n-1}$ be a smooth map.
+We assume that the following conditions are satisfied:
+  • $R\geq 0$ holds on $M$.
+  • $H\geq\|dN\|_{\mathrm{tr}}$ holds along $\Sigma$.
+  • The map $N\colon\Sigma\to S^{n-1}$ has positive degree.
+Then $M$ can be isometrically embedded in $\mathbb{R}^{n}$ as a compact convex domain with smooth boundary.
+Moreover, the map $N$ agrees, up to an element of $O(n)$, with the exterior normal $\partial K\to S^{n-1}$.
+The proof of Theorem I uses a boundary value problem for the Dirac operator.
+Boundary value problems for the Dirac operator have been extensively studied thanks to the work of Hörmander [Hormander], and the first named author and Ballmann [Baer-Ballmann, Baer-Ballmann13].
+In the second part of this paper, we consider initial data sets of the form $(\Omega,g,q)$, where $\Omega$ is a compact convex polytope in $\mathbb{R}^{n}$ with nonempty interior.
+When $q=0$, the dominant energy condition simplifies to the nonnegativity of the scalar curvature $R\geq 0$, and the boundary condition reduces to mean convexity $H\geq 0$, allowing us to compare $(\Omega,g)$ with Euclidean polytopes.
+Gromov [Gromov14] initiated the study of scalar curvature comparison for polytopes with nonnegative scalar curvature to explore the notion of scalar curvature lower bounds in low-regularity spaces, and this problem has been extensively investigated; see [BW23, Gromov14, Gromov19, Gromov22, Li17, Li17Err, Li19, Wang-Xie-Yu].
+For $q=g$, the dominant energy condition reduces to the inequality $R\geq-n(n-1)$.
+Li [Li20] established a polyhedral comparison result for metrics with scalar curvature at least n(n1)-n(n-1).
+The case of general initial data sets was addressed by Tsang [Tsang], in the special case when $\Omega$ is a three-dimensional cube.
 We will employ the strategy developed by the second named author in [Brendle], adapting it to initial data sets.
-To achieve this, we utilize a boundary value problem for Dirac operators on smooth domains, which have been extensively studied thanks to the work of Hörmander [Hormander], and the first named author and Ballmann [BB, BB13].
 We write the given polytope in the form $\Omega=\bigcap_{i\in I}\{u_{i}\leq 0\}$, where $I$ is a finite set and $u_{i}$ are linear functions on $\mathbb{R}^{n}$.
@@
 Assume that $n\geq 2$ is an integer.
-Let $\Omega$ be a compact, convex polytope in $\mathbb{R}^{n}$ with nonempty interior.
+Let $\Omega$ be a compact convex polytope in $\mathbb{R}^{n}$ with nonempty interior.
 Let $g$ be a Riemannian metric which is defined on an open set containing $\Omega$.
 Let $q$ be a symmetric $(0,2)$-tensor, which is defined on an open set containing $\Omega$.
-We assume that
-  • $(\Omega,g,q)$ satisfies the dominant energy condition,
-  • $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|$ on the boundary faces of $\Omega$,
+We assume that the following conditions are satisfied:
+  • $(\Omega,g,q)$ satisfies the dominant energy condition.
+  • $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|$ on the boundary faces of $\Omega$.
   • The Matching Angle Hypothesis is satisfied.
@@
 Moreover, each boundary face of $\Omega$ is totally geodesic as a codimension 2 submanifold in the Minkowski spacetime $\mathbb{R}^{n,1}$.
-Conjecture 1.7.
-Under the assumptions of Theorem II, there exists a smooth map $f\colon\Omega\to\mathbb{R}$ satisfying
-  • $|df|<1$ on $\Omega$,
-  • $f=0$ along $\partial\Omega$,
-and an isometry of $(\Omega,g)$ to
-$\textrm{graph}(f)=\{(x,f(x))\colon x\in\Omega\} \subset\mathbb{R}^{n,1}$
-with the induced Riemannian metric. Furthermore, under this isometry, $q$ is the second fundamental form of $\textrm{graph}(f)$ in $\mathbb{R}^{n,1}$.
 2. A boundary value problem for the Dirac operator in odd dimensions
-Throughout this section, we assume that $M$ is a compact, connected spin manifold of dimension $n\geq 2$ with boundary $\partial M=:\Sigma$.
+Throughout this section, we assume that $M$ is a compact connected spin manifold of dimension $n\geq 2$ with boundary $\partial M=\Sigma$.
 Let $g$ be a Riemannian metric on $M$.
-Suppose that $\nu$ is the unit outward normal field along $\partial M$ and that $N\colon\partial M\to S^{n-1}$ is a smooth map.
+Suppose that $\nu$ is the unit outward normal field along $\partial M$ and that $N\colon\partial M\to S^{n-1}$ is a given smooth map.
 Let $m=2^{[\frac{n}{2}]}$.
@@
 Let $S_{0}$ denote the spinor module of the complex Clifford algebra $\mathbb{C}\mathrm{l}(\mathbb{R}^{n})$ of $\mathbb{R}^{n}$.
-Now both $S_{0}$ and its dual space $S_{0}^{*}$ are $m$-dimensional Hermitian vector spaces.
+Both $S_{0}$ and its dual space $S_{0}^{*}$ are $m$-dimensional Hermitian vector spaces.
 We define a complex vector bundle $\mathcal{E}$ over $M$ by $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$.
@@
 The fiber of $\mathcal{E}\to M$ over $x\in M$ is given by
-(2.1) $\mathcal{E}|_{x}=\mathcal{S}|_{x}\otimes S_{0}^{*}=\mathrm{Hom}(S_{0},\mathcal{S}|_{x}).$
+$\mathcal{E}|_{x}=\mathcal{S}|_{x}\otimes S_{0}^{*}=\mathrm{Hom}(S_{0},\mathcal{S}|_{x}).$
 Thus, a section $s$ of $\mathcal{E}$ can be considered as a field of homomorphisms.
-Given an element $\sigma\in S_{0}$ we can apply $s$ at each point $x\in M$ to it and obtain a section $x\mapsto s(x)\sigma$ of $\mathcal{S}$.
-Denote by $\gamma\colon TM\to\mathrm{End}(\mathcal{S})$ the Clifford multiplication of $(M,g)$ and by $\gamma_{0}\colon\mathbb{R}^{n}\to\mathrm{End}(S_{0})$ the one of $\mathbb{R}^{n}$.
-This gives an induced Clifford multiplication $\gamma_{0}^{*}\colon\mathbb{R}^{n}\to\mathrm{End}(S_{0}^{*})$, where $\gamma_{0}^{*}(v)$ is defined as the adjoint of $\gamma_{0}(v)$.
-We define
+Given an element $\sigma\in S_{0}$ we can apply $s$ at each point $x\in M$ to $\sigma$ and obtain a section $s\sigma$ of $\mathcal{S}$.
+We denote by $\gamma\colon TM\to\mathrm{End}(\mathcal{S})$ and $\gamma_{0}\colon\mathbb{R}^{n}\to\mathrm{End}(S_{0})$ the Clifford multiplication of $(M,g)$ and the one of $\mathbb{R}^{n}$, respectively.
+This gives an induced Clifford multiplication
+$\gamma_{0}^{*}\colon\mathbb{R}^{n}\to\mathrm{End}(S_{0}^{*}),$
+where $\gamma_{0}^{*}(\xi)$ is defined as the adjoint of $\gamma_{0}(\xi)$. We define
 (2.1) $\chi\colon\mathcal{E}|_{\Sigma}\to\mathcal{E}|_{\Sigma}\quad\mbox{ by }\quad\chi_{x}:=-\gamma(\nu(x))\otimes\gamma_{0}^{*}(N(x)).$
-Since $\gamma(\nu)^{2}=\gamma_{0}^{*}(N)^{2}=-1$ we have
-$\chi^{2}=1.$
-Since both $\gamma(\nu)$ and $\gamma_{0}^{*}(N)$ are skew-adjoint we get
-$\chi^{*}=\chi.$
-For $\xi\in T\partial M$ the linear map $\gamma(\xi)\otimes 1$ anticommutes with $\chi$ and is an isomorphism if $\xi\neq 0$. Therefore the ±$\pm 1$-eigenspaces of $\chi$ have the same dimension.
-Let $\mathcal{D}$ denote the spinorial Dirac operator on $M$, acting on sections of $\mathcal{E}$.
+Using the identity $\gamma(\nu)^{2}=\gamma_{0}^{*}(N)^{2}=-\textup{id}$, we obtain $\chi^{2}=\textup{id}$. Since both $\gamma(\nu)$ and $\gamma_{0}^{*}(N)$ are skew-adjoint, we have $\chi^{*}=\chi$.
+For each $X\in T\Sigma$, the linear map $\gamma(X)\otimes\textup{id}$ anticommutes with $\chi$ and is an isomorphism if $X\neq 0$. Therefore the $\pm 1$-eigenspaces of $\chi$ have the same dimension.
+Let $\mathcal{D}$ denote the Dirac operator on $M$, acting on sections of $\mathcal{E}$.
 This Dirac operator is formally selfadjoint.
-By Corollary 7.23 in [BB], the conditions $\chi s=s$ and $\chi s=-s$ both form elliptic boundary conditions for $\mathcal{D}$.
-Since $\gamma(\nu)\otimes 1$ commutes with $\chi$, the boundary conditions $\chi s=s$ and $\chi s=-s$ are adjoint to each other, see [BB]*Section 7.2.
+By [Baer-Ballmann]*Corollary 7.23, the conditions $\chi s=s$ and $\chi s=-s$ both form elliptic boundary conditions for $\mathcal{D}$.
+Since $\gamma(\nu)\otimes\textup{id}$ commutes with $\chi$, the boundary conditions $\chi s=s$ and $\chi s=-s$ are adjoint to each other, see [Baer-Ballmann]*Section 7.2.
 We denote by $L^{2}(M,\mathcal{E})$ the Hilbert space of square integrable sections of $\mathcal{E}$ and by $H^{1}(M,\mathcal{E})$ the Sobolev space of $L^{2}$-sections whose distributional derivatives are also square integrable.
-If $\mathcal{F}\subset\mathcal{E}|_{\Sigma}$ is a subbundle of $\mathcal{E}$ along the boundary, we write $H^{1}(M,\mathcal{E},\mathcal{F}):=\{u\in H^{1}(M,\mathcal{E})\mid u|_{\Sigma}\in L^{2}(\Sigma,\mathcal{F})\}$.
-Note that the restriction map extends uniquely to a bounded linear map $H^{1}(M,\mathcal{E})\to L^{2}(\Sigma,\mathcal{E})$ by the trace theorem.
+If $\mathcal{F}\subset\mathcal{E}|_{\Sigma}$ is a subbundle of $\mathcal{E}$ along the boundary, we write
+$H^{1}(M,\mathcal{E},\mathcal{F}):=\{u\in H^{1}(M,\mathcal{E}):u|_{\Sigma}\in L^{2}(\Sigma,\mathcal{F})\}.$
+By the trace theorem, the restriction map extends uniquely to a bounded linear map
+$H^{1}(M,\mathcal{E})\to L^{2}(\Sigma,\mathcal{E}).$
 We decompose the bundle
 (2.2) $\mathcal{E}|_{\Sigma}=\mathcal{F}^{+}\oplus\mathcal{F}^{-}$
-where $\mathcal{F}^{\pm}$ are the eigensubbundles of $\chi$ corresponding to the eigenvalues ±$\pm 1$. We need the following index computation, which generalizes the results in [Brendle].
+where $\mathcal{F}^{\pm}$ are the eigensubbundles of $\chi$ corresponding to the eigenvalues $\pm 1$.
+As usual, we denote by $\nu$ the outward unit normal along $\Sigma$.
+Let $H$ denote the unnormalized mean curvature.
+The sign convention is such that the boundary of a Euclidean ball has positive mean curvature.
+We define an operator $\mathcal{D}^{\Sigma}$ acting on sections of $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes S_{0}^{*}$ by
+$\mathcal{D}^{\Sigma}=\sum_{j=1}^{n-1}(\gamma(\nu)\otimes\textup{id})(\gamma(e_{j})\otimes\textup{id})\,\nabla_{e_{j}}^{\mathcal{E}}+\tfrac{1}{2}H,$
+where $e_{1},\ldots,e_{n-1}$ denotes a local orthonormal frame on $\Sigma$ and $\nabla^{\mathcal{E}}$ denotes the connection on $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$ defined above.
+Moreover, we define an operator $\mathcal{A}$ acting on sections of $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes S_{0}^{*}$ by
+$\mathcal{A}=\mathcal{D}^{\Sigma}+\tfrac{1}{2}\,\chi\circ\Big(\sum_{j=1}^{n-1}\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j}))\Big).$
+Let $\mathcal{S}_{0}$ denote the trivial bundle over $S^{n-1}$ with fiber $S_{0}$, and let $\mathcal{S}_{0}^{*}$ denote the dual bundle over $S^{n-1}$ with fiber $S_{0}^{*}$.
+With this understood, we may write $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes N^{*}\mathcal{S}_{0}^{*}$.
+By [Baer96]*Proposition 2.2, $\mathcal{D}^{\Sigma}$ can be identified with a twisted Dirac operator acting on sections of
+$\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes N^{*}\mathcal{S}_{0}^{*},$
+where the twist bundle $N^{*}\mathcal{S}_{0}^{*}$ carries the flat connection.
+Lemma 2.2.
+Along $\Sigma$ we have
+$\mathcal{D}^{\Sigma}\chi+\chi\mathcal{D}^{\Sigma}=-\sum_{j=1}^{n-1}\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j})),$
+where $e_{1},\dots,e_{n-1}$ is a local orthonormal tangent frame to $\Sigma$.
+Proof.
+Recall that $\mathcal{D}^{\Sigma}$ anticommutes with $\gamma(\nu)\otimes\textup{id}$. Hence, we compute for a smooth section $s$ of $\mathcal{E}|_{\Sigma}$:
+$\mathcal{D}^{\Sigma}\chi s$ $=-\mathcal{D}^{\Sigma}(\gamma(\nu)\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(N))s$
+$=(\gamma(\nu)\otimes\textup{id})\mathcal{D}^{\Sigma}(\textup{id}\otimes\gamma_{0}^{*}(N))s$
+$=-\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\textup{id})\,\nabla_{e_{j}}^{\mathcal{E}}((\textup{id}\otimes\gamma_{0}^{*}(N))s)+\tfrac{1}{2}H\,(\gamma(\nu)\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(N))s$
+$=-\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(N))\,\nabla_{e_{j}}^{\mathcal{E}}s+\tfrac{1}{2}H\,(\gamma(\nu)\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(N))s$
+$\quad-\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(dN(e_{j})))s$
+$=(\gamma(\nu)\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(N))\,\mathcal{D}^{\Sigma}s-\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(dN(e_{j})))s$
+$=-\chi\mathcal{D}^{\Sigma}s-\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\textup{id})(\textup{id}\otimes\gamma_{0}^{*}(dN(e_{j})))s.\qed$
+Lemma 2.3.
+Along $\Sigma$ we have $\mathcal{A}\chi+\chi\mathcal{A}=0$.
+Proof.
+This follows from Lemma 2.2. ∎
+The following index computation generalizes the results in [Brendle] and plays a key role in our arguments.
 Proposition 2.4.
-Let $M$ have odd dimension $n\geq 3$. Then the operator
+Suppose that $M$ has odd dimension $n\geq 3$. Then the operator
 $\mathcal{D}\colon H^{1}(M,\mathcal{E},\mathcal{F}^{+})\to L^{2}(M,\mathcal{E})$
@@
 Proof.
-Since $n$ is odd, the restriction $\mathcal{S}|_{\Sigma}$ can be canonically identified with the spinor bundle of $\Sigma$. The field of involutions $i\gamma(\nu)$ can be used to decompose
-$\mathcal{S}|_{\Sigma}=\mathcal{S}_{\Sigma}^{+}\oplus\mathcal{S}_{\Sigma}^{-}$
-where $\mathcal{S}_{\Sigma}^{\pm}$ are the eigensubbundles corresponding to the eigenvalues ±$\pm 1$.
-Let $\mathcal{S}_{0}$ denote the trivial bundle over $S^{n-1}$ with fiber $S_{0}$, and let $\mathcal{S}_{0}^{*}$ denote the dual bundle over $S^{n-1}$ with fiber $S_{0}^{*}$.
-With this understood, we may write $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes N^{*}\mathcal{S}_{0}^{*}$.
-We may decompose the bundle $\mathcal{S}_{0}^{*}$ as
+Note that $\mathcal{D}$ and $\mathcal{A}$ are formally self-adjoint, and $\mathcal{A}$ is an adapted boundary operator for $\mathcal{D}$.
+It follows from Lemma 2.3 that $\mathcal{A}$ is an odd operator in the sense that it interchanges the bundles $\mathcal{F}^{+}$ and $\mathcal{F}^{-}$.
+Moreover, the involution $i\gamma(\nu)\otimes\textup{id}$ preserves the splitting (2.2) and anticommutes with $\mathcal{A}$.
+Therefore, the holographic index theorem [Baer-Brendle-Hanke-Wang]*Theorem B.1 applies and tells us that the operator
+$\mathcal{D}\colon H^{1}(M,\mathcal{E},\mathcal{F}^{+})\to L^{2}(M,\mathcal{E})$
+is Fredholm and
+$\mathsf{ind}(\mathcal{D}\colon H^{1}(M,\mathcal{E},\mathcal{F}^{+})\to L^{2}(M,\mathcal{E}))=\tfrac{1}{2}\,\mathsf{ind}(\mathcal{A}\colon H^{1}(\Sigma,\mathcal{F}^{+})\to L^{2}(\Sigma,\mathcal{F}^{-})).$
+It remains to compute the index of the operator $\mathcal{A}\colon H^{1}(\Sigma,\mathcal{F}^{+})\to L^{2}(\Sigma,\mathcal{F}^{-})$.
+Since $n$ is odd, the restriction $\mathcal{S}|_{\Sigma}$ can be canonically identified with the spinor bundle of $\Sigma$.
+The field of involutions $i\gamma(\nu)$ gives a decomposition $\mathcal{S}|_{\Sigma}=\mathcal{S}_{\Sigma}^{+}\oplus\mathcal{S}_{\Sigma}^{-}$, where $\mathcal{S}_{\Sigma}^{\pm}$ are the eigensubbundles corresponding to the eigenvalues $\pm 1$.
+Moreover, we may decompose the bundle $\mathcal{S}_{0}^{*}$ as
 (2.3) $\mathcal{S}_{0}^{*}=\mathcal{S}_{0}^{*,+}\oplus\mathcal{S}_{0}^{*,-},$
-where, for each point $v\in S^{n-1}$, $\mathcal{S}_{0}^{*,\pm}$ denote the eigensubbundles of $i\gamma(v)$ corresponding to the eigenvalues ±$\pm 1$. This gives a decomposition
-(2.5) $N^{*}\mathcal{S}^{*}=N^{*}\mathcal{S}_{0}^{*,+}\oplus N^{*}\mathcal{S}_{0}^{*,-},$
-where $N^{*}\mathcal{S}_{0}^{*,\pm}$ are the eigensubbundles of $i\gamma_{0}^{*}(N)$ corresponding to the eigenvalues ±$\pm 1$. We then have
+where, for each point $v\in S^{n-1}$, $\mathcal{S}_{0}^{*,\pm}$ denote the eigensubbundles of $i\gamma(v)$ corresponding to the eigenvalues $\pm 1$. This gives a decomposition
+$N^{*}\mathcal{S}_{0}^{*}=N^{*}\mathcal{S}_{0}^{*,+}\oplus N^{*}\mathcal{S}_{0}^{*,-},$
+where $N^{*}\mathcal{S}_{0}^{*,\pm}$ are the eigensubbundles of $i\gamma_{0}^{*}(N)$ corresponding to the eigenvalues $\pm 1$. We then have
 $\mathcal{F}^{+}$ $=(\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,+})\oplus(\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,-}),$
@@
 We equip $\mathcal{S}_{0}^{*,\pm}$ with these latter connections and $N^{*}\mathcal{S}_{0}^{*,\pm}$ with the corresponding pull-back connections.
-We next consider the twisted Dirac operators
+Recall that $\mathcal{A}$ maps sections of
+$\mathcal{F}^{+}=(\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,+})\oplus(\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,-})$
+to sections of $\mathcal{F}^{-}=(\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,-})\oplus(\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,+})$.
+Since the involution $i\gamma(\nu)\otimes\textup{id}$ preserves the splitting (2.2) and anticommutes with $\mathcal{A}$, it follows that $\mathcal{A}$ maps sections of $\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,+}$ to sections of $\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,+}$, and $\mathcal{A}$ maps sections of $\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,-}$ to sections of $\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,-}$.
+Therefore, the operator $\mathcal{A}\colon\mathcal{F}^{+}\to\mathcal{F}^{-}$ can be written in the form
+$\begin{bmatrix}0&\mathcal{D}^{\Sigma}_{-}\\ \mathcal{D}^{\Sigma}_{+}&0\end{bmatrix}$,
+where
 $\mathcal{D}^{\Sigma}_{+}$ $\colon H^{1}(\Sigma,\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,+})\to L^{2}(\Sigma,\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,+}),$
-$\mathcal{D}^{\Sigma}_{-}$ $\colon H^{1}(\Sigma,\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,-})\to L^{2}(\Sigma,\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,-}).$
-Then the operator
-(2.6) $\mathcal{D}^{\Sigma}=\begin{pmatrix}0&\mathcal{D}^{\Sigma}_{-}\\ \mathcal{D}^{\Sigma}_{+}&0\end{pmatrix}$
-maps sections of $\mathcal{F}^{+}$ to sections of $\mathcal{F}^{-}$.
-In the following, $\hat{\mathsf{A}}(\Sigma)$ denotes the $\hat{\mathsf{A}}$-form of the tangent bundle of $\Sigma$ and $\mathsf{ch}(\mathcal{S}_{0}^{*,\pm})$ the Chern character form of the bundle $\mathcal{S}_{0}^{*,\pm}$.
-The lower index $k$ in $\hat{\mathsf{A}}(\Sigma)_{k}$ indicates the homogeneous part of degree $k$ and similarly for the Chern character.
-By the Atiyah-Singer index theorem, we have
+$\mathcal{D}^{\Sigma}_{-}$ $\colon H^{1}(\Sigma,\mathcal{S}_{\Sigma}^{-}\otimes N^{*}\mathcal{S}_{0}^{*,-})\to L^{2}(\Sigma,\mathcal{S}_{\Sigma}^{+}\otimes N^{*}\mathcal{S}_{0}^{*,-})$
+are Dirac-type operators. Thus,
+$\mathsf{ind}(\mathcal{A}\colon H^{1}(\Sigma,\mathcal{F}^{+})\to L^{2}(\Sigma,\mathcal{F}^{-}))=\mathsf{ind}(\mathcal{D}^{\Sigma}_{+})+\mathsf{ind}(\mathcal{D}^{\Sigma}_{-}).$
+We may compute $\mathsf{ind}(\mathcal{D}^{\Sigma}_{+})$ and $\mathsf{ind}(\mathcal{D}^{\Sigma}_{-})$ using the Atiyah-Singer index theorem.
+To that end, we denote by $\hat{\mathsf{A}}(\Sigma)$ the $\hat{\mathsf{A}}-$form of the tangent bundle of $\Sigma$, and by $\mathsf{ch}(\mathcal{S}_{0}^{*,\pm})$ the Chern character form of the bundle $\mathcal{S}_{0}^{*,\pm}$.
+The lower index $k$ in $\hat{\mathsf{A}}(\Sigma)_{k}$ indicates the homogeneous part of degree $k$.
+We use the analogous notation for the homogeneous parts of the Chern character.
+The Atiyah-Singer index theorem gives
 $\mathsf{ind}(\mathcal{D}^{\Sigma}_{+})$ $=\int_{\Sigma}\hat{\mathsf{A}}(\Sigma)\wedge N^{*}\mathsf{ch}(\mathcal{S}_{0}^{*,+})$
@@
 $=\mathsf{deg}(N)\int_{S^{n-1}}\mathsf{ch}(\mathcal{S}_{0}^{*,+})_{n-1}+\frac{m}{2}\int_{\Sigma}\hat{\mathsf{A}}(\Sigma)_{n-1}$
-and similarly
+and, similarly,
 $-\mathsf{ind}(\mathcal{D}^{\Sigma}_{-})$ $=\mathsf{deg}(N)\int_{S^{n-1}}\mathsf{ch}(\mathcal{S}_{0}^{*,-})_{n-1}+\frac{m}{2}\int_{\Sigma}\hat{\mathsf{A}}(\Sigma)_{n-1}.$
 Therefore,
-$\mathsf{ind}(\mathcal{D}^{\Sigma})$ $=\mathsf{ind}(\mathcal{D}^{\Sigma}_{+})+\mathsf{ind}(\mathcal{D}^{\Sigma}_{-})$
+$\mathsf{ind}(\mathcal{A}\colon$ $H^{1}(\Sigma,\mathcal{F}^{+})\to L^{2}(\Sigma,\mathcal{F}^{-}))$
+$=\mathsf{ind}(\mathcal{D}^{\Sigma}_{+})+\mathsf{ind}(\mathcal{D}^{\Sigma}_{-})$
 $=\mathsf{deg}(N)\int_{S^{n-1}}(\mathsf{ch}(\mathcal{S}_{0}^{*,+})-\mathsf{ch}(\mathcal{S}_{0}^{*,-}))_{n-1}$
 $=2\,\mathsf{deg}(N).$
-Now, we observe that $\mathcal{D}$ and $\mathcal{D}^{\Sigma}$ are formally selfadjoint, $\mathcal{D}^{\Sigma}$ is an adapted boundary operator for $\mathcal{D}$, the involution $i\gamma(\nu)\otimes 1$ preserves the splitting (2.3) and anticommutes with $\mathcal{D}^{\Sigma}$.
-Therefore, the holographic index theorem from Appendix B in [BBHW23] applies and tells us that the operator
-$\mathcal{D}\colon H^{1}(M,\mathcal{E},\mathcal{F}^{+})\to L^{2}(M,\mathcal{E})$
-is Fredholm and
-$\mathsf{ind}(\mathcal{D}\colon H^{1}(M,\mathcal{E},\mathcal{F}^{+})\to L^{2}(M,\mathcal{E}))=\tfrac{1}{2}\,\mathsf{ind}(\mathcal{D}^{\Sigma})=\mathsf{deg}(N).\qed$
+Putting these facts together, the assertion follows. ∎
 3. Spin geometry for initial data sets
@@
 Let $\chi$ be defined as in (2.1).
-We will denote the selfadjoint endomorphism field corresponding to $q$ also by $q\colon TM\to TM$, i.e.
+We will denote the self-adjoint endomorphism field corresponding to $q$ also by $q\colon TM\to TM$, i.e.
 $q(X,Y)=g(q(X),Y)$.
@@
 and
-(3.2) $J:={\mathrm{div}}(q)-\nabla\mathrm{tr}(q)$
-where $R$ is the scalar curvature of $g$. We use $q$ to modify the connections on $\mathcal{S}$ and on $\mathcal{E}$.
+(3.2) $J:=\text{\rm div}(q)-\nabla\mathrm{tr}(q)$
+where $R$ is the scalar curvature of $g$. In the first step, we use $q$ to define modified connections on $\mathcal{S}$ and on $\mathcal{E}$.
 Definition 3.1.
 We define a connection $\nabla^{\mathcal{S},q}$ on $\mathcal{S}$ by
-(3.3) $\nabla^{\mathcal{S},q}:=\nabla^{\mathcal{S}}+\frac{i}{2}\gamma\circ q$
-on $\mathcal{S}$. We next introduce the homomorphism field $Q\colon\mathcal{E}\to T^{*}M\otimes\mathcal{E}$ which maps spinors to spinor-valued $1$-forms by setting
-(3.4) $(Qs)(X):=(\gamma(q(X))\otimes 1)s$
+$\nabla_{X}^{\mathcal{S},q}:=\nabla_{X}^{\mathcal{S}}+\tfrac{i}{2}\gamma(q(X))$
+on $\mathcal{S}$. We next define a homomorphism field $Q\colon\mathcal{E}\to T^{*}M\otimes\mathcal{E}$ by setting
+$(Qs)(X):=(\gamma(q(X))\otimes\textup{id})s$
 for every section $s$ of $\mathcal{E}$ and every tangent vector $X$. With this understood, we define a connection $\nabla^{\mathcal{E},q}$ on $\mathcal{E}$ by
-(3.5) $\nabla^{\mathcal{E},q}:=\nabla^{\mathcal{E}}+\tfrac{i}{2}Q.$
-Note that the connection $\nabla^{\mathcal{E},q}$ on $\mathcal{E}$ is the tensor product connection of $\nabla^{\mathcal{S},q}$ with the flat connection on the trivial bundle over $M$ with fiber $S_{0}$.
+$\nabla_{X}^{\mathcal{E},q}s:=\nabla_{X}^{\mathcal{E}}s+\tfrac{i}{2}(Qs)(X).$
+Note that the connection $\nabla^{\mathcal{E},q}$ on $\mathcal{E}$ is the tensor product connection of $\nabla^{\mathcal{S},q}$ with the flat connection on the trivial bundle over $M$ with fiber $S_{0}^{*}$.
 In order to derive a modified Weitzenböck formula involving $\nabla^{\mathcal{E},q}$ we need some preparation.
 Lemma 3.2.
-Denote the (pointwise) adjoint of $Q$ by $Q^{*}\colon T^{*}M\otimes\mathcal{E}\to\mathcal{E}$. Then we have
+Let $Q^{*}\colon T^{*}M\otimes\mathcal{E}\to\mathcal{E}$ denote the pointwise adjoint of $Q$. Then the following statements hold:
   1. (a)
@@
   2. (b)
-    $Q^{*}\nabla^{\mathcal{E}}-(\nabla^{\mathcal{E}})^{*}Q=\gamma(\mathrm{div}(q))\otimes 1$.
+    $Q^{*}\nabla^{\mathcal{E}}-(\nabla^{\mathcal{E}})^{*}Q=\gamma(\text{\rm div}(q))\otimes\textup{id}$.
 Proof.
 Let $p\in M$ and let $e_{1},\dots,e_{n}$ be an orthonormal basis of $T_{p}M$.
-Denote the dual basis of $T^{*}_{p}M$ by $e^{1},\dots,e^{n}$.
-Then we have for $s\in\mathcal{E}|_{p}$ and $\psi\in T^{*}_{p}M\otimes\mathcal{E}|_{p}$:
-$\langle Q^{*}\psi,s\rangle$ $=\langle\psi,Qs\rangle=\Big{\langle}\psi,\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes 1)s\Big{\rangle}$
-$=\sum_{j}\langle\psi(e_{j}),(\gamma(q(e_{j}))\otimes 1)s\rangle$
-$=-\sum_{j}\langle(\gamma(q(e_{j}))\otimes 1)\psi(e_{j}),s\rangle$
-and hence
-$Q^{*}\psi=-\sum_{j}(\gamma(q(e_{j}))\otimes 1)\psi(e_{j}).$
+We denote the dual basis of $T^{*}_{p}M$ by $e^{1},\dots,e^{n}$.
+For $s\in\mathcal{E}|_{p}$ and $\psi\in T^{*}_{p}M\otimes\mathcal{E}|_{p}$, we compute
+$\langle Q^{*}\psi,s\rangle$ $=\langle\psi,Qs\rangle=\Big\langle\psi,\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes\textup{id})s\Big\rangle$
+$=\sum_{j}\langle\psi(e_{j}),(\gamma(q(e_{j}))\otimes\textup{id})s\rangle$
+$=-\sum_{j}\langle(\gamma(q(e_{j}))\otimes\textup{id})\psi(e_{j}),s\rangle.$
 Therefore,
-$Q^{*}Qs$ $=Q^{*}\Big{(}\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes 1)s\Big{)}$
-$=-\sum_{j}(\gamma(q(e_{j}))\otimes 1)(\gamma(q(e_{j}))\otimes 1)s$
+$Q^{*}\psi=-\sum_{j}(\gamma(q(e_{j}))\otimes\textup{id})\psi(e_{j}).$
+Consequently,
+$Q^{*}Qs$ $=Q^{*}\Big(\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes\textup{id})s\Big)$
+$=-\sum_{j}(\gamma(q(e_{j}))\otimes\textup{id})(\gamma(q(e_{j}))\otimes\textup{id})s$
 $=\sum_{j}|q(e_{j})|^{2}s=|q|^{2}s.$
 This proves (a).
-To show (b), extend the orthonormal basis to a neighborhood of $p$ such that it is synchronous at $p$.
-Let $s$ be a smooth section of $\mathcal{E}$, defined near $p$.
+To show (b), we extend the orthonormal basis to a neighborhood of $p$ such that $\nabla e_{j}=0$ at $p$ for $j=1,\ldots,n$.
+Let $s$ be a smooth section of $\mathcal{E}$, defined near pp.
 Then we compute at $p$:
-$(\nabla^{\mathcal{E}})^{*}Qs$ $=(\nabla^{\mathcal{E}})^{*}\Big{(}\sum_{j}e^{j}\otimes(\gamma(q(e_{j}))\otimes 1)s\Big{)}$
-$=-\sum_{j}(\gamma((D_{e_{j}}q)(e_{j}))\otimes 1)s-\sum_{j}(\gamma(q(e_{j}))\otimes 1)\nabla_{e_{j}}^{\mathcal{E}}s$
-$=-(\gamma(\mathrm{div}(q))\otimes 1)s+Q^{*}\nabla^{\mathcal{E}}s.\qed$
-The well-known Weitzenböck formula gives
-$\mathcal{D}^{2}=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}+\tfrac{1}{4}R.$
-We find a modification, taking $q$ into account.
+$(\nabla^{\mathcal{E}})^{*}Qs$ $=-\sum_{j}\nabla_{e_{j}}^{\mathcal{E}}\big((\gamma(q(e_{j}))\otimes\textup{id})s\big)$
+$=-\sum_{j}(\gamma(q(e_{j}))\otimes\textup{id})\nabla_{e_{j}}^{\mathcal{E}}s-\sum_{j}(\gamma((D_{e_{j}}q)(e_{j}))\otimes\textup{id})s$
+$=Q^{*}\nabla^{\mathcal{E}}s-(\gamma(\text{\rm div}(q))\otimes\textup{id})s.\qed$
+The well-known Weitzenböck formula gives $\mathcal{D}^{2}=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}+\tfrac{1}{4}R$. We find a modification, taking $q$ into account.
 Proposition 3.3 (Modified Weitzenböck formula).
 We have the following operator identity:
-$(\mathcal{D}+\tfrac{i}{2}\mathrm{tr}(q))(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))=(\nabla^{\mathcal{E},q})^{*}\nabla^{\mathcal{E},q}+\tfrac{1}{2}\mu+\tfrac{i}{2}\gamma(J)\otimes 1.$
+$(\mathcal{D}+\tfrac{i}{2}\mathrm{tr}(q))(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))=(\nabla^{\mathcal{E},q})^{*}\nabla^{\mathcal{E},q}+\tfrac{1}{2}\mu+\tfrac{i}{2}\gamma(J)\otimes\textup{id}.$
 Proof.
@@
 $=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}-\tfrac{i}{2}(Q^{*}\nabla^{\mathcal{E}}-(\nabla^{\mathcal{E}})^{*}Q)+\tfrac{1}{4}Q^{*}Q$
-$=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}-\tfrac{i}{2}\gamma(\mathrm{div}(q))\otimes 1+\tfrac{1}{4}|q|^{2}.$
+$=(\nabla^{\mathcal{E}})^{*}\nabla^{\mathcal{E}}-\tfrac{i}{2}\gamma(\text{\rm div}(q))\otimes\textup{id}+\tfrac{1}{4}|q|^{2}.$
 Furthermore,
 $(\mathcal{D}+\tfrac{i}{2}\mathrm{tr}(q))(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))$ $=\mathcal{D}^{2}-\tfrac{i}{2}[\mathcal{D},\mathrm{tr}(q)]+\tfrac{1}{4}\mathrm{tr}(q)^{2}$
-$=\mathcal{D}^{2}-\tfrac{i}{2}\gamma(\nabla\mathrm{tr}(q))\otimes 1+\tfrac{1}{4}\mathrm{tr}(q)^{2}.$
+$=\mathcal{D}^{2}-\tfrac{i}{2}\gamma(\nabla\mathrm{tr}(q))\otimes\textup{id}+\tfrac{1}{4}\mathrm{tr}(q)^{2}.$
 These two equations together with the standard Weitzenböck formula yield
 $(\mathcal{D}+\tfrac{i}{2}\mathrm{tr}(q))(\mathcal{D}$ $-\tfrac{i}{2}\mathrm{tr}(q))-(\nabla^{\mathcal{E},q})^{*}\nabla^{\mathcal{E},q}$
-$=\tfrac{1}{4}R+\tfrac{i}{2}\gamma(\mathrm{div}(q)-\nabla\mathrm{tr}(q))\otimes 1+\tfrac{1}{4}(\mathrm{tr}(q)^{2}-|q|^{2})$
-$=\tfrac{1}{2}\mu+\tfrac{i}{2}\gamma(J)\otimes 1.\qed$
-Again we denote $\partial M=\Sigma$ and the unit outward normal along $\Sigma$ by $\nu$.
-Let $H$ denote the (unnormalized) mean curvature.
-The sign convention is such that the boundary of a Euclidean ball has positive mean curvature.
-We define an operator $\mathcal{D}^{\Sigma}_{0}$ acting on sections of $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes S_{0}^{*}$ by
-$\mathcal{D}^{\Sigma}_{0}=-\sum_{j=1}^{n-1}(\gamma(\nu)\otimes 1)(\gamma(e_{j})\otimes 1)\,\nabla_{e_{j}}^{\mathcal{E}}-\tfrac{1}{2}H,$
-where $e_{1},\ldots,e_{n-1}$ denotes a local orthonormal frame on $\Sigma$ and $\nabla^{\mathcal{E}}$ denotes the connection on $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$ defined above.
-By Proposition 2.2 in [B96], $\mathcal{D}^{\Sigma}_{0}$ can be identified with a twisted Dirac operator acting on sections of $\mathcal{E}|_{\Sigma}=\mathcal{S}|_{\Sigma}\otimes N^{*}\mathcal{S}_{0}^{*}$, where the twist bundle $N^{*}\mathcal{S}_{0}^{*}$ carries the flat connection.
-Note that $\mathcal{D}^{\Sigma}_{0}$ differs from the operator $\mathcal{D}^{\Sigma}$ defined in (2.6).
-In the definition of $\mathcal{D}^{\Sigma}$, we used a nonflat connection on the twist bundle $N^{*}\mathcal{S}_{0}^{*}$ that preserves the splittting (2.5).
-The operators $\mathcal{D}^{\Sigma}$ and $\mathcal{D}^{\Sigma}_{0}$ both anticommute with $\gamma(\nu)\otimes 1$.
+$=\tfrac{1}{4}R+\tfrac{i}{2}\gamma(\text{\rm div}(q)-\nabla\mathrm{tr}(q))\otimes\textup{id}+\tfrac{1}{4}(\mathrm{tr}(q)^{2}-|q|^{2})$
+$=\tfrac{1}{2}\mu+\tfrac{i}{2}\gamma(J)\otimes\textup{id}.\qed$
 Corollary 3.4.
 For all $s\in H^{1}(M,\mathcal{E})$ we have
-$-\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}$ $+\int_{M}|\nabla^{\mathcal{E},q}s|^{2}+\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes 1)s,s\rangle$
-$=\int_{\Sigma}\langle(-\mathcal{D}^{\Sigma}_{0}-\tfrac{1}{2}H+\tfrac{i}{2}(\gamma(q(\nu)-\mathrm{tr}(q)\nu)\otimes 1))s,s\rangle.$
+$-\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}$ $+\int_{M}|\nabla^{\mathcal{E},q}s|^{2}+\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes\textup{id})s,s\rangle$
+$=\int_{\Sigma}\langle(\mathcal{D}^{\Sigma}-\tfrac{1}{2}H+\tfrac{i}{2}(\gamma(q(\nu)-\mathrm{tr}(q)\nu)\otimes\textup{id}))s,s\rangle.$
 Proof.
 Note that
-$(\gamma(\nu)\otimes 1)\mathcal{D}s=-\mathcal{D}^{\Sigma}_{0}s-\nabla_{\nu}^{\mathcal{E}}s-\tfrac{1}{2}Hs$
-along $\Sigma$. Using this and integrating by parts we find
+$(\gamma(\nu)\otimes\textup{id})\mathcal{D}s=\mathcal{D}^{\Sigma}s-\nabla_{\nu}^{\mathcal{E}}s-\tfrac{1}{2}Hs$
+along $\Sigma$. Using this identity and integration by parts we find
 $\int_{M}$ $\langle(\mathcal{D}+\tfrac{i}{2}\mathrm{tr}(q))(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s,s\rangle$
-$=\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{\Sigma}\langle(\gamma(\nu)\otimes 1)(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s,s\rangle$
-$=\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{\Sigma}\big{\langle}\big{(}-\mathcal{D}^{\Sigma}_{0}-\nabla_{\nu}^{\mathcal{E}}-\tfrac{1}{2}H-\tfrac{i}{2}\mathrm{tr}(q)(\gamma(\nu)\otimes 1)\big{)}s,s\big{\rangle}.$
+$=\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{\Sigma}\langle(\gamma(\nu)\otimes\textup{id})(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s,s\rangle$
+$=\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{\Sigma}\big\langle\big(\mathcal{D}^{\Sigma}-\nabla_{\nu}^{\mathcal{E}}-\tfrac{1}{2}H-\tfrac{i}{2}\mathrm{tr}(q)(\gamma(\nu)\otimes\textup{id})\big)s,s\big\rangle.$
 Similarly,
 $\int_{M}\langle(\nabla^{\mathcal{E},q})^{*}\nabla^{\mathcal{E},q}s,s\rangle$ $=\int_{M}|\nabla^{\mathcal{E},q}s|^{2}-\int_{\Sigma}\langle\nabla_{\nu}^{\mathcal{E},q}s,s\rangle$
-$=\int_{M}|\nabla^{\mathcal{E},q}s|^{2}-\int_{\Sigma}\langle(\nabla_{\nu}^{\mathcal{E}}+\tfrac{i}{2}\gamma(q(\nu))\otimes 1)s,s\rangle.$
-Inserting these two equations into the modified Weitzenböck formula concludes the proof. ∎
-Lemma 3.5.
-Along $\Sigma$ we have
-$\mathcal{D}^{\Sigma}_{0}\chi+\chi\mathcal{D}^{\Sigma}_{0}=\sum_{j=1}^{n-1}\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j}))$
-where $e_{1},\dots,e_{n-1}$ is a local orthonormal tangent frame to $\Sigma$.
-Proof.
-Recall that $\mathcal{D}^{\Sigma}_{0}$ anticommutes with $\gamma(\nu)\otimes 1$. Hence, we compute for a smooth section $s$ of $\mathcal{E}|_{\Sigma}$:
-$\mathcal{D}^{\Sigma}_{0}\chi s$ $=-\mathcal{D}^{\Sigma}_{0}(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))s$
-$=(\gamma(\nu)\otimes 1)\mathcal{D}^{\Sigma}_{0}(1\otimes\gamma_{0}^{*}(N))s$
-$=\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)\,\nabla_{e_{j}}^{\mathcal{E}}((1\otimes\gamma_{0}^{*}(N))s)-\tfrac{1}{2}H\,(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))s$
-$=\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(N))\,\nabla_{e_{j}}^{\mathcal{E}}s-\tfrac{1}{2}H\,(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))s$
-$\quad+\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(dN(e_{j})))s$
-$=(\gamma(\nu)\otimes 1)(1\otimes\gamma_{0}^{*}(N))\,\mathcal{D}^{\Sigma}_{0}s+\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(dN(e_{j})))s$
-$=-\chi\mathcal{D}^{\Sigma}_{0}s+\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes 1)(1\otimes\gamma_{0}^{*}(dN(e_{j})))s.\qed$
+$=\int_{M}|\nabla^{\mathcal{E},q}s|^{2}-\int_{\Sigma}\langle(\nabla_{\nu}^{\mathcal{E}}+\tfrac{i}{2}\gamma(q(\nu))\otimes\textup{id})s,s\rangle.$
+Substituting these equations into the modified Weitzenböck formula concludes the proof. ∎
 Lemma 3.5.
@@
 Then we have the estimate
-$-\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{M}$ $|\nabla^{\mathcal{E},q}s|^{2}+\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes 1)s,s\rangle$
-$\leq\tfrac{1}{2}\int_{\Sigma}(\|dN\|_{\mathrm{tr}}-H+|q(\nu,\nu)-\mathrm{tr}(q)|)|s|^{2}.$
+$-\int_{M}|(\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q))s|^{2}+\int_{M}$ $|\nabla^{\mathcal{E},q}s|^{2}+\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes\textup{id})s,s\rangle$
+$\leq\tfrac{1}{2}\int_{\Sigma}(|q(\nu,\nu)-\mathrm{tr}(q)|+\|dN\|_{\mathrm{tr}}-H)|s|^{2}.$
 Proof.
 We first observe that
-$2\,\langle\mathcal{D}^{\Sigma}_{0}s,s\rangle=\varepsilon\,\langle\mathcal{D}^{\Sigma}_{0}\chi s,s\rangle+\varepsilon\,\langle\mathcal{D}^{\Sigma}_{0}s,\chi s\rangle=\varepsilon\,\langle\mathcal{D}^{\Sigma}_{0}\chi s+\chi\mathcal{D}^{\Sigma}_{0}s,s\rangle$
+$2\,\langle\mathcal{D}^{\Sigma}s,s\rangle=\varepsilon\,\langle\mathcal{D}^{\Sigma}\chi s,s\rangle+\varepsilon\,\langle\mathcal{D}^{\Sigma}s,\chi s\rangle=\varepsilon\,\langle\mathcal{D}^{\Sigma}\chi s+\chi\mathcal{D}^{\Sigma}s,s\rangle$
 at each point on $\Sigma$. Using Lemma 2.2, we obtain
-$2\,|\langle\mathcal{D}^{\Sigma}_{0}s,s\rangle|=\Big{|}\Big{\langle}\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j})))s,s\Big{\rangle}\Big{|}$
+$2\,|\langle\mathcal{D}^{\Sigma}s,s\rangle|=\Big|\Big\langle\sum_{j=1}^{n-1}(\gamma(e_{j})\otimes\gamma_{0}^{*}(dN(e_{j})))s,s\Big\rangle\Big|$
 at each point on $\Sigma$. Lemma 3.5 implies
-(3.6) $2\,|\langle\mathcal{D}^{\Sigma}_{0}s,s\rangle|\leq\|dN\|_{\mathrm{tr}}\,|s|^{2}$
+(3.3) $2\,|\langle\mathcal{D}^{\Sigma}s,s\rangle|\leq\|dN\|_{\mathrm{tr}}\,|s|^{2}$
 at each point on $\Sigma$. Moreover, for each $X\in T\Sigma$ we have
-$\langle(\gamma(X)\otimes 1)s,s\rangle$ $=\varepsilon\langle(\gamma(X)\otimes 1)\chi s,s\rangle$
-$=-\varepsilon\langle\chi(\gamma(X)\otimes 1)s,s\rangle$
-$=-\varepsilon\langle(\gamma(X)\otimes 1)s,\chi s\rangle$
-$=-\langle(\gamma(X)\otimes 1)s,s\rangle,$
-hence
-$\langle(\gamma(X)\otimes 1)s,s\rangle=0.$
+$2\,\langle(\gamma(X)\otimes\textup{id})s,s\rangle$ $=\varepsilon\,\langle(\gamma(X)\otimes\textup{id})\chi s,s\rangle+\varepsilon\,\langle(\gamma(X)\otimes\textup{id})s,\chi s\rangle$
+$=\varepsilon\,\langle(\gamma(X)\otimes\textup{id})\chi s+\chi(\gamma(X)\otimes\textup{id})s,s\rangle$
+$=0.$
 Putting $X=q(\nu)^{\text{\rm tan}}$, the part of $q(\nu)$ tangential to $\Sigma$, gives
-$\langle(\gamma(q(\nu))\otimes 1)s,s\rangle=q(\nu,\nu)\,\langle(\gamma(\nu)\otimes 1)s,s\rangle.$
+$\langle(\gamma(q(\nu))\otimes\textup{id})s,s\rangle=q(\nu,\nu)\,\langle(\gamma(\nu)\otimes\textup{id})s,s\rangle.$
 We conclude
-$\big{|}\langle(\gamma(q(\nu)-\mathrm{tr}(q)\nu)\otimes 1)s,s\rangle\big{|}$ $=\big{|}(q(\nu,\nu)-\mathrm{tr}(q))\,\langle(\gamma(\nu)\otimes 1)s,s\rangle\big{|}$
+$\big|\langle(\gamma(q(\nu)-\mathrm{tr}(q)\nu)\otimes\textup{id})s,s\rangle\big|$ $=\big|(q(\nu,\nu)-\mathrm{tr}(q))\,\langle(\gamma(\nu)\otimes\textup{id})s,s\rangle\big|$
 (3.4) $\leq|q(\nu,\nu)-\mathrm{tr}(q)|\,|s|^{2}.$
-Inserting (3.6) and (3.7) into Corollary 3.4 concludes the proof. ∎
+Using (3.3) and (3.4) together with Corollary 3.4, the assertion follows. ∎
 Proposition 3.7.
-Suppose that $M$ has odd dimension $n\geq 3$ and that the map $N\colon\Sigma\to S^{n-1}$ has positive degree. Then there exists a non-trivial section $s\in C^{\infty}(M,\mathcal{E})$ such that
-$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}$
-$\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
-Moreover, $\chi s=s$ holds at each point on the boundary $\Sigma$.
+Suppose that $M$ has odd dimension $n\geq 3$ and the map $N\colon\Sigma\to S^{n-1}$ has positive degree. Then there exists a non-trivial section $s\in C^{\infty}(M,\mathcal{E})$ such that
+$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(|J|-\mu)|s|^{2}+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
+Moreover, $\chi s=s$ at each point on the boundary $\Sigma$.
 Proof.
 By Proposition 2.4, the operator $\mathcal{D}$ subject to the boundary conditions $\chi s=s$ has positive index.
-By deformation invariance of the Fredholm index, the same holds for the operator $\mathcal{D}-\frac{i}{2}\mathrm{tr}(q)$.
-Hence we can find a non-trivial section $s$ of $\mathcal{E}$ such that $\mathcal{D}s-\frac{i}{2}\,\mathrm{tr}(q)\,s=0$ on $M$ and $\chi s=s$ along $\Sigma$.
+By deformation invariance of the Fredholm index, the operator $\mathcal{D}-\tfrac{i}{2}\mathrm{tr}(q)$ has positive index.
+Hence, we can find a non-trivial section $s$ of $\mathcal{E}$ such that $\mathcal{D}s-\tfrac{i}{2}\,\mathrm{tr}(q)\,s=0$ on $M$ and $\chi s=s$ along $\Sigma$.
 By elliptic regularity, $s$ is smooth.
 Using Proposition 3.6, we obtain
-$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq-\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes 1)s,s\rangle$
-$\qquad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
+$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq-\tfrac{1}{2}\int_{M}\langle(\mu+i\gamma(J)\otimes\textup{id})s,s\rangle$
+$\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
 From this, the assertion follows easily. ∎
@@
   • $\varphi^{*}g_{S^{n}}\leq g_{S^{n-1}}+4g_{S^{1}}$.
-  • For each $t\in S^{1}$, the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is either Lipschitz continuous with a Lipschitz constant strictly less than $1$, or else $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n}$, $x\mapsto(x,0)$.
+  • For each $t\in S^{1}$, the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is either Lipschitz continuous with a Lipschitz constant strictly less than $1$, or else the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n}$.
 Proof.
-Fix a smooth $2$-Lipschitz function $\beta\colon[-\pi,\pi]\to[-\pi,\pi]$ such that $\beta(t)=-\pi$ for $t$ near $-\pi$, $\beta(t)=0$ for $t$ near 00, and $\beta(t)=\pi$ for $t$ near $\pi$.
-Moreover, let us fix a point $a\in S^{n-1}$.
+Let us fix a point $a\in S^{n-1}$ and a smooth function $\beta\colon[-\pi,\pi]\to[-\pi,\pi]$ with the following properties:
+  • $\beta(t)=-\pi$ for $t\in[-\pi,-\tfrac{7\pi}{8}]$.
+  • $\beta(t)=0$ for $t\in[-\tfrac{\pi}{8},\tfrac{\pi}{8}]$.
+  • $\beta(t)=\pi$ for $t\in[\tfrac{7\pi}{8},\pi]$.
+  • $0<\beta^{\prime}(t)\leq 2$ for $t\in(-\tfrac{7\pi}{8},-\tfrac{\pi}{8})\cup(\tfrac{\pi}{8},\tfrac{7\pi}{8})$.
 We consider the map
-$S^{n-1}\times[-\pi,\pi]\to S^{n},\,(x,t)\mapsto\begin{cases}(\sin\beta(t)\,a,\cos\beta(t))&\text{\rm for $t\in[-\pi,0]$},\\ (\sin\beta(t)\,x,\cos\beta(t))&\text{\rm for $t\in[0,\pi]$.}\end{cases}$
-This gives a map $\varphi\colon S^{n-1}\times S^{1}\to S^{n}$ with the required properties. ∎
+$\varphi\colon S^{n-1}\times[-\pi,\pi]\to S^{n},\,(\xi,t)\mapsto\begin{cases}(\sin\beta(t)\,a,\cos\beta(t))&\text{\rm for $t\in[-\pi,0]$},\\ (\sin\beta(t)\,\xi,\cos\beta(t))&\text{\rm for $t\in[0,\pi]$.}\end{cases}$
+It is easy to see that $\varphi$ induces a smooth map from $S^{n-1}\times S^{1}$ to $S^{n}$ with degree $\pm 1$. We next compute
+$\varphi^{*}g_{S^{n}}=\begin{cases}\beta^{\prime}(t)^{2}\,dt\otimes dt&\text{\rm for $t\in[-\pi,0]$},\\ \sin^{2}\beta(t)\,g_{S^{n-1}}+\beta^{\prime}(t)^{2}\,dt\otimes dt&\text{\rm for $t\in[0,\pi]$.}\end{cases}$
+Since $0\leq\beta^{\prime}(t)\leq 2$ for all $t\in[-\pi,\pi]$, it follows that $\varphi^{*}g_{S^{n}}\leq g_{S^{n-1}}+4\,dt\otimes dt$.
+If $t\in[-\pi,0]$, then the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is constant.
+If $t\in[0,\pi]$ and $\beta(t)\neq\tfrac{\pi}{2}$, then the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is Lipschitz continuous with Lipschitz constant $\sin\beta(t)<1$.
+If $t\in[0,\pi]$ and $\beta(t)=\tfrac{\pi}{2}$, the map $\varphi(\cdot,t)\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n},\,\xi\mapsto(\xi,0)$.
+∎
 In the following, we fix a smooth map $\varphi\colon S^{n-1}\times S^{1}\to S^{n}$ as in Lemma 3.8.
 Moreover, $r$ will denote a parameter which will be chosen sufficiently large.
-We consider the product $\tilde{M}=M\times S^{1}$, equipped with the metric $\tilde{g}=g+r^{2}\,dt^{2}$.
+We consider the product $\tilde{M}=M\times S^{1}$, equipped with the metric $\tilde{g}=g+r^{2}\,dt\otimes dt$.
 We denote by $\tilde{q}$ the pull-back $(0,2)$-tensor field of $q$ under the canonical projection $\tilde{M}\to M$.
-The boundary of $\tilde{M}$ can be identified with $\tilde{\Sigma}=\Sigma\times S^{1}$.
+We denote by $\tilde{\Sigma}=\Sigma\times S^{1}$ the boundary of $\tilde{M}$.
 We define a map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ by $\tilde{N}(x,t)=\varphi(N(x),t)$ for $x\in\Sigma$ and $t\in S^{1}$.
@@
 Lemma 3.9.
-The trace norms of $dN$ and $d\tilde{N}$ are related by
+The trace norm of $d\tilde{N}$ with respect to the metric $\tilde{g}$ satisfies
 $\|d\tilde{N}\|_{\mathrm{tr}}\leq\|dN\|_{\mathrm{tr}}+2r^{-1}.$
 Proof.
-By Lemma 3.9, we have $\varphi^{*}g_{S^{n}}\leq g_{S^{n-1}}+4g_{S^{1}}$.
-Using the definition of $\tilde{N}$, we obtain $\tilde{N}^{*}g_{S^{n}}\leq N^{*}g_{S^{n-1}}+4g_{S^{1}}$.
-Let us denote the singular values of $N\colon T_{x}\Sigma\to T_{N(x)}S^{n-1}$ by $\lambda_{1},\ldots,\lambda_{n-1}$.
+By Lemma 3.8, we have $\varphi^{*}g_{S^{n}}\leq g_{S^{n-1}}+4g_{S^{1}}$. Using the definition of $\tilde{N}$, we obtain
+$\tilde{N}^{*}g_{S^{n}}\leq N^{*}g_{S^{n-1}}+4g_{S^{1}}.$
+Let us denote the singular values of $dN_{x}\colon T_{x}\Sigma\to T_{N(x)}S^{n-1}$ by $\lambda_{1},\ldots,\lambda_{n-1}$.
 Then the eigenvalues of $N^{*}g_{S^{n-1}}$ with respect to the metric $g$ are given by $\lambda_{1}^{2},\ldots,\lambda_{n-1}^{2}$.
@@
 Note that $\tilde{\mathcal{S}}$ and $\tilde{\mathcal{E}}$ are bundles over $\tilde{M}$.
-The map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ induces a chirality operator $\tilde{\chi}$.
+Since $n$ is even, the map $\gamma_{0}\colon\mathbb{R}^{n}\to\text{\rm End}(S_{0})$ can be extended to a map $\gamma_{0}\colon\mathbb{R}^{n+1}\to\text{\rm End}(S_{0})$ satisfying the Clifford relations.
+With this understood, the map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ induces a chirality operator $\tilde{\chi}$.
 We define $\tilde{\mu}$ and $\tilde{J}$ as in (3.1) and (3.2) on $\tilde{M}$ using $\tilde{g}$ and $\tilde{q}$.
@@
 Proposition 3.10.
-Suppose that $M$ has even dimension $n\geq 2$ and that the map $N\colon\Sigma\to S^{n-1}$ has positive degree.
-For each $r>0$, we can find an element $\hat{t}\in S^{1}$ and a non-trivial section $s\in C^{\infty}(M,\mathcal{E})$ such that
+Suppose $M$ has even dimension $n\geq 2$ and the map $N\colon\Sigma\to S^{n-1}$ has positive degree. For each $r>0$, we can find a non-trivial section $s\in C^{\infty}(M,\mathcal{E})$ such that
 $\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}$
 $\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}+r^{-1}\int_{\Sigma}|s|^{2}.$
-Moreover, we have $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ along $\Sigma$.
-Here, the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$.
+[…]
 Proof.
-Since the map $N\colon\Sigma\to S^{n-1}$ has positive degree, it follows that the map $\tilde{N}\colon\tilde{\Sigma}\to S^{n}$ has positive degree as well.
-By Proposition 3.8, we can find a section $\tilde{s}\in C^{\infty}(\tilde{M},\tilde{\mathcal{E}})$ such that
-$\int_{\tilde{M}}|\tilde{s}|^{2}>0$
-and
-$\int_{\tilde{M}}|\nabla^{\tilde{\mathcal{E}},\tilde{q}}\tilde{s}|^{2}$ $\leq\tfrac{1}{2}\int_{\tilde{M}}(-\tilde{\mu}+|\tilde{J}|)|s|^{2}$
-$\quad+\tfrac{1}{2}\int_{\tilde{\Sigma}}(|\mathrm{tr}(\tilde{q})-\tilde{q}(\tilde{\nu},\tilde{\nu})|+\|d\tilde{N}\|_{\mathrm{tr}}-\tilde{H})\,|\tilde{s}|^{2}.$
-Moreover, $\tilde{s}$ satisfies $\tilde{\chi}\tilde{s}=\tilde{s}$ along the boundary $\tilde{\Sigma}$. Consequently, there exists an element $\hat{t}\in S^{1}$ such that
-$\int_{M\times\{\hat{t}\}}|\tilde{s}|^{2}>0$
-and
+[…]
+$\int_{\tilde{M}}|\nabla^{\tilde{\mathcal{E}},\tilde{q}}\tilde{s}|^{2}$ $\leq\tfrac{1}{2}\int_{\tilde{M}}(-\tilde{\mu}+|\tilde{J}|)|s|^{2}+\tfrac{1}{2}\int_{\tilde{\Sigma}}(|\mathrm{tr}(\tilde{q})-\tilde{q}(\tilde{\nu},\tilde{\nu})|+\|d\tilde{N}\|_{\mathrm{tr}}-\tilde{H})\,|\tilde{s}|^{2}.$
+Moreover, $\tilde{s}$ satisfies $\tilde{\chi}\tilde{s}=\tilde{s}$ along the boundary $\tilde{\Sigma}$.
+Consequently, there exists an element $\hat{t}\in S^{1}$ such that $\int_{M\times\{\hat{t}\}}|\tilde{s}|^{2}>0$ and
 $\int_{M\times\{\hat{t}\}}|\nabla^{\tilde{\mathcal{E}},\tilde{q}}\tilde{s}|^{2}$ $\leq\tfrac{1}{2}\int_{M\times\{\hat{t}\}}(-\tilde{\mu}+|\tilde{J}|)|s|^{2}$
-$\quad+\tfrac{1}{2}\int_{\tilde{\Sigma}}(|\mathrm{tr}(\tilde{q})-\tilde{q}(\tilde{\nu},\tilde{\nu})|+\|d\tilde{N}\|_{\mathrm{tr}}-\tilde{H})\,|\tilde{s}|^{2}.$
+$\quad+\tfrac{1}{2}\int_{\Sigma\times\{\hat{t}\}}(|\mathrm{tr}(\tilde{q})-\tilde{q}(\tilde{\nu},\tilde{\nu})|+\|d\tilde{N}\|_{\mathrm{tr}}-\tilde{H})\,|\tilde{s}|^{2}.$
 It follows from Lemma 3.9 that $\|d\tilde{N}\|_{\mathrm{tr}}\leq\|dN\|_{\mathrm{tr}}+2r^{-1}$.
-Restricting $\tilde{s}$ to $M\times\{\hat{t}\}$ yields a section $s\in C^{\infty}(M,\mathcal{E})$ such that
-$\int_{M}|s|^{2}>0$
-and
+Restricting $\tilde{s}$ to $M\times\{\hat{t}\}$ yields a section $s\in C^{\infty}(M,\mathcal{E})$ such that $\int_{M}|s|^{2}>0$ and
 $\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}$
 $\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}+r^{-1}\int_{\Sigma}|s|^{2}.$
-Moreover, since $\tilde{\chi}\tilde{s}=\tilde{s}$ along the boundary $\tilde{\Sigma}$, we conclude that $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ at each point on $\Sigma$, where the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$.
-∎
+[…]
 Corollary 3.11.
-Suppose that $M$ has even dimension $n\geq 2$ and that the map $N\colon\Sigma\to S^{n-1}$ has positive degree.
-We can find an element $\hat{t}\in S^{1}$ and a non-trivial section $s\in H^{1}(M,\mathcal{E})$ such that
-$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}$
-$\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
-Moreover, we have $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ along $\Sigma$.
-Here, the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$.
+Suppose that $M$ has even dimension $n\geq 2$ and the map $N\colon\Sigma\to S^{n-1}$ has positive degree. We can find a non-trivial section $s\in H^{1}(M,\mathcal{E})$ such that
+$\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
+[…]
 Proof.
-Consider an arbitrary sequence $r_{l}\to\infty$. For each $l$, Proposition 3.11 implies the existence of a section $s^{(l)}\in C^{\infty}(M,\mathcal{E})$ such that
-$\int_{M}|s^{(l)}|^{2}=1$
-and
+Consider an arbitrary sequence $r_{l}\to\infty$. For each $l$, Proposition 3.10 implies the existence of a section $s^{(l)}\in C^{\infty}(M,\mathcal{E})$ such that $\int_{M}|s^{(l)}|^{2}=1$ and
 $\int_{M}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s^{(l)}|^{2}$
 $\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}+r_{l}^{-1}\int_{\Sigma}|s^{(l)}|^{2}.$
-Moreover, we can find an element $\hat{t}_{l}\in S^{1}$ such that $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{t}_{l})$ along $\Sigma$, where the map $\hat{N}_{l}:\Sigma\to S^{n}$ is defined by $\hat{N}_{l}(x)=\tilde{N}(x,\hat{t}_{l})=\varphi(N(x),\hat{t}_{l})$ for all $x\in\Sigma$.
+Moreover, for each $l$, we can find an element $\hat{t}_{l}\in S^{1}$ such that $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}_{l})$ along $\Sigma$, where the map $\hat{N}_{l}\colon\Sigma\to S^{n}$ is defined by $\hat{N}_{l}(x)=\tilde{N}(x,\hat{t}_{l})=\varphi(N(x),\hat{t}_{l})$ for all $x\in\Sigma$.
 This implies
@@
 Moreover, $s_{l}|_{\Sigma}\to s|_{\Sigma}$ strongly in $L^{2}(\Sigma,\mathcal{E}|_{\Sigma})$.
-Thus, we conclude that
-$\int_{M}|s|^{2}=1$
-and
+Thus, we conclude that $\int_{M}|s|^{2}=1$ and
 $\int_{M}|\nabla^{\mathcal{E},q}s|^{2}$ $\leq\tfrac{1}{2}\int_{M}(-\mu+|J|)|s|^{2}$
 $\quad+\tfrac{1}{2}\int_{\Sigma}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}-H)\,|s|^{2}.$
-Finally, on the boundary $\Sigma$, we have $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}_{l})$ for each $l$.
-Passing to the limit as $l\to\infty$, we conclude that $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$, where the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$.
-∎
+[…]
 4. Proof of Theorem I
-Let $M$, $\Sigma$, $g$, $q$, $N$, $H$ satisfy the assumptions in Theorem I.
-In particular, we assume that $\mu\geq|J|$ at each point in $M$ and $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ at each point on the boundary $\Sigma$.
-Before we prove Theorem I, we prove some preliminary statements in Proposition 4.1.
+Throughout this section, we consider an initial data set $(M,g,q)$ together with a map $N\colon\Sigma\to S^{n-1}$ satisfying the assumptions in Theorem I.
+In particular, we assume that $\mu\geq|J|$ at each point in $M$ and $H\geq\|dN\|_{\mathrm{tr}}+|\mathrm{tr}(q)-q(\nu,\nu)|$ at each point on the boundary $\Sigma$.
 We denote the second fundamental form of $\Sigma$ in $M$ (viewed as a (1,1)-tensor) by $h$.
 For each $X\in T\Sigma$, we denote by $q(X)^{\text{\rm tan}}$ the part of $q(X)\in TM$ tangential to $\Sigma$.
+We begin by proving several preliminary statements.
 Proposition 4.1.
-The following assertions hold.
+The following statements hold.
   1. (a)
@@
   2. (b)
-    At each point $p\in\Sigma$, the eigenvalues of the second fundamental form $h$ agree with the singular values of $dN_{p}:T_{p}\Sigma\to T_{N(p)}S^{n-1}$.
+    At each point $p\in\Sigma$, the eigenvalues of the second fundamental form $h$ agree with the singular values of $dN_{p}\colon T_{p}\Sigma\to T_{N(p)}S^{n-1}$.
 In particular, the eigenvalues of $h$ are nonnegative.
   3. (c)
-    $q^{\rm tan}=0$ along $\Sigma$.
+    We have $q^{\rm tan}=0$ at each point on $\Sigma$.
 We prove Proposition 4.1 first in odd and then in even dimensions.
@@
 Suppose that $n\geq 3$ is an odd integer.
-Since by assumption $\mu\geq|J|$ on $M$ and $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ along $\Sigma$, Proposition 3.8 yields non-trivial $\nabla^{\mathcal{E},q}$-parallel section $s\in C^{\infty}(M,\mathcal{E})$ such that $\chi s=s$ holds along $\Sigma$.
+By assumption, we have $\mu\geq|J|$ on $M$ and $H\geq\|dN\|_{\mathrm{tr}}+|\mathrm{tr}(q)-q(\nu,\nu)|$ along $\Sigma$.
+Therefore, Proposition 3.7 yields non-trivial $\nabla^{\mathcal{E},q}$-parallel section $s\in C^{\infty}(M,\mathcal{E})$ such that $\chi s=s$ along $\Sigma$.
 Lemma 4.2.
@@
 Proof.
-Since $s$ is $\nabla^{\mathcal{E},q}$-parallel,
+Since $s$ is $\nabla^{\mathcal{E},q}$-parallel, we obtain
 (4.1) $\nabla_{X}^{\mathcal{E}}s=-\tfrac{i}{2}Qs(X)=-\tfrac{i}{2}\gamma(q(X))\circ s$
@@
 (4.2) $\gamma(\nu)\circ s=s\circ\gamma_{0}(N)$
-on $\Sigma$. In the next step, we differentiate (4.2) along a vector field $X\in C^{\infty}(\Sigma,T\Sigma)$. Using (4.1) we find for the left hand side
+on $\Sigma$. In the next step, we fix a tangential vector field $X$ along $\Sigma$ and differentiate (4.2) along $X$. Using (4.1) we find for the left hand side
 $\nabla_{X}^{\mathcal{E}}(\gamma(\nu)\circ s)$ $=\gamma(D_{X}\nu)\circ s+\gamma(\nu)\circ\nabla_{X}^{\mathcal{E}}s$
@@
 and for the right hand side
-$\nabla_{X}^{\mathcal{E}}(s\circ\gamma_{0}(N))$ $=(\nabla_{X}^{\mathcal{S}}s)\circ\gamma_{0}(N)+s\circ\gamma_{0}(dN(X))$
+$\nabla_{X}^{\mathcal{E}}(s\circ\gamma_{0}(N))$ $=(\nabla_{X}^{\mathcal{E}}s)\circ\gamma_{0}(N)+s\circ\gamma_{0}(dN(X))$
 $=-\tfrac{i}{2}\gamma(q(X))\circ s\circ\gamma_{0}(N)+s\circ\gamma_{0}(dN(X)).$
@@
 $=-i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\circ s.$
-For the last equality, we have used the fact that the normal component of $q(X)$ commutes with $\gamma(\nu)$ while the tangential component of $q(X)$ anticommutes with $\gamma(\nu)$. ∎
-Lemma 4.3.
-We have
-$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}=\|dN\|_{\mathrm{tr}}\,|s|^{2}$
-at each point on $\Sigma$.
-Proof.
-Let us fix a point $p\in\Sigma$. Let $e_{1},\dots,e_{n-1}$ be an orthonormal basis of $T_{p}\Sigma$. Using Lemma 4.2, we obtain
-[…]
-$=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}.$
-This implies
-$H\,|s|^{2}+i\,(\mathrm{tr}(q)-q(\nu,\nu))\,\langle\gamma(\nu)\circ s,s\rangle=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}.$
-By assumption, $H-|\mathrm{tr}(q)-q(\nu,\nu)|\geq\|dN\|_{\mathrm{tr}}$ at each point on $\Sigma$. Consequently,
-$\|dN\|_{\mathrm{tr}}\,|s|^{2}\leq H\,|s|^{2}-|\mathrm{tr}(q)-q(\nu,\nu)|\,|s|^{2}\leq-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}.$
-On the other hand, using Lemma 3.6, we obtain
-$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(dN(e_{j})),s\Big{\rangle}\leq\|dN\|_{\mathrm{tr}}\,|s|^{2}.$
-Putting these facts together, the assertion follows. ∎
+To justify the last equality, we decompose $q(X)$ into its tangential and normal components and observe that $\gamma(q(X)^{\text{\rm tan}})$ anticommutes with $\gamma(\nu)$. ∎
 Lemma 4.3.
@@
 ∎
-We define a linear subspace $L\subset S_{0}$ by
-$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ at each point on $M$}\}.$
+We define a linear subspace $L\subset S_{0}$ by $L:=\{\sigma\in S_{0}\colon\text{\rm$s\sigma=0$ at each point in $M$}\}$.
 Lemma 4.4.
@@
 Since the map $N\colon\Sigma\to S^{n-1}$ has non-zero degree, we know that the image of $N$ is all of $S^{n-1}$.
-Consequently, $L$ is invariant under $\gamma_{0}(v)$ for each $v\in\mathbb{R}^{n}$.
+Consequently, $L$ is invariant under $\gamma_{0}(\xi)$ for each $\xi\in\mathbb{R}^{n}$.
 The spinor representation of the Clifford algebra $\mathbb{C}\mathrm{l}(\mathbb{R}^{n})\to\mathrm{End}(S_{0})$ is surjective.
-Therefore, the elements $\gamma_{0}(v)$, $v\in\mathbb{R}^{n}$, generate all of $\mathrm{End}(S_{0})$.
+Therefore, the elements $\gamma_{0}(\xi)$, $\xi\in\mathbb{R}^{n}$, generate all of $\mathrm{End}(S_{0})$.
 Putting these facts together, we conclude that $L$ is invariant under $\mathrm{End}(S_{0})$.
@@
 Let $p$ be a point on the boundary $\Sigma$.
-Let $e_{1},\dots,e_{n-1}$ be an orthonormal basis of $T_{p}\Sigma$ and let $e^{0}_{1},\dots,e^{0}_{n-1}$ be an orthonormal basis of $T_{N(p)}S^{n-1}$ such that $dN(e_{j})=\lambda_{j}e^{0}_{j}$, where $\lambda_{j}\geq 0$ denote the singular values of $dN$.
+Let $e_{1},\dots,e_{n-1}$ be an orthonormal basis of $T_{p}\Sigma$ and $e^{0}_{1},\dots,e^{0}_{n-1}$ an orthonormal basis of $T_{N(p)}S^{n-1}$ such that $dN(e_{j})=\lambda_{j}e^{0}_{j}$, where $\lambda_{j}\geq 0$ denote the singular values of $dN$.
 Then $h(e_{j})=\lambda_{j}e_{j}$ and $q(e_{j})^{\text{\rm tan}}=0$ for $j=1,\ldots,n-1$.
 Proof.
-By Lemma 4.3, we know that
-$-\sum_{j=1}^{n-1}\lambda_{j}\,\langle\gamma(e_{j})\circ s\circ\gamma_{0}(e^{0}_{j}),s\rangle=\sum_{j=1}^{n-1}\lambda_{j}\,|s|^{2}.$
-This implies $-\gamma(e_{j})\circ s\circ\gamma_{0}(e^{0}_{j})=s$ whenever $\lambda_{j}>0$.
-In other words, $\gamma(e_{j})\circ s=s\circ\gamma_{0}(e^{0}_{j})$ whenever $\lambda_{j}>0$.
 Applying Lemma 4.2 with $X=e_{j}$, we obtain
+(4.3) $\gamma(h(e_{j}))\circ s+i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\circ s=\lambda_{j}\,s\circ\gamma_{0}(e^{0}_{j})$
+for $j=1,\ldots,n-1$. This implies
+$H\,$ $|s|^{2}+i\,(\text{\rm tr}(q)-q(\nu,\nu))\,\langle\gamma(\nu)\circ s,s\rangle$
+$=-\sum_{j}\langle\gamma(e_{j})\circ\gamma(h(e_{j}))\circ s,s\rangle-i\sum_{j}\,\langle\gamma(e_{j})\circ\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\circ s,s\rangle$
+$=-\sum_{j}\lambda_{j}\,\langle\gamma(e_{j})\circ s\circ\gamma_{0}(e^{0}_{j}),s\rangle.$
+In the next step, we take the absolute value on both sides. Using the triangle inequality, we obtain
+(4.4) $H\,|s|^{2}-|\text{\rm tr}(q)-q(\nu,\nu)|\,|s|^{2}\leq\|dN\|_{\mathrm{tr}}\,|s|^{2}.$
+On the other hand, our assumption implies that $\|dN\|_{\mathrm{tr}}\leq H-|\text{\rm tr}(q)-q(\nu,\nu)|$.
+Consequently, equality holds in (4.4).
+From this, we deduce that $\gamma(e_{j})\circ s\circ\gamma_{0}(e^{0}_{j})=s$ whenever $\lambda_{j}>0$.
+Substituting this into (4.3) gives
 $\gamma(h(e_{j}))\circ s+i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\circ s=\lambda_{j}\,s\circ\gamma_{0}(e^{0}_{j})=\lambda_{j}\,\gamma(e_{j})\circ s$
-for $j=1,\ldots,n-1$. By Lemma 4.6, $s$ is invertible at each point in $M$. Consequently,
+for $j=1,\ldots,n-1$. Since $s\in\text{\rm Hom}(S_{0},\mathcal{S}|_{p})$ is invertible by Lemma 4.5, it follows that
 $\gamma(h(e_{j}))+i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)=\lambda_{j}\,\gamma(e_{j})\in\text{\rm End}(\mathcal{S}|_{p})$
-for $j=1,\ldots,n-1$.
-Note that $i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\in\text{\rm End}(\mathcal{S}|_{p})$ is self-adjoint, while $\gamma(h(e_{j}))\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(e_{j})\in\text{\rm End}(\mathcal{S}|_{p})$ are skew-adjoint.
+for $j=1,\ldots,n-1$. Note that
+$i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\in\text{\rm End}(\mathcal{S}|_{p})$
+is self-adjoint, while $\gamma(h(e_{j}))\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(e_{j})\in\text{\rm End}(\mathcal{S}|_{p})$ are skew-adjoint.
 Thus, we conclude that $\gamma(h(e_{j}))=\lambda_{j}\,\gamma(e_{j})\in\text{\rm End}(\mathcal{S}|_{p})$ and $i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p})$ for $j=1,\ldots,n-1$.
@@
 4.2. Proof of Proposition 4.1 in even dimensions
-We now turn to the even-dimensional case of Proposition 4.1.
-Suppose that $n\geq 2$ is an even integer.
-Since $\mu\geq|J|$ on $M$ and $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ along $\Sigma$, Corollary 3.12 yields a non-trivial $\nabla^{\mathcal{E},q}$-parallel section $s\in H^{1}(M,\mathcal{E})$.
-Clearly, $s$ is smooth.
-Moreover, we can find an element $\hat{t}\in S^{1}$ such that $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ along $\Sigma$.
-Here, the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$.
+[…]
+[…]
+for all $x\in\Sigma$.
 Lemma 4.7.
@@
 Lemma 4.8.
-We have
-$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}=\|d\hat{N}\|_{\mathrm{tr}}\,|s|^{2}$
-at each point on $\Sigma$. Moreover, $\|d\hat{N}\|_{\mathrm{tr}}=\|dN\|_{\mathrm{tr}}$ at each point on $\Sigma$.
+We have $\|d\hat{N}\|_{\mathrm{tr}}=\|dN\|_{\mathrm{tr}}$ at each point on $\Sigma$.
 Proof.
 Let us fix a point $p\in\Sigma$. Let $e_{1},\dots,e_{n-1}$ be an orthonormal basis of $T_{p}\Sigma$. Using Lemma 4.7, we obtain
-[…]
-$=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}.$
-This implies
-$H\,|s|^{2}+i\,(\mathrm{tr}(q)-q(\nu,\nu))\,\langle\gamma(\nu)\circ s,s\rangle=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}.$
-Consequently,
-$\|dN\|_{\mathrm{tr}}\,|s|^{2}\leq H\,|s|^{2}-|\mathrm{tr}(q)-q(\nu,\nu)|\,|s|^{2}\leq-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}.$
-As in Lemma 3.6, we can bound
-$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}\leq\|d\hat{N}\|_{\mathrm{tr}}\,|s|^{2}.$
-Finally, we note that $\|d\hat{N}\|_{\mathrm{tr}}\leq\|dN\|_{\mathrm{tr}}$ by definition of $\hat{N}$. Since $|s|^{2}>0$ at each point on $M$, the assertion follows. ∎
+$H\,$ $|s|^{2}+i\,(\text{\rm tr}(q)-q(\nu,\nu))\,\langle\gamma(\nu)\circ s,s\rangle$
+$=-\sum_{j}\langle\gamma(e_{j})\circ\gamma(h(e_{j}))\circ s,s\rangle-i\sum_{j}\,\langle\gamma(e_{j})\circ\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\circ s,s\rangle$
+$=-\sum_{j}\langle\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\rangle.$
+In the next step, we take the absolute value on both sides. This gives
+$H\,|s|^{2}-|\text{\rm tr}(q)-q(\nu,\nu)|\,|s|^{2}\leq\|d\hat{N}\|_{\mathrm{tr}}\,|s|^{2}.$
+Since $|s|^{2}>0$ at each point in $M$, it follows that $H-|\text{\rm tr}(q)-q(\nu,\nu)|\leq\|d\hat{N}\|_{\mathrm{tr}}$.
+On the other hand, $\|d\hat{N}\|_{\mathrm{tr}}\leq\|dN\|_{\mathrm{tr}}$ by definition of $\hat{N}$.
+Finally, our assumption implies that
+$\|dN\|_{\mathrm{tr}}\leq H-|\text{\rm tr}(q)-q(\nu,\nu)|.$
+Putting these facts together, the assertion follows. ∎
 Lemma 4.9.
 The map $\varphi(\cdot,\hat{t})\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n}$.
-In particular, the map $\hat{N}:\Sigma\to S^{n}$ is the composition of the map $N:\Sigma\to S^{n-1}$ with the standard inclusion $S^{n-1}\hookrightarrow S^{n}$.
+In particular, the map $\hat{N}\colon\Sigma\to S^{n}$ is the composition of the map $N\colon\Sigma\to S^{n-1}$ with the standard inclusion $S^{n-1}\hookrightarrow S^{n}$.
 Proof.
-Suppose that the assertion is false.
-By Lemma 3.9, the map $\varphi(\cdot,\hat{t})$ is Lipschitz continuous with a Lipschitz constant strictly less than $1$.
-This implies $\|d\hat{N}\|_{\mathrm{tr}}<\|dN\|_{\mathrm{tr}}$ at each point on $\Sigma$ where $\|dN\|_{\mathrm{tr}}>0$.
-Using Lemma 4.9, we conclude that $\|dN\|_{\mathrm{tr}}=0$ at each point on $\Sigma$.
-This implies that $N$ is locally constant.
-This contradicts the assumption that $N$ has non-zero degree.
-∎
+[…]
 Lemma 4.10.
@@
 As above, we define a linear subspace $L\subset S_{0}$ by
-$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ at each point on $M$}\}.$
+$L:=\{\sigma\in S_{0}\colon\text{\rm$s\sigma=0$ at each point in $M$}\}.$
 Lemma 4.11.
@@
 Since the map $N\colon\Sigma\to S^{n-1}$ has non-zero degree, we know that the image of $N\colon\Sigma\to S^{n-1}$ is all of $S^{n-1}$.
-Using Lemma 4.10, we conclude that the image of the map $\hat{N}:\Sigma\to S^{n}$ is the equator in $S^{n}$.
-Thus, we conclude that $\text{\rm span}\{\hat{N}(p):p\in\Sigma\}=\mathbb{R}^{n}\subset\mathbb{R}^{n+1}$.
-By Lemma 4.12, $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n}$.
-Since the volume element acts as a scalar, we deduce that $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n+1}$.
+Using Lemma 4.9, we conclude that the image of the map $\hat{N}\colon\Sigma\to S^{n}$ is the equator in $S^{n}$.
+In particular, $\text{\rm span}\{\hat{N}(p)\colon p\in\Sigma\}=\mathbb{R}^{n}\subset\mathbb{R}^{n+1}$.
+By Lemma 4.11, $L$ is invariant under $\gamma_{0}(\xi)$ for all $\xi\in\mathbb{R}^{n}$.
+Since the volume element acts as a scalar multiple of the identity, we deduce that $L$ is invariant under $\gamma_{0}(\xi)$ for all $\xi\in\mathbb{R}^{n+1}$.
 Arguing as in the proof of Lemma 4.5, the assertion follows.
@@
 Definition 4.14.
-Let $\mathcal{T}=TM\oplus\mathbb{R}$ denote the extended tangent bundle. We define a connection $\nabla^{\mathcal{T},q}$ on $\mathcal{T}$ by
+We define a vector bundle $\mathcal{T}$ over $M$ by $\mathcal{T}=TM\oplus\mathbb{R}$. We define a bundle metric of signature $(n,1)$ on $\mathcal{T}$ by
+$\langle(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}=g(Y,Z)-\varphi\psi$
+for $(Y,\varphi),(Z,\psi)\in\mathcal{T}$. We define a connection $\nabla^{\mathcal{T},q}$ on $\mathcal{T}$ by
 $\nabla_{X}^{\mathcal{T},q}(Y,\varphi)=(D_{X}Y-\varphi\,q(X),X(\varphi)-q(X,Y))$
-for every vector field $X$ and every section $(Y,\varphi)$ of $\mathcal{T}$. We define a bundle metric of signature $(n,1)$ on $\mathcal{T}$ by
-$\langle(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}=g(Y,Z)-\varphi\psi$
-for $(Y,\varphi),(Z,\psi)\in\mathcal{T}$.
+for every vector field $X$ and every section $(Y,\varphi)$ of $\mathcal{T}$.
 Lemma 4.15.
-The connection $\nabla^{\mathcal{T},q}$ is compatible with $\langle\cdot,\cdot\rangle_{\mathcal{T}}$.
+The connection $\nabla^{\mathcal{T},q}$ is compatible with the bundle metric $\langle\cdot,\cdot\rangle_{\mathcal{T}}$.
 Proof.
-Let $X$ be a vector field, and let $(Y,\varphi)$, $(Z,\psi)$ be two sections of $\mathcal{T}$. Then
+Let $X$ be a vector field on $M$, and let $(Y,\varphi)$ and $(Z,\psi)$ be two sections of $\mathcal{T}$. Then
+$\langle\nabla_{X}^{\mathcal{T},q}(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}=g(D_{X}Y,Z)-\varphi\,q(X,Z)-X(\varphi)\,\psi+q(X,Y)\,\psi$
+and
+$\langle(Y,\varphi),\nabla_{X}^{\mathcal{T},q}(Z,\psi)\rangle_{\mathcal{T}}=g(Y,D_{X}Z)-q(X,Y)\,\psi-\varphi\,X(\psi)+\varphi\,q(X,Z).$
+This gives
 $\langle\nabla_{X}^{\mathcal{T},q}(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}$ $+\langle(Y,\varphi),\nabla_{X}^{\mathcal{T},q}(Z,\psi)\rangle_{\mathcal{T}}$
-$=g(D_{X}Y,Z)+g(Y,D_{X}Z)-X(\varphi)\,\psi-\varphi\,X(\psi)$
-$=X(g(Y,Z)-\varphi\psi)$
-$=X(\langle(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}).$
+$=X(g(Y,Z)-\varphi\psi)=X(\langle(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}).$
 This completes the proof of Lemma 4.15. ∎
@@
 Suppose that $u,v$ are $\nabla^{\mathcal{S},q}$-parallel sections of $\mathcal{S}$.
-We define a vector field $Y$ on $M$ by the requirement $g(Y,Z)=\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)$ for all $Z\in TM$.
+We define a vector field $Y$ on $M$ so that $g(Y,Z)=\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)$ for all $Z\in TM$.
 Moreover, we define a scalar function $\varphi$ on $M$ by $\varphi=-\text{\rm Re}(\langle u,v\rangle)$.
@@
 Proof.
-By assumption,
+Let $X$ be a vector field on $M$. By assumption,
 $\nabla_{X}^{\mathcal{S}}u+\tfrac{i}{2}\,\gamma(q(X))\,u=\nabla_{X}^{\mathcal{S}}v+\tfrac{i}{2}\,\gamma(q(X))\,v=0.$
@@
 $=\tfrac{1}{2}\,\text{\rm Re}(\langle\gamma(Z)\,\gamma(q(X))\,u,v\rangle)+\tfrac{1}{2}\,\text{\rm Re}(\langle\gamma(q(X))\,\gamma(Z)\,u,v\rangle)$
-$=-\text{\rm Re}(\langle u,v\rangle)\,q(X,Z)$
-$=\varphi\,q(X,Z)$
+$=-q(X,Z)\,\text{\rm Re}(\langle u,v\rangle)$
+$=q(X,Z)\,\varphi$
 and
@@
 $=\text{\rm Re}(i\,\langle\gamma(q(X))\,u,v\rangle)$
-$=g(Y,q(X))$
 $=q(X,Y).$
@@
 Lemma 4.17.
-Let $p$ be an arbitrary point in $M$.
-Consider the set of all elements $(Y,\varphi)\in\mathcal{T}|_{p}$ which can be written in the form $g(Y,Z)=\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)$ and $\varphi=-\text{\rm Re}(\langle u,v\rangle)$, where $u,v\in\mathcal{S}|_{p}$.
+Suppose that $p$ is an arbitrary point in $M$.
+Consider the set of all pairs $(Y,\varphi)\in\mathcal{T}|_{p}$ with the property that there exist $u,v\in\mathcal{S}|_{p}$ such that $g(Y,Z)=\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)$ for all $Z\in TM$ and $\varphi=-\text{\rm Re}(\langle u,v\rangle)$.
 This set spans $\mathcal{T}|_{p}$.
 Proof.
-Suppose that the assertion is false. Then there exists a non-zero element $(Z,\psi)\in\mathcal{T}|_{p}$ such that
+Suppose that the assertion is false. Then there exists a non-zero pair $(Z,\psi)\in\mathcal{T}|_{p}$ such that
 $\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)+\text{\rm Re}(\psi\,\langle u,v\rangle)=0$
-for all $u,v\in\mathcal{S}|_{p}$. From this, we deduce that $i\,\gamma(Z)+\psi\,\text{\rm id}=0\in\text{\rm End}(\mathcal{S}|_{p})$. This implies $Z=0$ and $\psi=0$. This is a contradiction. ∎
+for all $u,v\in\mathcal{S}|_{p}$. This implies that $i\,\gamma(Z)+\psi\,\textup{id}=0\in\text{\rm End}(\mathcal{S}|_{p})$. From this, we deduce that $Z=0$ and $\psi=0$. This is a contradiction. ∎
 Proposition 4.18.
@@
 Proof.
-By Proposition 4.1, there exists a $\nabla^{\mathcal{E},q}$-parallel section $s$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $s$ is invertible at each point in $M$.
+According to Proposition 4.1, there exists a $\nabla^{\mathcal{E},q}$-parallel section $s$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $s$ is invertible at each point in $M$.
 Consequently, the spinor bundle $\mathcal{S}$ admits a trivialization by $\nabla^{\mathcal{S},q}$-parallel sections.
@@
 $\langle(E^{(\alpha)},\zeta^{(\alpha)}),(E^{(\beta)},\zeta^{(\beta)})\rangle_{\mathcal{T}}=m_{\alpha\beta},$
-where $m_{\alpha\beta}$ is a constant $(n+1)\times(n+1)$-matrix. The matrix $m_{\alpha\beta}$ is invertible with signature $(n,1)$. By taking suitable linear combinations, we may arrange that
+where $m_{\alpha\beta}$ is a constant $(n+1)\times(n+1)$-matrix. The matrix $m_{\alpha\beta}$ is invertible and has signature $(n,1)$. By taking suitable linear combinations, we may arrange that
 $m_{\alpha\beta}=\begin{cases}-1&\text{\rm for $\alpha=\beta=0$}\\ 1&\text{\rm for $\alpha=\beta>0$}\\ 0&\text{\rm for $\alpha\neq\beta$}\end{cases}$
@@
 for all $Y,Z\in TM$. Moreover, choosing $\varphi=0$ and $Z=0$, we find
-(4.4) $00=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,g(E^{(\alpha)},Y)\,\zeta^{(\beta)}=$ 0
+(4.6) $0=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,g(E^{(\alpha)},Y)\,\zeta^{(\beta)}=0$
 for all $Y\in TM$. Finally, choosing $Y=Z=0$ and $\varphi=\psi=1$ yields
-(4.5) $-1=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,\zeta^{(\alpha)}\,\zeta^{(\beta)}.$
-Since $(E^{(\alpha)},\zeta^{(\alpha)})$ is $\nabla^{\mathcal{T},q}$-parallel, it follows that the covariant derivative satisfies $DE^{(\alpha)}=\zeta^{(\alpha)}q$ and is hence symmetric.
-In particular, the one-form $g(E^{(\alpha)},\cdot)$ is closed.
+(4.7) $1=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,\zeta^{(\alpha)}\,\zeta^{(\beta)}.$
+Since $(E^{(\alpha)},\zeta^{(\alpha)})$ is $\nabla^{\mathcal{T},q}$-parallel, it follows that the covariant derivative of $E^{(\alpha)}$ is given by $\zeta^{(\alpha)}q$ and is hence symmetric.
+Thus, the one-form $g(E^{(\alpha)},\cdot)$ is closed.
 We may locally write $g(E^{(\alpha)},\cdot)=df^{(\alpha)}$ for some scalar function $f^{(\alpha)}$.
@@
 $=(g(E^{(0)},X),g(E^{(1)},X),\ldots,g(E^{(n)},X))\in\mathbb{R}^{n,1}.$
-It follows from (4.4) and (4.5) that the unit normal vector to the immersion $F$ is given by $(\zeta^{(0)},\zeta^{(1)},\ldots,\zeta^{(n)})\in\mathbb{R}^{n,1}$.
-Using the fact that $(E^{(\alpha)},\zeta^{(\alpha)})$ is $\nabla^{\mathcal{T},q}$-parallel, it is easy to see that the second fundamental form associated with $F$ is equal to $q$.
-Note that the isometric immersion $F=(f^{(0)},f^{(1)},\ldots,f^{(n)})$ is indeed globally well-defined if $M$ is simply connected.
+It follows from (4.6) and (4.7) that the unit normal vector to the immersion $F$ in $\mathbb{R}^{n,1}$ is given by $(\zeta^{(0)},\zeta^{(1)},\ldots,\zeta^{(n)})$.
+Since $(E^{(\alpha)},\zeta^{(\alpha)})$ is $\nabla^{\mathcal{T},q}$-parallel, it is easy to see that the second fundamental form associated with the immersion $F$ is equal to $q$.
+If $M$ is simply connected, the isometric immersion $F=(f^{(0)},f^{(1)},\ldots,f^{(n)})$ is globally well-defined.
 If $M$ is not simply connected, then $F$ may not be globally well-defined, but its differential is globally well-defined.
@@
 Definition 4.19.
-We define $\mathcal{W}=\mathcal{S}\oplus\mathcal{S}$. We define a connection $\nabla^{\mathcal{W}}$ on $\mathcal{W}$ by
+We define a vector bundle $\mathcal{W}$ over $M$ by $\mathcal{W}=\mathcal{S}\oplus\mathcal{S}$. We define a connection $\nabla^{\mathcal{W}}$ on $\mathcal{W}$ by
 $\nabla_{X}^{\mathcal{W}}(u,v)=(\nabla_{X}^{\mathcal{S},q}u,\nabla_{X}^{\mathcal{S},-q}v),$
@@
 Definition 4.20.
-For each point $p\in M$, we define a map $\gamma^{\mathcal{W}}:\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ by
+For each point $p\in M$, we define a map $\gamma^{\mathcal{W}}\colon\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ by
 $\gamma^{\mathcal{W}}(T)\,(u,v)=(-iv,iu)$
@@
 $\gamma^{\mathcal{W}}(X)\,(u,v)=(\gamma(X)\,v,\gamma(X)\,u),$
-where $X\in TM$ and $u,v\in\mathcal{S}$.
+where $X$ denotes a vector field on $M$ and $u,v\in\mathcal{S}|_{p}$.
 Lemma 4.21.
-The map $\gamma^{\mathcal{W}}:\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ satisfies the Clifford relations with respect to the Lorentzian inner product on $\mathbb{R}^{n,1}$. In other words,
-$\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)+\gamma^{\mathcal{W}}(b)\,\gamma^{\mathcal{W}}(a)=-2\,\langle a,b\rangle_{\mathbb{R}^{n,1}}\,\text{\rm id}$
+The map $\gamma^{\mathcal{W}}\colon\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ satisfies the Clifford relations with respect to the Lorentzian inner product on $\mathbb{R}^{n,1}$.
+In other words,
+$\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)+\gamma^{\mathcal{W}}(b)\,\gamma^{\mathcal{W}}(a)=-2\,\langle a,b\rangle_{\mathbb{R}^{n,1}}\,\textup{id}$
 for all $a,b\in\mathbb{R}^{n,1}$.
 Proof.
-Note that
-$\gamma^{\mathcal{W}}(X)\,\gamma^{\mathcal{W}}(Y)+\gamma^{\mathcal{W}}(Y)\,\gamma^{\mathcal{W}}(X)=-2\,g(X,Y)\,\text{\rm id}$
+We compute
+$\gamma^{\mathcal{W}}(X)\,\gamma^{\mathcal{W}}(Y)+\gamma^{\mathcal{W}}(Y)\,\gamma^{\mathcal{W}}(X)=-2\,g(X,Y)\,\textup{id}$
 for all $X,Y\in TM$. Moreover,
 $\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(X)+\gamma^{\mathcal{W}}(X)\,\gamma^{\mathcal{W}}(T)=0$
-for all $X\in TM$ and
-$(\gamma^{\mathcal{W}}(T))^{2}=\text{\rm id}.$
-This completes the proof of Lemma 4.22. ∎
+for all $X\in TM$. Finally, $(\gamma^{\mathcal{W}}(T))^{2}=\textup{id}$. This completes the proof of Lemma 4.21. ∎
 Lemma 4.22.
@@
 Proof.
+Let $X$ be a section of $TM$.
 We decompose $a=Y+fT$, where $Y$ is a section of $TM$ and $f$ is a scalar function on $M$.
 Then $\bar{D}_{X}Y+\bar{D}_{X}(fT)=0$, where $\bar{D}$ denotes the standard flat connection on $\mathbb{R}^{n,1}$.
-Using the formula $\bar{D}_{X}Y=D_{X}Y+q(X,Y)\,T$, we obtain
+Using the identity $\bar{D}_{X}Y=D_{X}Y+q(X,Y)\,T$, we obtain
 $\nabla_{X}^{\mathcal{W}}$ $(\gamma^{\mathcal{W}}(Y)\,(u,v))-\gamma^{\mathcal{W}}(Y)\,\nabla_{X}^{\mathcal{W}}(u,v)-\gamma^{\mathcal{W}}(\bar{D}_{X}Y)\,(u,v)$
 $=\nabla_{X}^{\mathcal{W}}(\gamma(Y)\,v,\gamma(Y)\,u)-\gamma^{\mathcal{W}}(Y)\,(\nabla_{X}^{\mathcal{S},q}u,\nabla_{X}^{\mathcal{S},-q}v)$
-$-\gamma^{\mathcal{W}}(D_{X}Y)\,(u,v)-q(X,Y)\,\gamma^{\mathcal{W}}(T)\,(u,v)$
+$\quad-\gamma^{\mathcal{W}}(D_{X}Y)\,(u,v)-q(X,Y)\,\gamma^{\mathcal{W}}(T)\,(u,v)$
 (4.8) $=(\nabla_{X}^{\mathcal{S},q}(\gamma(Y)\,v),\nabla_{X}^{\mathcal{S},-q}(\gamma(Y)\,u))-(\gamma(Y)\,\nabla_{X}^{\mathcal{S},-q}v,\gamma(Y)\,\nabla_{X}^{\mathcal{S},q}u)$
-$-(\gamma(D_{X}Y)\,v,\gamma(D_{X}Y)\,u)-q(X,Y)\,(-iv,iu)$
-$=0$
-for all $X,Y\in TM$. Using the formula $\bar{D}_{X}T=q(X)$, we obtain
+$\quad-(\gamma(D_{X}Y)\,v,\gamma(D_{X}Y)\,u)-q(X,Y)\,(-iv,iu)$
+$=0.$
+Using the identity $\bar{D}_{X}T=q(X)$, we obtain
 $\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(T)\,(u,v))-\gamma^{\mathcal{W}}(T)\,\nabla_{X}^{\mathcal{W}}(u,v)-\gamma^{\mathcal{W}}(\bar{D}_{X}T)\,(u,v)$
@@
 $=(-i\,\nabla_{X}^{\mathcal{S},q}v,i\,\nabla_{X}^{\mathcal{S},-q}u)-(-i\,\nabla_{X}^{\mathcal{S},-q}v,i\,\nabla_{X}^{\mathcal{S},q}u)-(\gamma(q(X))\,v,\gamma(q(X))\,u)$
-$=0$
-for all $X\in TM$. This implies
-(4.7) $\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(fT)\,(u,v))-\gamma^{\mathcal{W}}(fT)\,\nabla_{X}^{\mathcal{W}}(u,v)-\gamma^{\mathcal{W}}(\bar{D}_{X}(fT))\,(u,v)=0$
-for all $X\in TM$. In the next step, we add (4.4) and (4.7). Using the identity $\bar{D}_{X}Y+\bar{D}_{X}(fT)=0$, we conclude that
-$\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(a)\,(u,v))-\gamma^{\mathcal{W}}(a)\,\nabla_{X}^{\mathcal{W}}(u,v)=0$
-for all $X\in TM$. This completes the proof of Lemma 4.23. ∎
+$=0.$
+This implies
+(4.9) $\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(fT)\,(u,v))-\gamma^{\mathcal{W}}(fT)\,\nabla_{X}^{\mathcal{W}}(u,v)-\gamma^{\mathcal{W}}(\bar{D}_{X}(fT))\,(u,v)=0.$
+In the next step, we add (4.4) and (4.9). Using the identity $\bar{D}_{X}Y+\bar{D}_{X}(fT)=0$, we conclude that
+$\nabla_{X}^{\mathcal{W}}(\gamma^{\mathcal{W}}(a)\,(u,v))-\gamma^{\mathcal{W}}(a)\,\nabla_{X}^{\mathcal{W}}(u,v)=0.$
+This completes the proof of Lemma 4.22. ∎
 Definition 4.23.
-For each point $p\in M$, we define a map $\gamma^{\mathcal{W}}:\Lambda^{2}\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ by
-$\gamma^{\mathcal{W}}(a\wedge b)=\frac{1}{2}\,(\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)-\gamma^{\mathcal{W}}(b)\,\gamma^{\mathcal{W}}(a)).$
+For each point $p\in M$, we define a map $\gamma^{\mathcal{W}}\colon\Lambda^{2}\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$ by
+$\gamma^{\mathcal{W}}(a\wedge b)=\tfrac{1}{2}\,(\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)-\gamma^{\mathcal{W}}(b)\,\gamma^{\mathcal{W}}(a)).$
 Note that $\gamma^{\mathcal{W}}(a\wedge b)=\gamma^{\mathcal{W}}(a)\,\gamma^{\mathcal{W}}(b)$ whenever $\langle a,b\rangle_{\mathbb{R}^{n,1}}=0$.
+By Proposition 4.1, there exists a $\nabla^{\mathcal{E},q}$-parallel section $s$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $s$ is invertible at each point in $M$.
+Moreover, $s$ satisfies the boundary condition $\gamma(\nu)\circ s=s\circ\gamma_{0}(N)$ at each point on $\Sigma$.
+Replacing $q$ by q-q, we can find a $\nabla^{\mathcal{E},-q}$-parallel section $t$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $t$ is invertible at each point in $M$.
+Moreover, $t$ satisfies the boundary condition $\gamma(\nu)\circ t=t\circ\gamma_{0}(N)$ at each point on $\Sigma$.
+In the following, we fix sections $s$ and $t$ of $\text{\rm Hom}(S_{0},\mathcal{S})$ satisfying the properties above.
+For each $\sigma\in S_{0}$ and each $\tau\in S_{0}$, the pair $(s\sigma,t\tau)$ defines a $\nabla^{\mathcal{W}}$-parallel section of the vector bundle $\mathcal{W}$.
+Definition 4.24.
+We denote by $\mathcal{Y}$ the set of all pairs $(\xi,\omega)\in\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}$ with the property that, for all $\sigma,\tau\in S_{0}$,
+$(s\gamma_{0}(\xi)\sigma,-t\gamma_{0}(\xi)\tau)=i\,\gamma^{\mathcal{W}}(\omega)\,(s\sigma,t\tau)$
+at each point on $M$.
+Let $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ denote the image of $\mathcal{Y}$ under the projection $\mathcal{Y}\to\Lambda^{2}\mathbb{R}^{n,1},\,(\xi,\omega)\mapsto\omega$.
+Clearly, $\mathcal{Y}$ is a linear subspace of $\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}$, and $\mathcal{Z}$ is a linear subspace of $\Lambda^{2}\mathbb{R}^{n,1}$.
 Lemma 4.25.
-Suppose that $p$ is a point in $M$ and $\omega\in\Lambda^{2}\mathbb{R}^{n,1}$ is a two-form with the property that $\gamma^{\mathcal{W}}(\omega)=0\in\text{\rm End}(\mathcal{W}|_{p})$. Then $\omega=0$.
+Suppose that $\xi\in\mathbb{R}^{n}$ satisfies $(\xi,0)\in\mathcal{Y}$. Then $\xi=0$.
 Proof.
-We decompose $\omega=\mu+T\wedge\eta$, where $\mu\in\Lambda^{2}T_{p}M$ and $\eta\in T_{p}M$. By assumption,
-00 0 $=\gamma^{\mathcal{W}}(\omega)\,(u,v)$
-$=\gamma^{\mathcal{W}}(\mu)\,(u,v)+\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\eta)\,(u,v)$
-$=(\gamma(\mu)\,u-i\,\gamma(\eta)\,u,\gamma(\mu)\,v+i\,\gamma(\eta)\,v)$
-for all $u,v\in\mathcal{S}|_{p}$.
-Therefore, $\gamma(\mu)-i\,\gamma(\eta)=0\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(\mu)+i\,\gamma(\eta)=0\in\text{\rm End}(\mathcal{S}|_{p})$.
-This implies $\gamma(\mu)=0\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(\eta)=0\in\text{\rm End}(\mathcal{S}|_{p})$.
-Thus, $\mu=0$ and $\eta=0$.
-This completes the proof of Lemma 4.25.
+Let us fix an arbitrary point $p\in M$. Since $(\xi,0)\in\mathcal{Y}$, we obtain
+$(s\gamma_{0}(\xi)\sigma,-t\gamma_{0}(\xi)\tau)=(0,0)$
+for all $\sigma,\tau\in S_{0}$. Since $s,t\in\text{\rm Hom}(S_{0},\mathcal{S}|_{p})$ are invertible, it follows that $\gamma_{0}(\xi)=0\in\text{\rm End}(S_{0})$. Thus, $\xi=0$. ∎
+Lemma 4.26.
+Suppose that $\omega\in\Lambda^{2}\mathbb{R}^{n,1}$ satisfies $(0,\omega)\in\mathcal{Y}$. Then $\omega=0$.
+Proof.
+Fix an arbitrary point $p\in M$.
+We may write $\omega=\rho+T\wedge\eta$, where $\rho\in\Lambda^{2}T_{p}M$ and $\eta\in T_{p}M$.
+Since $\langle T,\eta\rangle_{\mathbb{R}^{n,1}}=0$, we obtain $\gamma^{\mathcal{W}}(T\wedge\eta)=\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\eta)$.
+Since $(0,\omega)\in\mathcal{Y}$, we compute
+$(0,0)$ $=\gamma^{\mathcal{W}}(\omega)\,(s\sigma,t\tau)$
+$=\gamma^{\mathcal{W}}(\rho)\,(s\sigma,t\tau)+\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\eta)\,(s\sigma,t\tau)$
+$=(\gamma(\rho)s\sigma-i\,\gamma(\eta)s\sigma,\gamma(\rho)t\tau+i\,\gamma(\eta)t\tau)$
+for all $\sigma,\tau\in S_{0}$. Since $s,t\in\text{\rm Hom}(S_{0},\mathcal{S}|_{p})$ are invertible, it follows that
+[…]
+This gives $\gamma(\rho)=0\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(\eta)=0\in\text{\rm End}(\mathcal{S}|_{p})$. Thus, $\rho=0$ and $\eta=0$. This finally implies $\omega=0$. ∎
+Lemma 4.27.
+Let $p\in\Sigma$ be a boundary point. Then $(N(p),T(p)\wedge\nu(p))\in\mathcal{Y}$.
+Proof.
+Let us fix an arbitrary point $p\in\Sigma$, and let us fix elements $\sigma,\tau\in S_{0}$.
+The boundary conditions for $s$ and $t$ imply that $\gamma(\nu)\circ s=s\circ\gamma_{0}(N)$ and $\gamma(\nu)\circ t=t\circ\gamma_{0}(N)$ at the point $p$.
+Since $\langle T,\nu\rangle_{\mathbb{R}^{n,1}}=0$, we obtain $\gamma^{\mathcal{W}}(T\wedge\nu)=\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\nu)$.
+This gives
+$(s\gamma_{0}(N)\sigma,-t\gamma_{0}(N)\tau)$ $=(\gamma(\nu)s\sigma,-\gamma(\nu)t\tau)$
+$=i\,\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\nu)\,(s\sigma,t\tau)$
+$=i\,\gamma^{\mathcal{W}}(T\wedge\nu)\,(s\sigma,t\tau)$
+at the point $p$. For abbreviation, let $\xi=N(p)\in\mathbb{R}^{n}$, $a=T(p)\in\mathbb{R}^{n,1}$, and $b=\nu(p)\in\mathbb{R}^{n,1}$. Then
+$(s\gamma_{0}(\xi)\sigma,-t\gamma_{0}(\xi)\tau)=i\,\gamma^{\mathcal{W}}(a\wedge b)\,(s\sigma,t\tau)$
+at the point $p$.
+Note that $(s\gamma_{0}(\xi)\sigma,-t\gamma_{0}(\xi)\tau)$ is a $\nabla^{\mathcal{W}}$-parallel section of the vector bundle $\mathcal{W}$.
+Moreover, it follows from Lemma 4.22 that $\gamma^{\mathcal{W}}(a\wedge b)\,(s\sigma,t\tau)$ is a $\nabla^{\mathcal{W}}$-parallel section of the vector bundle $\mathcal{W}$.
+Putting these facts together, we conclude that
+$(s\gamma_{0}(\xi)\sigma,-t\gamma_{0}(\xi)\tau)=i\,\gamma^{\mathcal{W}}(a\wedge b)\,(s\sigma,t\tau)$
+at each point in $M$. Since $\sigma,\tau\in S_{0}$ are arbitrary, it follows that $(\xi,a\wedge b)\in\mathcal{Y}$. ∎
+Proposition 4.28.
+We have
+$\mathcal{Y}=\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p))\colon p\in\Sigma\}.$
+The projection $\mathcal{Y}\to\mathbb{R}^{n},\,(\xi,\omega)\mapsto\xi$ is bijective. In particular, $\mathcal{Y}$ has dimension $n$.
+Proof.
+We consider the set
+$\mathcal{Y}_{0}=\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p))\colon p\in\Sigma\}\subset\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}.$
+It follows from Lemma 4.27 that $\mathcal{Y}_{0}\subset\mathcal{Y}$.
+By Lemma 4.26, the projection $\mathcal{Y}\to\mathbb{R}^{n},\,(\xi,\omega)\mapsto\xi$ is injective.
+On the other hand, since the map $N\colon\Sigma\to S^{n-1}$ is surjective, it follows that the projection $\mathcal{Y}_{0}\to\mathbb{R}^{n},\,(\xi,\omega)\mapsto\xi$ is surjective.
+Putting these facts together, we conclude that $\mathcal{Y}_{0}=\mathcal{Y}$ and the projection $\mathcal{Y}\to\mathbb{R}^{n},\,(\xi,\omega)\mapsto\xi$ is bijective.
 ∎
-By Proposition 4.1, there exists a $\nabla^{\mathcal{E},q}$-parallel section $s$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $s$ is invertible at each point in $M$.
-Moreover, $s$ satisfies the boundary condition $\gamma(\nu)\circ s=s\circ\gamma_{0}(N)$ at each point on the boundary $\Sigma$.
-Replacing $q$ by $-q$, we can find a $\nabla^{\mathcal{E},-q}$-parallel section $t$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $t$ is invertible at each point in $M$.
-Moreover, $t$ satisfies the boundary condition $\gamma(\nu)\circ t=t\circ\gamma_{0}(N)$ at each point on the boundary $\Sigma$.
-We denote by $\mathcal{Y}$ the set of all pairs $(v,\omega)\in\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}$ with the property that
-$i\,\gamma^{\mathcal{W}}(\omega)\circ(s,t)=(s\circ\gamma_{0}(v),-t\circ\gamma_{0}(v))$
-at each point on $M$.
-Clearly, $\mathcal{Y}$ is a linear subspace of $\mathbb{R}^{n}\times\Lambda^{2}\mathbb{R}^{n,1}$.
-Let $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ denote the image of $\mathcal{Y}$ under the projection $\mathcal{Y}\to\Lambda^{2}\mathbb{R}^{n,1},\,(v,\omega)\mapsto\omega$.
-Proposition 4.26.
-We have $\mathcal{Y}=\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p)):p\in\Sigma\}$.
-The projection $\mathcal{Y}\to\mathbb{R}^{n},\,(v,\omega)\mapsto v$ is bijective.
-In particular, $\mathcal{Y}$ has dimension $n$.
+Corollary 4.29.
+We have $\mathcal{Z}=\{\mathbb{R}\cdot(T(p)\wedge\nu(p))\colon p\in\Sigma\}$. Moreover, $\mathcal{Z}$ has dimension $n$.
 Proof.
-Using the boundary conditions satisfied by $s$ and $t$, we compute
-$i\,\gamma^{\mathcal{W}}(T\wedge\nu)\circ(s,t)$ $=i\,\gamma^{\mathcal{W}}(T)\,\gamma^{\mathcal{W}}(\nu)\circ(s,t)$
-$=(\gamma(\nu)\circ s,-\gamma(\nu)\circ t)$
-$=(s\circ\gamma_{0}(N),-t\circ\gamma_{0}(N))$
-at each point on $\Sigma$. Let us fix a boundary point $p\in\Sigma$, and let $v=N(p)\in\mathbb{R}^{n}$, $a=T(p)\in\mathbb{R}^{n,1}$, and $b=\nu(p)\in\mathbb{R}^{n,1}$. Then
-$i\,\gamma^{\mathcal{W}}(a\wedge b)\circ(s,t)=(s\circ\gamma_{0}(v),-t\circ\gamma_{0}(v))$
-at the point $p$. Since the left hand side and the right hand side are parallel sections with respect to the same connection, we conclude that
-$i\,\gamma^{\mathcal{W}}(a\wedge b)\circ(s,t)=(s\circ\gamma_{0}(v),-t\circ\gamma_{0}(v))$
-at each point in $M$. Thus, $(v,a\wedge b)\in\mathcal{Y}$. To summarize, we have shown that
-$\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p)):p\in\Sigma\}\subset\mathcal{Y}.$
-Since the map $N:\Sigma\to S^{n-1}$ is surjective, it follows that the projection $\mathcal{Y}\to\mathbb{R}^{n},\,(v,\omega)\mapsto v$ is surjective.
-On the other hand, it follows from Lemma 4.25 that the projection $\mathcal{Y}\to\mathbb{R}^{n},\,(v,\omega)\mapsto v$ is injective.
-Putting these facts together, we conclude that
-$\{\mathbb{R}\cdot(N(p),T(p)\wedge\nu(p)):p\in\Sigma\}=\mathcal{Y}.$
-This completes the proof of Proposition 4.26. ∎
-Corollary 4.27.
-We have $\mathcal{Z}=\{\mathbb{R}\cdot(T(p)\wedge\nu(p)):p\in\Sigma\}$. Moreover, $\mathcal{Z}$ has dimension $n$.
-Proof.
-The first statement follows directly from Proposition 4.26.
-The second statement follows from the fact that $\mathcal{Y}$ has dimension $n$ and the projection $\mathcal{Y}\to\Lambda^{2}\mathbb{R}^{n,1},\,(v,\omega)\mapsto\omega$ is injective.
+The first statement follows immediately from Proposition 4.28.
+The second statement follows from the fact that $\mathcal{Y}$ has dimension $n$ and the projection $\mathcal{Y}\to\Lambda^{2}\mathbb{R}^{n,1},\,(\xi,\omega)\mapsto\omega$ is injective (see Lemma 4.25).
 ∎
@@
 Proof.
-By Theorem 6 in [LA], there are two possibilities:
-Case 1: We have $n=3$ and we can find three linearly independent vectors $b_{0},b_{1},b_{2}\in\mathbb{R}^{3,1}$ with the property that $\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{1}\wedge b_{2},b_{2}\wedge b_{0}\}$.
-We can find two linearly independent vectors $\tilde{b}_{1},\tilde{b}_{2}\in\text{\rm span}\{b_{0},b_{1},b_{2}\}$ such that $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}$ is spacelike.
-Then $\tilde{b}_{1}\wedge\tilde{b}_{2}\in\mathcal{Z}$.
-By Corollary 4.27, we can find a point $p\in\Sigma$ with the property that $\tilde{b}_{1}\wedge\tilde{b}_{2}$ is a non-zero scalar multiple of $T(p)\wedge\nu(p)$.
-In particular, $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}=\text{\rm span}\{T(p),\nu(p)\}$.
-This contradicts the fact that $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}$ does not contain a timelike vector.
-Case 2: There exists a basis $\{b_{0},b_{1},\ldots,b_{n}\}$ of $\mathbb{R}^{n,1}$ with the property that $\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{0}\wedge b_{2},\ldots,b_{0}\wedge b_{n}\}$.
+It follows from Corollary 4.29 that the subspace $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ is decomposable in the sense of [Ghorpade-Patil-Pillai].
+(See [Ghorpade-Patil-Pillai]*p.
+55 for the definition.) Applying [Ghorpade-Patil-Pillai]*Theorem 6 with $l=2$ and $r=n$, we conclude that the subspace $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ is close in the sense of [Ghorpade-Patil-Pillai].
+(See [Ghorpade-Patil-Pillai]*p.
+57 for the definition.) Thus, there are two possibilities.
+Case 1: The subspace $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ is close of Type I (cf.
+[Ghorpade-Patil-Pillai], p.
+57).
+In this case, there exists a basis $\{b_{0},b_{1},\ldots,b_{n}\}$ of $\mathbb{R}^{n,1}$ with the property that
+$\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{0}\wedge b_{2},\ldots,b_{0}\wedge b_{n}\}.$
 We claim that $b_{0}$ is timelike.
 Suppose that $b_{0}$ is spacelike or null.
-We can find a vector $\tilde{b}\in\mathbb{R}^{n,1}$ such that $\tilde{b}$ is orthogonal to $b_{0}$ and $\tilde{b}$ is spacelike.
+We can find a spacelike vector $\tilde{b}\in\mathbb{R}^{n,1}$ such that $\langle b_{0},\tilde{b}\rangle_{\mathbb{R}^{n,1}}=0$.
 This implies that $b_{0}$ and $\tilde{b}$ are linearly independent.
@@
 This contradicts the fact that $\text{\rm span}\{b_{0},\tilde{b}\}$ does not contain a timelike vector.
-∎
+Case 2: The subspace $\mathcal{Z}\subset\Lambda^{2}\mathbb{R}^{n,1}$ is close of Type II, but not close of Type I (cf.
+[Ghorpade-Patil-Pillai]*p.
+57).
+In this case, $n=3$.
+Moreover, there exists a collection of three linearly independent vectors $b_{0},b_{1},b_{2}\in\mathbb{R}^{3,1}$ with the property that $\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{1}\wedge b_{2},b_{2}\wedge b_{0}\}$.
+We can find two linearly independent vectors $\tilde{b}_{1},\tilde{b}_{2}\in\text{\rm span}\{b_{0},b_{1},b_{2}\}$ such that $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}$ is spacelike.
+Then $\tilde{b}_{1}\wedge\tilde{b}_{2}\in\mathcal{Z}$.
+By Corollary 4.29, we can find a point $p\in\Sigma$ with the property that $\tilde{b}_{1}\wedge\tilde{b}_{2}$ is a non-zero scalar multiple of $T(p)\wedge\nu(p)$.
+In particular,
+$\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}=\text{\rm span}\{T(p),\nu(p)\}.$
+This contradicts the fact that $\text{\rm span}\{\tilde{b}_{1},\tilde{b}_{2}\}$ does not contain a timelike vector. This case cannot occur. ∎
 By scaling, we may assume that $\langle b_{0},b_{0}\rangle_{\mathbb{R}^{n,1}}=-1$. By applying a Lorentz transformation, we may assume that $b_{0}=(1,0,\ldots,0)\in\mathbb{R}^{n,1}$.
-For each point $p\in\Sigma$, we know that $T(p)\wedge\nu(p)\in\text{\rm span}\{b_{0}\wedge b_{1},b_{0}\wedge b_{2},\ldots,b_{0}\wedge b_{n}\}$, hence $b_{0}\in\text{\rm span}\{T(p),\nu(p)\}$.
-Consequently, $\langle b_{0},X\rangle_{\mathbb{R}^{n,1}}=0$ for every tangent vector $X\in T\Sigma$.
+For each point $p\in\Sigma$, Lemma 4.27 implies that
+$T(p)\wedge\nu(p)\in\mathcal{Z}=\text{\rm span}\{b_{0}\wedge b_{1},b_{0}\wedge b_{2},\ldots,b_{0}\wedge b_{n}\};$
+hence $b_{0}\in\text{\rm span}\{T(p),\nu(p)\}$.
+Consequently, we have $\langle b_{0},X\rangle_{\mathbb{R}^{n,1}}=0$ for every tangent vector $X\in T\Sigma$.
 In particular, each connected component of $\Sigma$ is contained in a hyperplane in $\mathbb{R}^{n,1}$ which is orthogonal to $b_{0}$.
-For each point $p\in\Sigma$, we can find a spacelike unit vector $B(p)\in\mathbb{R}^{n,1}$ such that $B(p)$ is orthogonal to $b_{0}$ and $T(p)\wedge\nu(p)=b_{0}\wedge B(p)$.
-This defines a smooth map $B:\Sigma\to S^{n-1}\subset\mathbb{R}^{n,1}$.
-The map $B$ can be interpreted as the unit normal to $\Sigma$ within the hyperplane orthogonal to $b_{0}$.
-Using Proposition 4.26, we obtain $\mathcal{Y}=\{\mathbb{R}\cdot(N(p),b_{0}\wedge B(p)):p\in\Sigma\}$.
-This implies that the map $B:\Sigma\to S^{n-1}\subset\mathbb{R}^{n,1}$ is the composition of the map $N:\Sigma\to S^{n-1}$ with a linear isometry of $\mathbb{R}^{n}$.
+For each point $p\in\Sigma$, there is a unique spacelike unit vector $B(p)\in\mathbb{R}^{n,1}$ such that $\langle b_{0},B(p)\rangle_{\mathbb{R}^{n,1}}=0$ and $T(p)\wedge\nu(p)=b_{0}\wedge B(p)$.
+This defines a smooth map $B$ from $\Sigma$ to the (n1)(n-1)-dimensional unit sphere
+$S^{n-1}=\{b\in\mathbb{R}^{n,1}\colon\langle b_{0},b\rangle_{\mathbb{R}^{n,1}}=0,\,\langle b,b\rangle_{\mathbb{R}^{n,1}}=1\}.$
+The map $B$ can be interpreted as the unit normal to $\Sigma$ within a hyperplane orthogonal to $b_{0}$.
+In view of Lemma 4.26, we can find a linear transformation $A\colon\mathbb{R}^{n}\to\mathcal{Z}$ such that
+$\mathcal{Y}\subset\{(\xi,A\xi)\colon\xi\in\mathbb{R}^{n}\}.$
+Using Lemma 4.27, it follows that the map $\Sigma\to\mathcal{Z},\,p\mapsto T(p)\wedge\nu(p)$ is the composition of the map $N\colon\Sigma\to S^{n-1}$ with the linear transformation $A\colon\mathbb{R}^{n}\to\mathcal{Z}$.
+This implies that the map $B\colon\Sigma\to S^{n-1}$ is the composition of the map $N\colon\Sigma\to S^{n-1}$ with a linear isometry of $\mathbb{R}^{n}$.
 In particular, $\|dB\|_{\text{\rm tr}}=\|dN\|_{\text{\rm tr}}$ at each point on $\Sigma$.
 Proposition 4.31.
-If $p$ is a point on the boundary $\Sigma$ with $\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\neq 0$, then $h=0$, $dN=0$, and $dB=0$ at the point $p$.
+Suppose that $p$ is a point on the boundary $\Sigma$ with $\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\neq 0$.
+Then $h=0$, $dN=0$, and $dB=0$ at the point $p$.
+Moreover, the second fundamental form of $\Sigma$ in $\mathbb{R}^{n,1}$ vanishes at the point $p$.
 Proof.
-Let us fix a point $p\in\Sigma$, and let $\{e_{1},\ldots,e_{n-1}\}$ denote an orthonormal basis of $T_{p}\Sigma$. Differentiating the identity $T\wedge\nu=b_{0}\wedge B$ gives
+Let $\{e_{1},\ldots,e_{n-1}\}$ denote an orthonormal basis of $T_{p}\Sigma$. Differentiating the identity $T\wedge\nu=b_{0}\wedge B$ gives
 $\sum_{j=1}^{n-1}h(e_{i},e_{j})\,T\wedge e_{j}+\sum_{j=1}^{n-1}q(e_{i},e_{j})\,e_{j}\wedge\nu=b_{0}\wedge dB(e_{i})$
-for each $i=1,\ldots,n-1$. By Proposition 4.1, we know that $q(e_{i},e_{j})=0$ for $i,j=1,\ldots,n-1$. This gives
-00 0 $=\sum_{j=1}^{n-1}h(e_{i},e_{j})\,\langle T\wedge e_{j},\nu\wedge e_{k}\rangle_{\mathbb{R}^{n,1}}$
-$=\langle b_{0}\wedge dB(e_{i}),\nu\wedge e_{k}\rangle_{\mathbb{R}^{n,1}}$
-$=\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\,\langle dB(e_{i}),e_{k}\rangle_{\mathbb{R}^{n,1}}.$
-Hence, if $\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\neq 0$ at the point $p$, then $dB=0$ at the point $p$.
-This implies $dN=0$ at the point $p$.
-Using Proposition 4.1, it follows that $h=0$ at the point $p$.
+for each $i=1,\ldots,n-1$. Proposition 4.1 implies that $q(e_{i},e_{j})=0$ for $i,j=1,\ldots,n-1$. This gives
+$0$ $=\sum_{j=1}^{n-1}h(e_{i},e_{j})\,\langle T\wedge e_{j},e_{k}\wedge\nu\rangle_{\mathbb{R}^{n,1}}+\sum_{j=1}^{n-1}q(e_{i},e_{j})\,\langle e_{j}\wedge\nu,e_{k}\wedge\nu\rangle_{\mathbb{R}^{n,1}}$
+$=\langle b_{0}\wedge dB(e_{i}),e_{k}\wedge\nu\rangle_{\mathbb{R}^{n,1}}$
+$=-\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\,\langle dB(e_{i}),e_{k}\rangle_{\mathbb{R}^{n,1}}$
+for $i,k=1,\ldots,n-1$.
+By assumption, $\langle b_{0},\nu\rangle_{\mathbb{R}^{n,1}}\neq 0$ at the point $p$.
+This implies $dB=0$ at the point $p$.
+Consequently, $dN=0$ at the point $p$.
+Using Proposition 4.1, we conclude that $h=0$ at the point $p$.
+Since $q^{\text{\rm tan}}=0$ by Proposition 4.1, it follows that the second fundamental form of $\Sigma$ in $\mathbb{R}^{n,1}$ vanishes at the point $p$.
 ∎
 Proposition 4.32.
-The composition of the isometric immersion $F\colon(M,g)\to\mathbb{R}^{n,1}$ with the projection $\pi\colon\mathbb{R}^{n,1}\to\mathbb{R}^{n}$, $(t,x)\mapsto x$, is a diffeomorphism onto a compact convex domain $K\subset\mathbb{R}^{n}$ with smooth boundary.
+Let $\pi$ denote the projection from $\mathbb{R}^{n,1}$ to
+$\mathbb{R}^{n}=\{b\in\mathbb{R}^{n,1}\colon\langle b_{0},b\rangle_{\mathbb{R}^{n,1}}=0\}.$
+The composition of the isometric immersion
+$F\colon(M,g)\to\mathbb{R}^{n,1}$
+with the projection $\pi\colon\mathbb{R}^{n,1}\to\mathbb{R}^{n}$ is a diffeomorphism from $M$ onto a compact convex domain $K\subset\mathbb{R}^{n}$ with smooth boundary.
 In particular, $F$ is an isometric embedding onto the graph of a smooth real valued function defined on $K$.
@@
 for $X,Y\in TM$.
-In other words, $g$ is the pull-back of the Euclidean metric under the immersion $\pi\circ F\colon M\to\mathbb{R}^{n}$.
+In other words, $g_{0}$ is the pull-back of the Euclidean metric under the immersion $\pi\circ F\colon M\to\mathbb{R}^{n}$.
 In particular, the metric $g_{0}$ is flat.
 It follows from Proposition 4.31 that the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ coincides with the second fundamental form of $\partial M$ with respect to the metric $g$.
-Using Proposition 4.1, we conclude that the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ is positive semi-definite at each point on $\partial M$.
-Moreover, since $N$ has non-zero degree, we can find a point on $\partial M$ where the Jacobian determinant of $N$ is non-zero, and at that point the second fundamental of $\partial M$ with respect to $g_{0}$ is positive definite.
+Using Proposition 4.1, we conclude that the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ is weakly positive definite at each point on $\partial M$.
+Moreover, since $N$ has non-zero degree, we can find a point on $\partial M$ where the Jacobian determinant of $N$ is non-zero, and at that point the second fundamental of $\partial M$ with respect to $g_{0}$ is positive definite by Proposition 4.1.
 Theorem A.1 now implies that there exists an isometry of $(M,g_{0})$ onto a compact convex domain in $\mathbb{R}^{n}$ with smooth boundary.
@@
 In this final subsection, we treat the case when $M$ is not simply connected.
-In this case, the isometric immersion $F:M\to\mathbb{R}^{n,1}$ is only locally defined, but its differential its globally defined.
-This is sufficient to define the bundle $\mathcal{W}$, the connection $\nabla^{\mathcal{W}}$, and the map $\gamma^{\mathcal{W}}:\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$.
-The arguments in Subsection 4.4 now imply that there exists a vector $b_{0}\in\mathbb{R}^{n,1}$ such that $\langle b_{0},b_{0}\rangle_{\mathbb{R}^{n,1}}=-1$ and $\langle b_{0},dF(X)\rangle_{\mathbb{R}^{n,1}}=0$ for every tangent vector $X\in T\Sigma$.
-We now define a Riemannian metric $g_{0}$ on $M$ by
+In this case, the isometric immersion $F\colon M\to\mathbb{R}^{n,1}$ is only locally defined, but its differential is globally defined.
+This is sufficient to define the bundle $\mathcal{W}$, the connection $\nabla^{\mathcal{W}}$, and the map $\gamma^{\mathcal{W}}\colon\mathbb{R}^{n,1}\to\text{\rm End}(\mathcal{W}|_{p})$.
+The arguments in Section 4.4 now imply that there exists a vector $b_{0}\in\mathbb{R}^{n,1}$ such that
+$\langle b_{0},b_{0}\rangle_{\mathbb{R}^{n,1}}=-1\quad\text{ and }\quad\langle b_{0},dF(X)\rangle_{\mathbb{R}^{n,1}}=0$
+for every tangent vector $X\in T\Sigma$.
+As above, we define a Riemannian metric $g_{0}$ on $M$ by
 $g_{0}(X,Y)=g(X,Y)+\langle b_{0},dF(X)\rangle_{\mathbb{R}^{n,1}}\,\langle b_{0},dF(Y)\rangle_{\mathbb{R}^{n,1}}$
@@
 Moreover, the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ coincides with the second fundamental form of $\partial M$ with respect to the metric $g$.
-In particular, the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ is positive semi-definite at each point on $\partial M$, and we can find a point on $\partial M$ where the second fundamental of $\partial M$ with respect to $g_{0}$ is positive definite.
+In particular, the second fundamental form of $\partial M$ with respect to the metric $g_{0}$ is weakly positive definite at each point on $\partial M$, and we can find a point on $\partial M$ where the second fundamental of $\partial M$ with respect to $g_{0}$ is positive definite.
 Using Theorem A.1, it follows that $M$ is diffeomorphic to a ball.
 This contradicts the assumption that $M$ is not simply connected.
-5. Approximating a compact, convex polytope by smooth domains
+5. Approximating a compact convex polytope by smooth domains
 Throughout this section, we fix an integer $n\geq 2$.
-We further assume that $\Omega\subset\mathbb{R}^{n}$ is a compact, convex polytope with non-empty interior.
-We write $\Omega=\bigcap_{i\in I}\{u_{i}\leq 0\}$, where $I$ is a finite set and the $u_{i}$ are linear functions on $\mathbb{R}^{n}$.
+We further assume that $\Omega\subset\mathbb{R}^{n}$ is a compact convex polytope with nonempty interior.
+We write $\Omega=\bigcap_{i\in I}\{u_{i}\leq 0\}$, where $I$ is a finite set and $u_{i}$, $i\in I$, is a collection of linear functions on $\mathbb{R}^{n}$.
 After eliminating redundant inequalities, we may assume that the following condition is satisfied.
 Assumption 5.1.
-For each $i_{0}\in I$, the set
-$\{u_{i_{0}}>0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}\leq 0\}$
-is non-empty.
+For each $i_{0}\in I$, the following set is nonempty:
+$\{u_{i_{0}}>0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}\leq 0\}.$
 Let $g$ be a Riemannian metric which is defined on an open set containing $\Omega$, and let $q$ be a symmetric $(0,2)$-tensor.
-For each $i\in I$, $\nabla u_{i}$ will denote the gradient of $u_{i}$ with respect to the metric $g$, $D^{2}u_{i}$ will denote the Hessian of $u_{i}$ with respect to the metric $g$, $|\nabla u_{i}|_{g}$ will denote the norm of the gradient of $u_{i}$ with respect to the metric $g$, and $\nu_{i}=\frac{\nabla u_{i}}{|\nabla u_{i}|_{g}}$ will denote the unit normal vector field, with respect to the metric $g$, to the level sets of $u_{i}$.
+For each $i\in I$, $\nabla u_{i}$ will denote the gradient of $u_{i}$ with respect to the metric $g$, $D^{2}u_{i}$ will denote the Hessian of $u_{i}$ with respect to the metric $g$, $|\nabla u_{i}|_{g}$ will denote the norm of the gradient of $u_{i}$ with respect to the metric $g$, and $\nu_{i}={\nabla u_{i}}/{|\nabla u_{i}|_{g}}$ will denote the unit normal vector field, with respect to the metric $g$, to the level sets of $u_{i}$.
 For each $i\in I$, we denote by $N_{i}\in S^{n-1}$ the outward-pointing unit normal vector to the halfspace $\{u_{i}\leq 0\}$ with respect to the Euclidean metric.
@@
 $\Omega_{\lambda}=\bigg\{\sum_{i\in I}e^{\lambda u_{i}}\leq 1\bigg\}\subset\Omega.$
-If $\lambda$ is sufficiently large, then $\Omega_{\lambda}$ is a compact, convex domain in $\mathbb{R}^{n}$ with smooth boundary $\Sigma_{\lambda}=\partial\Omega_{\lambda}$.
+If $\lambda$ is sufficiently large, then $\Omega_{\lambda}$ is a compact convex domain in $\mathbb{R}^{n}$ with smooth boundary $\Sigma_{\lambda}=\partial\Omega_{\lambda}$.
 The sets $\Omega_{\lambda}$ form an increasing family of sets.
@@
 $\bigcup_{\lambda>\lambda_{0}}\Omega_{\lambda}=\bigcap_{i\in I}\{u_{i}<0\}.$
-In the remainder of this section, $|\cdot|_{g}$ will denote the norm taken with respect to the metric $g$, and ||$\cdot$|| ⋅ | will denote the norm taken with respect to the Euclidean metric.
+In the remainder of this section, $|\cdot|_{g}$ will denote the norm taken with respect to the metric $g$, and $|\cdot|$ will denote the norm taken with respect to the Euclidean metric.
 The outward-pointing unit normal vector to the domain $\Omega_{\lambda}$ with respect to the metric $g$ is given by
@@
 The map $N\colon\Sigma_{\lambda}\to S^{n-1}$ is homotopic to the Gauss map of $\Sigma_{\lambda}$ with respect to the Euclidean metric.
-A homotopy is induced by deforming $g$ to the Euclidean metric.
-In particular, the map $N$ has degree $1$.
-Proposition 5.1 (Prop. 3.9 in [Brendle]).
+A homotopy can be constructed by deforming $g$ to the Euclidean metric.
+In particular, the map $N\colon\Sigma_{\lambda}\to S^{n-1}$ has degree $1$.
+Proposition 5.2 (cf. [Brendle]*Proposition 3.9).
 Let $x\in\Sigma_{\lambda}$.
@@
 $\quad+\frac{\sum_{i\in I}e^{\lambda u_{i}}\,(\Delta u_{i}-(D^{2}u_{i})(\nu,\nu))}{\big|\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}\,\nu_{i}\big|_{g}}-\frac{\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla(|\nabla u_{i}|_{g})|_{g}\,|P(N_{i})|}{\big|\sum_{i\in I}e^{\lambda u_{i}}\,|\nabla u_{i}|_{g}\,N_{i}\big|}.$
-We define a function $W_{\lambda}\colon\Sigma_{\lambda}\to\mathbb{R}$ by
-$W_{\lambda}=V_{\lambda}-|\mathrm{tr}(q)-q(\nu,\nu)|$
+We define a function $W_{\lambda}\colon\Sigma_{\lambda}\to\mathbb{R}$ by $W_{\lambda}=V_{\lambda}-|\mathrm{tr}(q)-q(\nu,\nu)|$.
 Proposition 5.3.
-Suppose that the Matching Angle Hypothesis is satisfied. Then $\sup_{\Sigma_{\lambda}}\max\{-W_{\lambda},0\}\leq o(\lambda)$ as $\lambda\to\infty$.
+Suppose that the Matching Angle Hypothesis is satisfied. Then
+$\sup_{\Sigma_{\lambda}}\max\{-W_{\lambda},0\}\leq o(\lambda)\quad\text{ as }\lambda\to\infty.$
 Proof.
-It follows from Proposition 3.10 in [Brendle] that $\sup_{\Sigma_{\lambda}}\max\{-V_{\lambda},0\}\leq o(\lambda)$ as $\lambda\to\infty$.
+It follows from [Brendle]*Proposition 3.10 that $\sup_{\Sigma_{\lambda}}\max\{-V_{\lambda},0\}\leq o(\lambda)$ as $\lambda\to\infty$.
 Since $W_{\lambda}-V_{\lambda}$ is uniformly bounded, the assertion follows.
 ∎
-In the remainder of this section, we will estimate the $L^{\sigma}$-norm $\max\{-W_{\lambda},0\}$ on $\Sigma_{\lambda}\cap B_{r}(p)$, where $\sigma\in[1,\frac{3}{2})$ is a fixed exponent and $B_{r}(p)$ denotes a Euclidean ball of radius $r$.
-We begin by recalling a basic fact about the area of convex hypersurfaces in $\mathbb{R}^{n}$.
+In the remaining part of this section, we will estimate the $L^{\sigma}$-norm $\max\{-W_{\lambda},0\}$ on $\Sigma_{\lambda}\cap B_{r}(p)$, where $\sigma\in[1,\tfrac{3}{2})$ is a fixed exponent and $B_{r}(p)$ denotes a Euclidean ball of radius $r$.
 Definition 5.4.
 Consider three pairwise distinct elements $i_{1},i_{2},i_{3}\in I$.
-We denote by $G_{\lambda}^{(i_{1},i_{2},i_{3})}$ the set of all points $x\in\Sigma_{\lambda}$ with the property that $u_{i_{1}}(x)\geq u_{i_{2}}(x)\geq u_{i_{3}}(x)$ and $u_{i_{3}}(x)\geq u_{i}(x)$ for $i\in I\setminus\{i_{1},i_{2},i_{3}\}$.
+We denote by $G_{\lambda}^{(i_{1},i_{2},i_{3})}$ the set of all points $x\in\Sigma_{\lambda}$ with the property that $u_{i_{1}}(x)\geq u_{i_{2}}(x)\geq u_{i_{3}}(x)$ and $u_{i_{3}}(x)\geq u_{i}(x)$ for each $i\in I\setminus\{i_{1},i_{2},i_{3}\}$.
 Lemma 5.5.
-For each $i\in I$, we assume that the hypersurface $\{u_{i}=0\}$ satisfies the condition $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$.
-Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$.
+For each $i\in I$, we assume that $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$, where $H$ denotes the mean curvature of the hypersurface $\{u_{i}=0\}$ with respect to the metric $g$.
+Let us fix an exponent $\sigma\in[1,\tfrac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$.
 If $\lambda r$ is sufficiently large, then
@@
 Proof.
-We adapt the proof of Lemma 3.13 in [Brendle].
-Let us consider an arbitrary point $x\in G_{\lambda}^{(i_{1},i_{2},i_{3})}$ with $u_{i_{2}}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$.
-By definition of $G_{\lambda}^{(i_{1},i_{2},i_{3})}$, it follows that $u_{i}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$ for all $i\in I\setminus\{i_{1}\}$.
-Using the identity $\sum_{i\in I}e^{\lambda u_{i}(x)}=1$, we obtain $e^{\lambda u_{i_{1}}(x)}\geq 1-C\,e^{-(\lambda r)^{\frac{1}{8}}}$, hence $u_{i_{1}}(x)\geq-C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$.
-Arguing as in [Brendle], we obtain
+We adapt the proof of [Brendle]*Lemma 3.13. Let us consider an arbitrary point
+$x\in G_{\lambda}^{(i_{1},i_{2},i_{3})}\quad\text{ with }u_{i_{2}}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}.$
+By definition of $G_{\lambda}^{(i_{1},i_{2},i_{3})}$, it follows that $u_{i}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$ for all $i\in I\setminus\{i_{1}\}$. Using the identity
+$\sum_{i\in I}e^{\lambda u_{i}(x)}=1,$
+we obtain $e^{\lambda u_{i_{1}}(x)}\geq 1-C\,e^{-(\lambda r)^{\frac{1}{8}}}$, hence $u_{i_{1}}(x)\geq-C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$. Arguing as in [Brendle], we obtain
 $V_{\lambda}\geq\frac{\Delta u_{i_{1}}-(D^{2}u_{i_{1}})(\nu_{i_{1}},\nu_{i_{1}})}{|\nabla u_{i_{1}}|_{g}}-C\lambda\,e^{-(\lambda r)^{\frac{1}{8}}}$
@@
 at the point $x$.
-Since $u_{i_{1}}(x)\geq-C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$ and $u_{i}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$ for all $i\in I\setminus\{i_{1}\}$, we can find a point $y\in\Omega$ such that $u_{i_{1}}(y)=0$ and $d(x,y)\leq C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$.
-By assumption, $H-|\mathrm{tr}(q)-q(\nu_{i_{1}},\nu_{i_{1}})|\geq 0$ at the point $y$.
+Since $u_{i_{1}}(x)\geq-C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}$ and $u_{i}(x)\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}$ for all $i\in I\setminus\{i_{1}\}$, we can find a point $y\in\Omega$ such that
+$u_{i_{1}}(y)=0\quad\text{ and }\quad d(x,y)\leq C\lambda^{-1}\,e^{-(\lambda r)^{\frac{1}{8}}}.$
+By assumption, $H-|\mathrm{tr}(q)-q(\nu_{i_{1}},\nu_{i_{1}})|\geq 0$ at the point $y$, where $H$ denotes the mean curvature of the hypersurface $\{u_{i_{1}}=0\}$ with respect to the metric $g$.
 This implies
@@
 $W_{\lambda}(x)\geq-C\lambda\,e^{-(\lambda r)^{\frac{1}{8}}}$
-for each point $x\in G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}$.
-On the other hand, the intersection $\Sigma_{\lambda}\cap B_{r}(p)$ has area at most $Cr^{n-1}$.
-Consequently,
+for each point
+$x\in G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}.$
+On the other hand, the intersection $\Sigma_{\lambda}\cap B_{r}(p)$ has area at most $Cr^{n-1}$ (see [Brendle]*Lemma 3.11). Consequently,
 $\bigg(r^{\sigma+1-n}\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\leq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}.$
@@
 Assume that the Matching Angle Hypothesis holds.
-Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$.
+Let us fix an exponent $\sigma\in[1,\tfrac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$.
 If $\lambda r$ is sufficiently large, then
-$\bigg{(}r^{\sigma+1-n}\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\geq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap\{u_{i_{3}}\leq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}$
-$\leq C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}$
+$\bigg(r^{\sigma+1-n}$ $\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\geq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap\{u_{i_{3}}\leq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\leq C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}$
 for all pairwise distinct elements $i_{1},i_{2},i_{3}\in I$.
 Proof.
-The corresponding estimate for $V_{\lambda}$ follows from Lemma 3.14 in [Brendle].
-We next observe that $|W_{\lambda}-V_{\lambda}|\leq C\leq C\lambda^{\frac{1}{8}}r^{-\frac{7}{8}}$.
+The corresponding estimate for $V_{\lambda}$ follows from [Brendle]*Lemma 3.14.
 Moreover, it is shown in [Brendle] that the set $G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{2}}\geq-\lambda^{-\frac{7}{8}}r^{\frac{1}{8}}\}\cap\{u_{i_{3}}\leq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}$ has area at most $C\,(\lambda r)^{-\frac{7}{8}}\,r^{n-1}$.
-From this, the assertion follows.
+Since $|W_{\lambda}-V_{\lambda}|\leq C\leq C\lambda^{\frac{1}{8}}r^{-\frac{7}{8}}$, the assertion follows.
 ∎
 Lemma 5.7.
-Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
+Let us fix an exponent $\sigma\in[1,\tfrac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$. If $\lambda r$ is sufficiently large, then
 $\bigg(r^{\sigma+1-n}\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{3}}\geq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\leq C\,(\lambda r)^{1-\frac{3}{2\sigma}}$
@@
 Proof.
-The corresponding estimate for $V_{\lambda}$ follows from Lemma 3.15 in [Brendle].
-We next observe that $|W_{\lambda}-V_{\lambda}|\leq C\leq C\lambda$.
+The corresponding estimate for $V_{\lambda}$ follows from [Brendle]*Lemma 3.15.
 Moreover, it is shown in [Brendle] that the set $G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap\{u_{i_{3}}\geq-\lambda^{-\frac{3}{4}}r^{\frac{1}{4}}\}\cap B_{r}(p)$ has area at most $C\,(\lambda r)^{-\frac{3}{2}}\,r^{n-1}$.
-From this, the assertion follows.
+Since $|W_{\lambda}-V_{\lambda}|\leq C\leq C\lambda$, the assertion follows.
 ∎
 Proposition 5.8.
-For each $i\in I$, we assume that the hypersurface $\{u_{i}=0\}$ satisfies the condition $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$.
+For each $i\in I$, we assume that $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$, where $H$ denotes the mean curvature of the hypersurface $\{u_{i}=0\}$ with respect to the metric $g$.
 Moreover, we assume that the Matching Angle Hypothesis is satisfied.
-Let us fix an exponent $\sigma\in[1,\frac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$.
+Let us fix an exponent $\sigma\in[1,\tfrac{3}{2})$, and let $B_{r}(p)$ denote a Euclidean ball of radius $r\leq 1$.
 If $\lambda r$ is sufficiently large, then
-$\bigg{(}r^{\sigma+1-n}\int_{\Sigma_{\lambda}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}$
-$\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}+C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}+C\,(\lambda r)^{1-\frac{3}{2\sigma}}.$
+$\bigg(r^{\sigma+1-n}$ $\int_{\Sigma_{\lambda}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}+C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}+C\,(\lambda r)^{1-\frac{3}{2\sigma}}.$
 Proof.
 Combining Lemma 5.5, Lemma 5.6, and Lemma 5.7, we conclude that
-$\bigg{(}r^{\sigma+1-n}\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}$
-$\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}+C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}+C\,(\lambda r)^{1-\frac{3}{2\sigma}}$
+$\bigg(r^{\sigma+1-n}$ $\int_{G_{\lambda}^{(i_{1},i_{2},i_{3})}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\leq C\lambda r\,e^{-(\lambda r)^{\frac{1}{8}}}+C\,(\lambda r)^{\frac{1}{8}-\frac{7}{8\sigma}}+C\,(\lambda r)^{1-\frac{3}{2\sigma}}$
 for all pairwise distinct elements $i_{1},i_{2},i_{3}\in I$.
@@
 Corollary 5.9.
-For each $i\in I$, we assume that the hypersurface $\{u_{i}=0\}$ satisfies the condition $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$.
+For each $i\in I$, we assume that $H-|\mathrm{tr}(q)-q(\nu_{i},\nu_{i})|\geq 0$ at each point in $\Omega\cap\{u_{i}=0\}$, where $H$ denotes the mean curvature of the hypersurface $\{u_{i}=0\}$ with respect to the metric $g$.
 Moreover, we assume that the Matching Angle Hypothesis is satisfied.
-Let us fix an exponent $\sigma\in[1,\frac{3}{2})$.
+Let us fix an exponent $\sigma\in[1,\tfrac{3}{2})$.
 Then
-$\sup_{p\in\mathbb{R}^{n}}\sup_{r\leq 1}\bigg{(}r^{\sigma+1-n}\int_{\Sigma_{\lambda}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}\to 0$
-as $\lambda\to\infty$.
+$\sup_{p\in\mathbb{R}^{n}}\sup_{r\leq 1}\bigg(r^{\sigma+1-n}\int_{\Sigma_{\lambda}\cap B_{r}(p)}\max\{-W_{\lambda},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\to 0\quad\text{ as }\lambda\to\infty.$
 Proof.
 Let us consider an arbitrary sequence $\lambda_{l}\to\infty$. By Proposition 5.3, we can find a sequence of positive real numbers $\delta_{l}\to 0$ such that
-$(\delta_{l}\lambda_{l})^{-1}\sup_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}\to 0$
-as $l\to\infty$. This implies
+$(\delta_{l}\lambda_{l})^{-1}\sup_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}\to 0\quad\text{ as }l\to\infty.$
+This implies
 $\sup_{p\in\mathbb{R}^{n}}\sup_{r\leq(\delta_{l}\lambda_{l})^{-1}}\bigg(r^{\sigma+1-n}$ $\int_{\Sigma_{\lambda_{l}}\cap B_{r}(p)}\max\{-W_{\lambda_{l}},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}$
-$\leq(\delta_{l}\lambda_{l})^{-1}\sup_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}\to 0$
-as $l\to\infty$. On the other hand, it follows from Proposition 5.7 that
-$\sup_{p\in\mathbb{R}^{n}}\sup_{(\delta_{l}\lambda_{l})^{-1}\leq r\leq 1}\bigg{(}r^{\sigma+1-n}\int_{\Sigma_{\lambda_{l}}\cap B_{r}(p)}\max\{-W_{\lambda_{l}},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}\to 0$
-as $l\to\infty$. Putting these facts together, the assertion follows. ∎
+$\leq(\delta_{l}\lambda_{l})^{-1}\sup_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}\to 0\quad\text{ as }l\to\infty.$
+On the other hand, it follows from Proposition 5.8 that
+$\sup_{p\in\mathbb{R}^{n}}\sup_{(\delta_{l}\lambda_{l})^{-1}\leq r\leq 1}\bigg(r^{\sigma+1-n}\int_{\Sigma_{\lambda_{l}}\cap B_{r}(p)}\max\{-W_{\lambda_{l}},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\to 0\quad\text{ as }l\to\infty.$
+Putting these facts together, the assertion follows. ∎
 6. Proof of Theorem II
-Let $\Omega$, $g$, $q$, $H$ satisfy the assumptions in Theorem II.
-Furthermore assume that $\Omega$ satisfies the Matching Angle hypothesis.
-We obtain the spinor bundle $\mathcal{S}$ and the twisted spinor bundle $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$ over an open set containing $\Omega$.
+Throughout this section, we assume that $\Omega$, $g$, $q$ satisfy the assumptions in Theorem II.
+In particular, we assume that $\Omega$ satisfies the Matching Angle hypothesis.
+We consider the spinor bundle $\mathcal{S}$ and the twisted spinor bundle $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$ over an open set containing $\Omega$.
 Let $U$ denote a Euclidean ball such that the closure of $U$ is contained in the interior of $\Omega$.
 Consider a sequence $\lambda_{l}\to\infty$.
-Given $l$, let $N^{(l)}\colon\Sigma_{\lambda_{l}}\to S^{n-1}$ denote the map from (5.1).
+Given $l$, let $N^{(l)}\colon\Sigma_{\lambda_{l}}\to S^{n-1}$ denote the map defined in (5.1).
 Recall that $N^{(l)}$ has degree $1$.
@@
 ∎
-Corollary 6.2.
-For each section $s\in C^{\infty}(\Omega_{\lambda_{l}},\mathcal{E})$, we have
-$\int_{\Omega_{\lambda_{l}}}|s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s|^{2}+C\int_{U}|s|^{2},$
-where $C$ is a constant that does not depend on $l$.
-Proof.
-If we choose $K$ to be a large multiple of $\sup_{\Omega}|q|$, then we can bound
-$(|\nabla^{\mathcal{E}}s|-K\,|s|)_{+}\leq|\nabla^{\mathcal{E},q}s|.$
-We now apply Proposition 6.1 to the function $F=(\delta^{2}+|s|^{2})^{\frac{1}{2}}$, and send $\delta\to 0$. This gives
-$\int_{\Omega_{\lambda_{l}}}|s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}(|\nabla^{\mathcal{E}}s|-K\,|s|)_{+}^{2}+C\int_{U}|s|^{2}.$
-Putting these facts together, the assertion follows. ∎
 Proposition 6.2.
@@
 ∎
+Proposition 6.3.
+We have
+$\int_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}F^{2}\leq o(1)\int_{\Omega_{\lambda_{l}}}|\nabla F|^{2}+o(1)\int_{\Omega_{\lambda_{l}}}F^{2}$
+for every smooth function $F$ on $\Omega_{\lambda_{l}}$.
+Proof.
+The statement follows by combining [Brendle]*Corollary A.7 (see also [Fefferman-Phong]) with Corollary 5.9. ∎
+Corollary 6.4.
+For each section $s\in C^{\infty}(\Omega_{\lambda_{l}},\mathcal{E})$, we have
+$\int_{\Omega_{\lambda_{l}}}|s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s|^{2}+C\int_{U}|s|^{2},$
+where $C$ is a constant that does not depend on $l$.
+Proof.
+By the triangle inequality, we can bound $|\nabla^{\mathcal{E}}s|\leq|\nabla^{\mathcal{E},q}s|+K\,|s|$, where $K$ is a constant that does not depend on $l$.
+This implies $(|\nabla^{\mathcal{E}}s|-K\,|s|)_{+}\leq|\nabla^{\mathcal{E},q}s|$.
+We now apply Proposition 6.1 to the function $F=(\delta^{2}+|s|^{2})^{\frac{1}{2}}$, and send $\delta\to 0$. This gives
+$\int_{\Omega_{\lambda_{l}}}|s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}(|\nabla^{\mathcal{E}}s|-K\,|s|)_{+}^{2}+C\int_{U}|s|^{2}.$
+Putting these facts together, the assertion follows. ∎
 Corollary 6.5.
@@
 Hence, the assertion follows from Corollary 6.4. ∎
-Proposition 6.5.
-We have
-$\int_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}F^{2}\leq o(1)\int_{\Omega_{\lambda_{l}}}|\nabla F|^{2}+o(1)\int_{\Omega_{\lambda_{l}}}F^{2}$
-for every smooth function $F$ on $\Omega_{\lambda_{l}}$.
-Proof.
-This follows from Corollary 5.8 together with the Fefferman-Phong estimate (cf. [Brendle], Appendix A). ∎
 Corollary 6.6.
@@
 Proposition 6.7.
-The following assertions hold.
+The following statements hold.
   1. (a)
@@
   3. (c)
-    $q^{\rm tan}=0$ along each of the boundary faces.
-Having established Proposition 6.7, the arguments in Subsection 4.3 imply that $(\Omega,g,q)$ can be isometrically immersed as a spacelike hypersurface in the Minkowski spacetime $\mathbb{R}^{n,1}$ such that $q$ is the second fundamental form.
+    Along each of the boundary faces, we have $q^{\rm tan}=0$.
+Having established Proposition 6.7, the arguments in Section 4.3 imply that $(\Omega,g,q)$ can be isometrically immersed as a spacelike hypersurface in the Minkowski spacetime $\mathbb{R}^{n,1}$, with second fundamental form equal to $q$.
 In the remainder of this section, we will give the proof of Proposition 6.7. We first treat the odd-dimensional case, and then the even-dimensional case.
 6.1. Proof of Proposition 6.7 in odd dimensions
-Suppose that $n\geq 3$ is an odd integer. By Proposition 3.8, we can find a non-trivial section $s^{(l)}\in C^{\infty}(\Omega_{\lambda_{l}},\mathcal{E})$ such that
+Suppose that $n\geq 3$ is an odd integer. For each $l$, Proposition 3.7 implies the existence of a non-trivial section $s^{(l)}\in C^{\infty}(\Omega_{\lambda_{l}},\mathcal{E})$ such that
 $\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}$ $\leq\tfrac{1}{2}\int_{\Omega_{\lambda_{l}}}(-\mu+|J|)|s^{(l)}|^{2}$
@@
 $\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}\leq o(1)\int_{U}|s^{(l)}|^{2}$
-if $l$ is sufficiently large. Using Corollary 6.2, we obtain
+if $l$ is sufficiently large. Using this estimate together with Corollary 6.4, we obtain
 $\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}\leq C\int_{U}|s^{(l)}|^{2}$
 if $l$ is sufficiently large. Since $\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}>0$, it follows that $\int_{U}|s^{(l)}|^{2}>0$ if $l$ is sufficiently large.
-By scaling, we can arrange that $\int_{U}|s^{(l)}|^{2}=1$ if $l$ is sufficiently large. Then
+By scaling, we may arrange that $\int_{U}|s^{(l)}|^{2}=1$ if $l$ is sufficiently large. Then
 (6.1) $\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}\leq o(1)$
@@
 $\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}\leq C.$
-After passing to a subsequence if necessary, the sequence $s^{(l)}$ converges weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$.
+Passing to a subsequence if necessary, the sequence $s^{(l)}$ converges weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$.
 The limit $s$ is defined on the interior of $\Omega$ and satisfies $\int_{U}|s|^{2}=1$.
@@
 We have
-$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\to 0$
-as $l\to\infty$.
+$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\to 0\quad\text{ as }l\to\infty.$
 Proof.
@@
 $\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}(s^{(l)}-s)|^{2}+C\int_{U}|s^{(l)}-s|^{2}.$
-Since $\nabla^{\mathcal{E},q}s=0$, we conclude that
+Since $\nabla^{\mathcal{E},q}s=0$, it follows that
 $\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\leq C\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}+C\int_{U}|s^{(l)}-s|^{2}.$
-The assertion now follows from (6.1) and the fact that $s^{(l)}\to s$ weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$. ∎
+The assertion now follows from (6.1) together with the fact that $s^{(l)}\to s$ in $L_{\text{\rm loc}}^{2}(\Omega\setminus\partial\Omega)$. ∎
 Corollary 6.9.
 We have
-$\int_{\Sigma_{\lambda_{l}}}|\chi^{(l)}s-s|^{2}\to 0$
-as $l\to\infty$. Here, $\chi^{(l)}$ denotes the boundary operator on $\Sigma_{\lambda_{l}}$.
+$\int_{\Sigma_{\lambda_{l}}}|\chi^{(l)}s-s|^{2}\to 0\quad\text{ as }l\to\infty.$
+Here, $\chi^{(l)}$ denotes the boundary operator on $\Sigma_{\lambda_{l}}$.
 Proof.
-Recall that $s^{(l)}$ satisfies the boundary condition $\chi^{(l)}s^{(l)}=s^{(l)}$ at each point on $\Sigma_{\lambda_{l}}$. Hence, the assertion follows from Lemma 6.8. ∎
-For each $i_{0}\in I$, Assumption 5.1 implies that the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is non-empty.
-Moreover, it follows from Lemma 6.9 that $s$ satisfies the boundary condition
-$\gamma(\nu_{i_{0}})\circ s=s\circ\gamma_{0}(N_{i_{0}})$
-on $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$.
+Recall that $s^{(l)}$ satisfies the boundary condition $\chi^{(l)}s^{(l)}=s^{(l)}$ at each point on $\Sigma_{\lambda_{l}}$. Hence, the assertion follows from Proposition 6.8. ∎
+For each $i_{0}\in I$, Assumption 5.1 implies that the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is nonempty (see [Brendle]*Lemma 3.2).
+Moreover, it follows from Corollary 6.9 that $s$ satisfies the boundary condition $\gamma(\nu_{i_{0}})\circ s=s\circ\gamma_{0}(N_{i_{0}})$ on $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$.
 Lemma 6.10.
@@
 The proof is analogous to the proof of Lemma 4.3. ∎
-As above, we define
-$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ each point on $\Omega$}\}.$
+As above, we define $L:=\{\sigma\in S_{0}\colon\text{\rm$s\sigma=0$ at each point in $\Omega$}\}$.
 Lemma 6.11.
@@
 Proof.
-This follows from Lemma 6.11 together with the fact that $\text{\rm span}\{N_{i}:i\in I\}=\mathbb{R}^{n}$. ∎
+This follows from Lemma 6.11 together with the fact that
+$\text{\rm span}\{N_{i}\colon i\in I\}=\mathbb{R}^{n}.\qed$
 Lemma 6.13.
@@
 $\gamma(\nu_{i_{0}})\circ s=s\circ\gamma_{0}(N_{i_{0}})$
-on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$.
-In the next step, we differentiate this relation in tangential direction.
+on the boundary face
+$\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}.$
+In the next step, we differentiate this identity in tangential direction.
+Let $p$ be a point on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$, and let $X$ be a tangent vector to that boundary face.
 Arguing as in the proof of Lemma 4.2, we obtain
-$\gamma(h(X))\circ s+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\circ s=0$
-whenever $p$ is a point on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ and $X$ is a tangent vector to that boundary face. Using Lemma 6.12, we conclude that
-$\gamma(h(X))+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p})$
-whenever $p$ is a point on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ and $X$ is a tangent vector to that boundary face.
+$\gamma(h(X))\circ s+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\circ s=0.$
+Since $s\in\text{\rm Hom}(S_{0},\mathcal{S}|_{p})$ is invertible by Lemma 6.12, it follows that
+$\gamma(h(X))+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p}).$
 Note that $i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\in\text{\rm End}(\mathcal{S}|_{p})$ is self-adjoint, while $\gamma(h(X))\in\text{\rm End}(\mathcal{S}|_{p})$ is skew-adjoint.
 Thus, we conclude that
-$\gamma(h(X))=0\in\text{\rm End}(\mathcal{S}|_{p}),\quad\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p})$
-whenever $p$ is a point on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ and $X$ is a tangent vector to that boundary face. From this, the assertion follows. ∎
+$\gamma(h(X))=0\in\text{\rm End}(\mathcal{S}|_{p})\quad\text{ and }\quad\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p}).$
+From this, the assertion follows. ∎
 6.2. Proof of Proposition 6.7 in even dimensions
-Suppose now that $n\geq 2$ is an even integer. By Corollary 3.12, we can find a non-trivial section $s^{(l)}\in H^{1}(\Omega_{\lambda_{l}},\mathcal{E})$ such that
+Suppose that $n\geq 2$ is an even integer. For each $l$, Corollary 3.11 implies the existence of a non-trivial section $s^{(l)}\in H^{1}(\Omega_{\lambda_{l}},\mathcal{E})$ such that
 $\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}$ $\leq\tfrac{1}{2}\int_{\Omega_{\lambda_{l}}}(-\mu+|J|)|s^{(l)}|^{2}$
-$+\tfrac{1}{2}\int_{\Sigma_{\lambda_{l}}}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN^{(l)}\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}.$
-Moreover, we can find an element $\hat{t}_{l}\in S^{1}$ such that $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}^{(l)})$ on $\Sigma_{\lambda_{l}}$, where the map $\hat{N}^{(l)}:\Sigma_{\lambda_{l}}\to S^{n}$ is defined by $\hat{N}^{(l)}(x)=\varphi(N^{(l)}(x),\hat{t}_{l})$ for all $x\in\Sigma_{\lambda_{l}}$.
+$\quad+\tfrac{1}{2}\int_{\Sigma_{\lambda_{l}}}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN^{(l)}\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}.$
+Moreover, for each $l$, we can find an element $\hat{t}_{l}\in S^{1}$ such that $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}^{(l)})$ on $\Sigma_{\lambda_{l}}$, where the map $\hat{N}^{(l)}\colon\Sigma_{\lambda_{l}}\to S^{n}$ is defined by $\hat{N}^{(l)}(x)=\varphi(N^{(l)}(x),\hat{t}_{l})$ for all $x\in\Sigma_{\lambda_{l}}$.
 By assumption, $\mu-|J|\geq 0$.
 Using Corollary 6.6, we obtain
-$\int_{\Sigma_{\lambda_{l}}}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN^{(l)}\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}$
-$\leq o(1)\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}+o(1)\int_{U}|s^{(l)}|^{2}.$
+$\int_{\Sigma_{\lambda_{l}}}(|\mathrm{tr}(q)-q(\nu,\nu)|$ $+\|dN^{(l)}\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}\leq o(1)\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}+o(1)\int_{U}|s^{(l)}|^{2}.$
 Putting these facts together, we conclude that
 $\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}\leq o(1)\int_{U}|s^{(l)}|^{2}$
-if $l$ is sufficiently large. Using Corollary 6.2, we obtain
+if $l$ is sufficiently large. Using this estimate together with Corollary 6.4, we obtain
 $\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}\leq C\int_{U}|s^{(l)}|^{2}$
 if $l$ is sufficiently large. Since $\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}>0$, it follows that $\int_{U}|s^{(l)}|^{2}>0$ if $l$ is sufficiently large.
-By scaling, we can arrange that $\int_{U}|s^{(l)}|^{2}=1$ if $l$ is sufficiently large. Then
+By scaling, we may arrange that $\int_{U}|s^{(l)}|^{2}=1$ if $l$ is sufficiently large. Then
 (6.2) $\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}\leq o(1)$
-and
-$\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}\leq C.$
-After passing to a subsequence if necessary, we may assume that $\hat{t}_{l}\to\hat{t}$.
+and $\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}\leq C$.
+After passing to a subsequence if necessary, we may assume that $\hat{t}_{l}\to\hat{t}$ for some $\hat{t}\in S^{1}$.
 Moreover, the sequence $s^{(l)}$ converges weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$.
@@
 We have
-$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\to 0$
-as $l\to\infty$.
+$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\to 0\quad\text{ as }l\to\infty.$
 Proof.
@@
 We have
-$\int_{\Sigma_{\lambda_{l}}}|\gamma(\nu)\circ s-s\circ\gamma_{0}(\hat{N}^{(l)})|^{2}\to 0$
-as $l\to\infty$. Here, $\chi^{(l)}$ denotes the boundary operator on $\Sigma_{\lambda_{l}}$.
+$\int_{\Sigma_{\lambda_{l}}}|\gamma(\nu)\circ s-s\circ\gamma_{0}(\hat{N}^{(l)})|^{2}\to 0\quad\text{ as }l\to\infty.$
 Proof.
 Recall that $s^{(l)}$ satisfies the boundary condition $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}^{(l)})$ on $\Sigma_{\lambda_{l}}$ at each point on $\Sigma_{\lambda_{l}}$.
-Hence, the assertion follows from Lemma 6.14.
+Hence, the assertion follows from Proposition 6.14.
 ∎
 For each $i\in I$, we define $\hat{N}_{i}=\varphi(N_{i},\hat{t})\in S^{n}$.
-For each $i_{0}\in I$, Assumption 5.1 implies that the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is non-empty.
-Moreover, it follows from Lemma 6.15 that $s$ satisfies the boundary condition
+For each $i_{0}\in I$, Assumption 5.1 implies that the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is nonempty (see [Brendle]*Lemma 3.2).
+Moreover, it follows from Corollary 6.15 that $s$ satisfies the boundary condition
 $\gamma(\nu_{i_{0}})\circ s=s\circ\gamma_{0}(\hat{N}_{i_{0}})$
 on $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$.
-By Lemma 3.2 in [Brendle], the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is a dense subset of $\Omega\cap\{u_{i_{0}}=0\}$.
+By [Brendle]*Lemma 3.2, the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is a dense subset of $\Omega\cap\{u_{i_{0}}=0\}$.
 Thus, we conclude that
@@
 Proof.
-We consider a point $p\in\Omega\cap\{u_{i_{1}}=0\}\cap\{u_{i_{2}}=0\}$. Then
+By assumption, we can find a point $p\in\Omega\cap\{u_{i_{1}}=0\}\cap\{u_{i_{2}}=0\}$. Then
 $\gamma(\nu_{i_{1}})\circ s=s\circ\gamma_{0}(\hat{N}_{i_{1}})$
@@
 $=\langle\hat{N}_{i_{1}},\hat{N}_{i_{2}}\rangle\,|s|^{2}$
+at the point $p$. On the other hand, the Matching Angle Hypothesis implies that
+$\langle\nu_{i_{1}},\nu_{i_{2}}\rangle_{g}\,|s|^{2}=\langle N_{i_{1}},N_{i_{2}}\rangle\,|s|^{2}$
 at the point $p$.
-On the other hand, the Matching Angle Hypothesis implies that $\langle\nu_{i_{1}},\nu_{i_{2}}\rangle_{g}\,|s|^{2}=\langle N_{i_{1}},N_{i_{2}}\rangle\,|s|^{2}$ at the point $p$.
 Putting these facts together, we conclude that $\langle N_{i_{1}},N_{i_{2}}\rangle\,|s|^{2}=\langle\hat{N}_{i_{1}},\hat{N}_{i_{2}}\rangle\,|s|^{2}$ at the point $p$.
@@
 Proof.
-Suppose that the assertion is false. By Lemma 3.9, the map $\varphi(\cdot,\hat{t})$ is Lipschitz continuous with a Lipschitz constant strictly less than $1$. This contradicts Lemma 6.16. ∎
+[…]
+$|N_{i_{1}}-N_{i_{2}}|^{2}>|\varphi(N_{i_{1}},\hat{t})-\varphi(N_{i_{2}},\hat{t})|^{2}=|\hat{N}_{i_{1}}-\hat{N}_{i_{2}}|^{2}$
+for $i_{1}\neq i_{2}$. This contradicts Lemma 6.16. ∎
 Lemma 6.18.
@@
 The proof is analogous to the proof of Lemma 4.3. ∎
-We again define
-$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ each point on $\Omega$}\}.$
+We again define $L:=\{\sigma\in S_{0}\colon\text{\rm$s\sigma=0$ at each point in $\Omega$}\}$.
 Lemma 6.19.
@@
 Proof.
-Note that $\text{\rm span}\{N_{i}:i\in I\}=\mathbb{R}^{n}$.
-Using Lemma 6.17, it follows that $\text{\rm span}\{\hat{N}_{i}:i\in I\}=\mathbb{R}^{n}\subset\mathbb{R}^{n+1}$.
-By Lemma 6.19, $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n}$.
-Since the volume element acts as a scalar, we deduce that $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n+1}$.
+Note that $\text{\rm span}\{N_{i}\colon i\in I\}=\mathbb{R}^{n}$. Using Lemma 6.17, it follows that
+$\text{\rm span}\{\hat{N}_{i}\colon i\in I\}=\mathbb{R}^{n}\subset\mathbb{R}^{n+1}.$
+By Lemma 6.19, $L$ is invariant under $\gamma_{0}(\xi)$ for all $\xi\in\mathbb{R}^{n}$.
+Since the volume element acts as a scalar multiple of the identity, we deduce that $L$ is invariant under $\gamma_{0}(\xi)$ for all $\xi\in\mathbb{R}^{n+1}$.
 Arguing as in the proof of Lemma 4.5, the assertion follows.
@@
   • $g$ is flat.
-  • The second fundamental form of $\partial M$ is positive semi-definite.
-  • There exists a point in $\partial M$, where the second fundamental form is positive definite.
+  • The second fundamental form of $\partial M$ is weakly positive definite.
+  • There exists a point in $\partial M$, where the second fundamental form is strictly positive definite.
 Then $(M,g)$ is isometric to a compact convex domain in $\mathbb{R}^{n}$ with smooth boundary. In particular, $M$ is diffeomorphic to a ball.
 Proof.
-For small enough $\varepsilon>0$, we consider the smooth manifold
+For $\varepsilon>0$ small enough, we consider the smooth manifold
 $M^{\prime}:=M\cup_{\sim}(\partial M\times(-\varepsilon,\infty)),$
 where $\sim$ identifies $\partial M\times(-\varepsilon,0]\subset\partial M\times(-\varepsilon,\infty)$ with an open neighborhood of $\partial M\subset M$ by the normal exponential map along $\partial M$ with respect to $g$.
-Denote the metric on $\partial M$ induced by $g$ by $g_{0}$ and the second fundamental form of $\partial M$ (considered as a $(1,1)$-tensor field) by $h$.
-The Gauss lemma, the Riccati equation for the second fundamental form and the vanishing of the curvature of $M$ imply that on $\partial M\times(-\varepsilon,0]$ the metric $g$ takes the form $g=dt^{2}+g_{t}$ where $g_{t}$ is the metric on $\partial M$ given by $g_{t}(X,Y)=g_{0}((\textup{id}+th)^{2}X,Y)$.
+We denote the metric on $\partial M$ induced by $g$ by $g_{0}$ and the second fundamental form of $\partial M$ (considered as a $(1,1)$-tensor field) by $h$.
+The Gauss lemma, the Riccati equation for the second fundamental form and the vanishing of the curvature of $M$ imply that on $\partial M\times(-\varepsilon,0]$ the metric $g$ takes the form $g=dt\otimes dt+g_{t}$ where $g_{t}$ is the metric on $\partial M$ given by $g_{t}(X,Y)=g_{0}((\textup{id}+th)^{2}X,Y)$.
 Here $t\in(-\varepsilon,\infty)$ is the standard parameter.
-Since $h$ is positive semi-definite, we can use this formula to extend $g$ to $\partial M\times[0,\infty)$ and hence to a smooth metric $g^{\prime}$ on all of $M^{\prime}$.
-By Theorem 7.2 in [BGM05], the manifold $(M^{\prime},g^{\prime})$ is flat.
-Since the projection $(\partial M\times[0,\infty),dt^{2}+g_{t})\to(\partial M,g_{0})$ is $1$-Lipschitz, it is easy to see that $(M^{\prime},g^{\prime})$ is complete.
-Let $\tilde{M}^{\prime}$ be the universal covering of $M^{\prime}$ equipped with the complete flat Riemannian metric $\tilde{g}^{\prime}$ induced by $g^{\prime}$.
-Since $(\tilde{M}^{\prime},\tilde{g}^{\prime})$ is complete, flat and simply connected, it is isometric to Euclidean space $\mathbb{R}^{n}$.
-Using this isometry, we regard the preimage $\tilde{M}\subset\tilde{M}^{\prime}$ of $M\subset M^{\prime}$ as a closed subset of $\mathbb{R}^{n}$ with smooth boundary $\partial\tilde{M}$.
-Since the embedding $M\hookrightarrow M^{\prime}$ is a homotopy equivalence, $\tilde{M}$ is connected and $\tilde{M}$ is the universal covering of $M$.
-We first assume that $n\geq 3$.
-By assumption, the second fundamental form of $\partial\tilde{M}\subset\mathbb{R}^{n}$ is positive semi-definite and there exists a connected component $\Sigma\subset\partial\tilde{M}$ with the property that the second fundamental form is positive definite at some point on $\Sigma$.
-This implies that the sectional curvature of the induced Riemannian metric on $\Sigma\subset\tilde{M}$ is non-negative and positive at some point.
-This Riemannian metric of $\Sigma$ is complete because it is pulled back from a compact manifold.
-By the main theorem in [dCL] (see also [Sacksteder]), $\Sigma=\partial K\subset\mathbb{R}^{n}$ where $K\subset\mathbb{R}^{n}$ is convex closed subset in $\mathbb{R}^{n}$ with smooth boundary.
-Furthermore, $\Sigma$ is diffeomorphic to $S^{n-1}$ or to $\mathbb{R}^{n-1}$, hence it is the universal covering of a component of $\partial M$ with the induced metric.
-In the case $\Sigma\approx\mathbb{R}^{n-1}$, the Cheeger-Gromoll splitting theorem (Theorem 3 in [CG]) implies that the induced metric on $\Sigma$ is flat, which is a contradiction.
-Hence $\Sigma\approx S^{n-1}$ and $K$ is a compact convex domain in $\mathbb{R}^{n}$ with smooth boundary.
-As $\tilde{M}$ is connected and complete, we have $\tilde{M}\subset K$ as a closed, hence compact subset.
-Assume $K\setminus\tilde{M}\neq\emptyset$.
-Then there exists a connected component $\Sigma^{\prime}\neq\Sigma$ of $\partial\tilde{M}$ and, again by [dCL], $\Sigma^{\prime}=\partial K^{\prime}$ for a convex closed subset $K^{\prime}\subset{\rm int}(K)\subset\mathbb{R}^{n}$.
-Since $\Sigma^{\prime}\subset\mathrm{int}(K)$, we have that $K^{\prime}\subset\mathrm{int}(K)$.
-In particular, $K^{\prime}\subset K$ and hence $K^{\prime}$ is compact.
-Arguing as before, we obtain $\tilde{M}\subset K^{\prime}$.
-Thus $\tilde{M}\subset\mathrm{int}(K)$, which is a contradiction.
-Hence $\tilde{M}=K$ and $\tilde{M}$ is diffeomorphic to a ball. By the Brouwer fixed point theorem, we have $M=\tilde{M}$. This implies the assertion.
-Now assume $n=2$.
-Then the Gauss-Bonnet formula shows that the Euler characteristic of $M$ is positive, which implies that $M$ is diffeomorphic to a $2$-ball and, in particular, simply connected.
-Since the inclusion $M\hookrightarrow M^{\prime}$ is a homotopy equivalence, we get $\tilde{M}^{\prime}=M^{\prime}$.
-In particular, $(M^{\prime},g^{\prime})$ is isometric to Euclidean space $\mathbb{R}^{2}$ and $(M,g)\subset(M^{\prime},g^{\prime})$ is isometric to a compact domain $K\subset\mathbb{R}^{2}$ with smooth, mean-convex boundary.
-This implies the assertion.
+Since $h$ is weakly positive definite, we can use this formula to extend $g$ to $\partial M\times[0,\infty)$ and hence to a smooth metric $g^{\prime}$ on all of $M^{\prime}$.
+By [Baer-Gauduchon-Moroianu]*Theorem 7.2, the manifold $(M^{\prime},g^{\prime})$ is flat.
+It is easy to see that $(M^{\prime},g^{\prime})$ is complete.
+[…]
+We define $\tilde{M}=\pi^{-1}(M)$.
+Then $\tilde{M}$ is a closed domain in $\tilde{M}^{\prime}$ with smooth boundary.
+Since the embedding $M\hookrightarrow M^{\prime}$ is a homotopy equivalence, it follows that $\tilde{M}^{\prime}$ is connected.
+We distinguish two cases:
+Case 1: Suppose that $n\geq 3$.
+By assumption, the second fundamental form of $\partial M$ is weakly positive definite at each point on $\partial M$, and is strictly positive definite at some point $p\in\partial M$.
+We denote by $\Sigma$ the connected component of $\partial M$ that contains the point $p$.
+Since $n\geq 3$, the Gauss equations imply that the sectional curvature of $\Sigma$ is nonnegative at each point on $\Sigma$, and is strictly positive at the point $p\in\Sigma$.
+Using the Cheeger-Gromoll splitting theorem [Cheeger-Gromoll]*Theorem 3, we conclude that every covering of $\Sigma$ is compact.
+In the next step, we fix a point $\tilde{p}\in\partial\tilde{M}$ with $\pi(\tilde{p})=p$.
+We denote by $\tilde{\Sigma}$ the connected component of $\partial\tilde{M}$ that contains the point $\tilde{p}$.
+Then $\tilde{\Sigma}$ is a covering of $\Sigma$.
+In view of the discussion above, $\tilde{\Sigma}$ is compact.
+[…]
+If $M$ is not simply connected, then there exists an isometry of $\tilde{M}^{\prime}$ which maps $\tilde{M}$ to itself and has no fixed points.
+This contradicts the Brouwer fixed point theorem.
+Thus, $M$ is simply connected, and so is $M^{\prime}$.
+From this, the assertion follows.
+Case 2: Suppose that $n=2$.
+In this case, the Euler characteristic of $M$ is strictly positive by the Gauss-Bonnet theorem.
+Since $M$ is connected, it follows that $M$ is diffeomorphic to a disk $B^{2}$.
+In particular, $M$ is simply connected, and so is $M^{\prime}$.
+Consequently, $(M^{\prime},g^{\prime})$ is isometric to $\mathbb{R}^{2}$ and $(M,g)$ can be identified with a compact convex domain in $R^{2}$ with smooth boundary.
 ∎
@@
 Proposition B.1.
-Let $\Omega\subset\mathbb{R}^{n}$ be a compact, connected Lipschitz domain. Assume that $U$ is a non-empty open subset of $\Omega$. Suppose that $K>0$ is a given positive real number. Then
+Let $\Omega\subset\mathbb{R}^{n}$ be a compact connected Lipschitz domain. Assume that $U$ is a nonempty open subset of $\Omega$. Suppose that $K>0$ is a given positive real number. Then
 $\int_{\Omega}F^{2}\leq C\int_{\Omega}(|\nabla F|-KF)_{+}^{2}+C\int_{U}F^{2}$
@@
 $\int_{\Omega}(|\nabla F_{j}|-KF_{j})_{+}^{2}\to 0.$
-This implies
-$\int_{\Omega}(-F_{j})_{+}^{2}\to 0.$
-Moreover, using the triangle inequality, we obtain
+Using the triangle inequality, we obtain
 $\int_{\Omega}|\nabla F_{j}|^{2}\leq C.$
-After passing to a subsequence, the sequence $F_{j}$ converges to a function $F$ in $L^{2}(\Omega)$.
+After passing to a subsequence, the sequence $F_{j}$ converges in $L^{2}(\Omega)$ to a nonnegative function $F$.
 The function $F$ satisfies $\int_{\Omega}F^{2}=1$.
-Moreover, $F=0$ almost everywhere in $U$, and $F\geq 0$ almost everywhere in $\Omega$.
-Let $\eta:\mathbb{R}^{n}\to[0,\infty)$ denote a smooth function which is supported in the unit ball and satisfies $\int_{\mathbb{R}^{n}}\eta=1$.
-For each $\varepsilon>0$, we define a smooth function $\eta_{\varepsilon}:\mathbb{R}^{n}\to[0,\infty)$ by
-$\eta_{\varepsilon}(x)=\frac{1}{\varepsilon^{n}}\eta\left(\frac{x}{\varepsilon}\right)$
+Moreover, $F=0$ almost everywhere in $U$.
+Let $\eta\colon\mathbb{R}^{n}\to[0,\infty)$ denote a smooth function which is supported in the unit ball and satisfies $\int_{\mathbb{R}^{n}}\eta=1$.
+For each $\varepsilon>0$, we define a smooth function $\eta_{\varepsilon}\colon\mathbb{R}^{n}\to[0,\infty)$ by
+$\eta_{\varepsilon}(x)=\varepsilon^{-n}\,\eta(\varepsilon^{-1}x)$
 for all $x\in\mathbb{R}^{n}$. The function $\eta_{\varepsilon}$ is supported in a ball of radius $\varepsilon$ and $\int_{\mathbb{R}^{n}}\eta_{\varepsilon}=1$.
@@
 Then $d(p,\partial\Omega)>\varepsilon$.
-We denote by $\Omega_{\varepsilon}$ the connected component of the set $\{x\in\Omega:d(x,\partial\Omega)>\varepsilon\}$ containing $p$.
+We denote by $\Omega_{\varepsilon}$ the connected component of the set
+$\{x\in\Omega\colon d(x,\partial\Omega)>\varepsilon\}$
+containing $p$.
 Let $F_{j}^{\varepsilon}\colon\Omega_{\varepsilon}\to\mathbb{R}$ denote the convolution of $F_{j}$ with $\eta_{\varepsilon}$, and let $F^{\varepsilon}\colon\Omega_{\varepsilon}\to\mathbb{R}$ denote the convolution of $F$ with $\eta_{\varepsilon}$.
@@
 Moreover, since $(|\nabla F_{j}|-KF_{j}\big)_{+}\to 0$ in $L^{2}(\Omega)$, it follows that $\eta_{\varepsilon}*(|\nabla F_{j}|-KF_{j})_{+}\to 0$ at each point in $\Omega_{\varepsilon}$.
-Putting these facts together, we conclude that
-$|\nabla F^{\varepsilon}|-KF^{\varepsilon}\leq 0$
-at each point in $\Omega_{\varepsilon}$.
+Putting these facts together, we conclude that $|\nabla F^{\varepsilon}|-KF^{\varepsilon}\leq 0$ at each point in $\Omega_{\varepsilon}$.
 On the other hand, we know that $F=0$ almost everywhere in $U$.